### v2

# Efficient Transformed Gaussian Process State-Space Models for Non-Stationary High-Dimensional Dynamical Systems

Abstract
Gaussian process state-space models (GPSSMs) offer a principled framework for learning and inference in nonlinear dynamical systems with uncertainty quantification. However, existing GPSSMs are limited by the use of multiple independent stationary Gaussian processes (GPs), leading to prohibitive computational and parametric complexity in high-dimensional settings and restricted modeling capacity for non-stationary dynamics. To address these challenges, we propose an efficient transformed Gaussian process state-space model (ETGPSSM) for scalable and flexible modeling of high-dimensional, non-stationary dynamical systems. Specifically, our ETGPSSM integrates a single shared GP with input-dependent normalizing flows, yielding an expressive implicit process prior that captures complex, non-stationary transition dynamics while significantly reducing model complexity. For the inference of the implicit process, we develop a variational inference algorithm that jointly approximates the posterior over the underlying GP and the neural network parameters defining the normalizing flows. To avoid explicit variational parameterization of the latent states, we further incorporate the ensemble Kalman filter (EnKF) into the variational framework, enabling accurate and efficient state estimation. Extensive empirical evaluations on synthetic and real-world datasets demonstrate the superior performance of our ETGPSSM in system dynamics learning, high-dimensional state estimation, and time-series forecasting, outperforming existing GPSSMs and neural network-based SSMs in terms of computational efficiency and accuracy.
Index Terms:
Gaussian process, state-space model, normalizing flows, variational inference, high-dimensional dynamical systems.
I Introduction
State-space models (SSMs) offer a versatile framework for capturing dynamical systems where latent states evolve based on internal dynamics and external inputs [1]. For systems with well-defined dynamics, Bayesian filtering methods—including the Kalman filter (KF), extended Kalman filter (EKF), ensemble Kalman filter (EnKF), and particle filter (PF)—enable sequential state estimation by integrating prior knowledge with new observations, thus facilitating reliable predictions in uncertain conditions. This capability makes SSMs highly applicable in fields such as robotics, finance, control systems, and signal processing, where modeling temporal dependencies is essential [2, 3, 4]. Yet, in many complex real-world scenarios, the lack of precise knowledge about system dynamics significantly limits traditional SSM methods, particularly for nonlinear, high-dimensional systems like climate models, physiological processes, robotics, and neural dynamics [5].
To address this challenge, numerous data-driven modeling and system identification approaches have been proposed. Deterministic models, such as deep state-space models (DSSMs), employ neural networks to learn system dynamics and have gained popularity for their ability to capture high-dimensional, nonlinear behaviors [6]. However, these models often demand extensive training datasets to mitigate overfitting, exhibit limited generalization in uncertain environments, and lack interpretability in their predictions [7]. Consequently, despite their theoretical potential, DSSMs frequently underperform in practical applications [8].
In contrast, stochastic models incorporating flexible random process priors offer a robust alternative within a Bayesian learning framework [9]. A prominent example is the class of Gaussian process state-space models (GPSSMs), which utilize Gaussian processes (GPs) to provide interpretable uncertainty quantification and inherent regularization, making them extraordinary for modeling safety-critical dynamical systems [10, 11]. Consequently, GPSSMs have been extensively utilized in various applications, including human pose and motion learning [12], robotics and control learning [13], reinforcement learning [14, 15], target tracking and navigation [16], and magnetic-field sensing [17].
Research on GPSSMs has also focused on advancing learning and inference methodologies. Early works either assumed a pre-learned GPSSM and concentrated solely on latent state inference—an impractical assumption for many real-world applications [18, 19, 20]—or employed maximum a posteriori (MAP) estimation to jointly infer the model and the latent states [12, 21], which is prone to overfitting. To address these issues, a pivotal development came with the fully Bayesian treatment of GPSSMs using particle Markov chain Monte Carlo (PMCMC) methods [22], though the inherent cubic computational complexity of GP remains prohibitive for long time series. Subsequent efforts [23, 24, 25, 26] adopted reduced-rank GP approximations to alleviate the computational demands, yet PMCMC scalability remains a challenge. This has led to a growing interest in variational inference methods [10, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39], which leverage sparse GP approximations with inducing points to enhance scalability [40, 41].
However, existing GPSSMs, which employ multiple independent GPs to model system dynamics, remain limited to low-dimensional state spaces and face fundamental challenges when scaling to high-dimensional systems. Specifically, in variational GPSSMs, the inherent cubic computational complexity of the GPs scales linearly with the state dimension, while the number of parameters grows quadratically. This computational and parametric complexity seriously hinders the applicability of GPSSMs to high-dimensional systems. Additionally, all GPSSMs inherit a fundamental limitation, namely the use of stationary random process priors for modeling system dynamics. Such stationary assumptions are often inadequate for capturing time-varying and complex transition functions, ultimately limiting GPSSMs’ ability to model complex dynamical systems accurately.
To address these challenges, we propose a novel approach for learning and inference in GPSSMs tailored to non-stationary, high-dimensional dynamical systems. Our primary contribution is the development and integration of an efficient yet flexible implicit process prior [42] for the state transition function. This prior combines the strengths of GPs, normalizing flows [43], and (Bayesian) neural networks, significantly enhancing modeling capabilities in high-dimensional state spaces. Our key contributions are summarized as follows:
  • Instead of employing multiple independent GPs for the transition function, we propose applying an efficient transformed Gaussian process (ETGP), an implicit random process that combines a shared GP with normalizing flows. Crucially, the flow parameters are made input-dependent through neural network modeling. This novel modeling approach offers two key advantages: (1) it captures complex, non-stationary transition functions that standard GP priors cannot represent, and (2) by utilizing only a single shared GP, it mitigates the computational and parametric complexities of traditional GPSSMs in high-dimensional spaces, enabling effective scaling to higher dimensions.
  • For efficient inference, we develop a scalable variational algorithm to jointly approximate the posterior distributions of both the ETGP and the latent states. Unlike standard GPs, the implicit nature of the ETGP prevents direct parametrization of the function-space posterior. To address this challenge, we explicitly approximate the posterior over the underlying GP and the associated neural network parameters, which together define the ETGP posterior. Additionally, we incorporate the EnKF into the variational inference framework, eliminating the need to parameterize the variational distributions over latent states. This hybrid formulation thus not only offers accurate state estimation but also substantially reduces the computational cost of the parameter optimization.
  • Finally, extensive empirical evaluations across system dynamics learning, high-dimensional state estimation, and time-series prediction demonstrate the superiority of our proposed method. It achieves significantly lower computational and parametric complexity than conventional GPSSMs while outperforming state-of-the-art approaches, including deep neural network-based SSMs, in terms of learning and inference accuracy.
Part of this work was presented at IEEE ICASSP 2024 [38]. The current paper introduces input-dependent normalizing flows to further enhance model flexibility and proposes an EnKF-aided variational inference algorithm for efficient state estimation. More extensive empirical validations are also included to demonstrate the superiority of the proposed method.
The remainder of this paper is structured as follows. Section II introduces the preliminaries of state-space models and outlines key challenges in existing GPSSMs. Section III presents our proposed efficient GPSSM framework for high-dimensional dynamical systems, including the associated variational inference algorithm. Experimental results are detailed in Section IV, followed by concluding remarks in Section V. Technical proofs and derivations are deferred to the Appendix.
II Preliminaries
Section II-A introduces the SSM considered in this work. A brief review of GPSSM is presented in Section II-B. The challenges associated with existing GPSSMs are discussed in detail in Section II-C.
II-A State-Space Models (SSMs)
A generic SSM describes the probabilistic relationship between the latent state ${\mathbf{x}}_{t}\in\mathbb{R}^{d_{x}}$ and the observation ${\mathbf{y}}_{t}\in\mathbb{R}^{d_{y}}$. It can be mathematically expressed as:
(Transition): ${\mathbf{x}}_{t+1}=\bm{f}({\mathbf{x}}_{t})+\mathbf{v}_{t},$ $\mathbf{v}_{t}\sim{\cal N}(\bm{0},\mathbf{Q}),$ (1a)
(Emission): ${\mathbf{y}}_{t}=\bm{C}{\mathbf{x}}_{t}+\mathbf{e}_{t},$ $\mathbf{e}_{t}\sim{\cal N}(\bm{0},\mathbf{R}).$ (1b)
Here, the latent states follow a Markov process, meaning that at any time step $t\in\mathbb{N}$, the next state ${\mathbf{x}}_{t+1}$ depends only on the current state ${\mathbf{x}}_{t}$ and the transition function $\bm{f}(\cdot):\mathbb{R}^{d_{x}}\mapsto\mathbb{R}^{d_{x}}$. In this paper, the emission function is assumed to be linear, with the coefficient matrix $\bm{C}\in\mathbb{R}^{d_{y}\times d_{x}}$. For systems with nonlinear emission functions, the latent state can be augmented to a higher dimension, transforming the system into one with a linear emission function [30], which mitigates the non-identifiability issues commonly encountered in data-driven SSMs [10]. Both the latent states and observations are assumed to be perturbed by zero-mean Gaussian noise, with covariance matrices $\mathbf{Q}$ and $\mathbf{R}$, respectively.
II-B Gaussian Process State-Space Models (GPSSMs)
As discussed in Section I, system dynamics in real-world complex scenarios are often unknown, with only a set of noisy observations, ${\mathbf{y}}_{1:T}\!=\!\{{\mathbf{y}}_{t}\}_{t=1}^{T}$, available. Consequently, there has been growing interest in adopting data-driven approaches, such as GPs [9], to model these dynamics and quantify the underlying uncertainties.
II-B1 Gaussian process (GP)
The GP is a generalization of the Gaussian distribution over infinite index sets, enabling the specification of distributions over functions $\tilde{f}(\cdot):\mathbb{R}^{d_{x}}\mapsto\mathbb{R}$. A GP is fully characterized by its mean function, often set to zero, and its covariance (kernel) function $k({\mathbf{x}},{\mathbf{x}}^{\prime})$, which includes a set of hyperparameters $\bm{\theta}_{gp}$ that need to be optimized for model selection [9]. According to the definition of a GP, the function values $\tilde{{\mathbf{f}}}=\{\tilde{f}({\mathbf{x}}_{i})\}_{i=1}^{N}\in\mathbb{R}^{N}$ at any finite set of points $\mathbf{X}=\{{\mathbf{x}}_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times d_{x}}$ follow a joint Gaussian distribution, i.e.,
$p(\tilde{{\mathbf{f}}}\mid\mathbf{X})=\mathcal{N}(\tilde{{\mathbf{f}}}\mid\bm{0},\mathbf{K}),$ (2)
where $\mathbf{K}\in\mathbb{R}^{N\times N}$ is the covariance matrix evaluated on the finite input $\mathbf{X}$ with $[\mathbf{K}]_{i,j}\!=\!k({{{\mathbf{x}}}}_{i},{{\mathbf{x}}}_{j})$.
Given the observed function values $\tilde{{\mathbf{f}}}$ at the input $\mathbf{X}$, the GP prediction distribution, $p(\tilde{f}({\mathbf{x}}_{*})|\tilde{{\mathbf{f}}},\mathbf{X})$, at any new input ${{\mathbf{x}}}_{*}$, is also Gaussian, and fully characterized by the posterior mean $\xi({{\mathbf{x}}}_{*})$ and the posterior variance $\Xi({{\mathbf{x}}}_{*})$. Concretely,
$\xi(\mathbf{x}_{*})=\mathbf{K}_{\mathbf{x}_{*},\mathbf{X}}\mathbf{K}^{-1}\tilde{{\mathbf{f}}},$ (3a)
$\Xi(\mathbf{x}_{*})=k(\mathbf{x}_{*},\mathbf{x}_{*})-\mathbf{K}_{\mathbf{x}_{*},\mathbf{X}}\mathbf{K}^{-1}\mathbf{K}_{\mathbf{x}_{*},\mathbf{X}}^{\top},$ (3b)
where $\mathbf{K}_{\mathbf{x}_{*},\mathbf{X}}$ is the cross covariance matrix evaluated on the new input ${{\mathbf{x}}}_{*}$ and the observed input ${\mathbf{X}}$.
II-B2 Gaussian process state-space model
Placing a GP prior over the transition function $\bm{f}(\cdot)$ in the classic SSM (see Eq. (1)) leads to the following salient GPSSM [10]:
$\bm{f}(\cdot)\sim\mathcal{GP}\left(\bm{0},\bm{k}(\cdot,\cdot)\right),$ (4a)
${{\mathbf{f}}}_{t}=\bm{f}(\mathbf{x}_{t-1}),$ (4b)
$\mathbf{x}_{0}\sim p(\mathbf{x}_{0}),$ (4c)
$\mathbf{x}_{t}|{{\mathbf{f}}}_{t}\sim\mathcal{N}\left({\mathbf{x}}_{t}|{{\mathbf{f}}}_{t},\mathbf{Q}\right),$ (4d)
$\mathbf{y}_{t}|\mathbf{x}_{t}\sim{\cal N}\left({\mathbf{y}}_{t}|\bm{C}\mathbf{x}_{t},\mathbf{R}\right),$ (4e)
where the initial state prior distribution $p({\mathbf{x}}_{0})$ is assumed to be known and follows a Gaussian distribution. The graphical representation is depicted in Fig. 1.
In GPSSMs, when the state dimension $d_{x}>1$, the transition $\bm{f}(\cdot):\mathbb{R}^{d_{x}}\mapsto\mathbb{R}^{d_{x}}$ is typically modeled using $d_{x}$ mutually independent GPs. Each independent GP represents a dimension-specific function ${f}_{d}(\cdot):\mathbb{R}^{d_{x}}\mapsto\mathbb{R}$, and the multivariate output is expressed as
${\mathbf{f}}_{t}=\bm{f}({\mathbf{x}}_{t-1})=\{{f}_{d}({\mathbf{x}}_{t-1})\}_{d=1}^{d_{x}}=\{\mathrm{f}_{t,d}\}_{d=1}^{d_{x}},$ (5)
where each independent GP is characterized by a distinct kernel function and associated hyperparameters [30]. The main challenge in GPSSMs is the simultaneous learning of the transition function and noise parameters, i.e., learning $[{{\bm{\theta}}}_{gp},\mathbf{Q},\mathbf{R}]$, while inferring both the GPs and the latent states of interest.
${{\mathbf{x}}}_{0}$ ${{\mathbf{x}}}_{1}$cdots $\!\!{{\mathbf{x}}}_{t-1}\!\!$ ${{\mathbf{x}}}_{t}$cdots $\cdots$ ${{\mathbf{f}}}_{1}$cdots $\!\!{{\mathbf{f}}}_{t-1}\!\!$ ${{\mathbf{f}}}_{t}$cdots ${{\mathbf{y}}}_{1}$cdots $\!\!{{\mathbf{y}}}_{t-1}\!\!$ ${{\mathbf{y}}}_{t}$cdots
Figure 1: Graphical representation of GPSSM. The white circles represent the latent variables, while the gray circles represent the observable variables. The thick horizontal bar represents a set of fully connected, mutually correlated nodes, specifically, the GP.
II-C Challenges in Existing Variational GPSSMs
Note that in Eq. (3), evaluating the GP posterior involves inverting the covariance matrix $\mathbf{K}$, resulting in a computational cost of $\mathcal{O}(N^{3})$, which becomes prohibitive for large $N$. To address this issue, variational GPSSMs typically adopt sparse GP approximations [40, 41], a ubiquitous technique in GP-based models [44, 10] due to its analytical and computational tractability. The general idea of the sparse GP involves augmenting the corresponding GP by introducing a set of inducing points $\mathbf{Z}\!=\!\{{\mathbf{z}}_{i}\}_{i=1}^{M}$ and ${\mathbf{u}}\!=\!\{\mathrm{u}_{i}\}_{i=1}^{M},M\ll N,$ that will serve as a surrogate for the GP. Here, ${\mathbf{u}}\!\sim\!{\cal N}(\bm{0},\mathbf{K}_{\mathbf{Z},\mathbf{Z}})$ and ${\mathbf{z}}_{i}\!\in\!\mathbb{R}^{d_{x}}$. Specifically, the augmented GP prior is given by [^1] With slight notation abuse, we denote the distribution of GP prior as $p(\tilde{f})$.,
$p(\tilde{f},{\mathbf{u}})=p(\tilde{f}|\mathbf{u})p(\mathbf{u}).$ (6)
We can approximate the corresponding GP posterior by explicitly specifying a free Gaussian density on the function values $\mathbf{u}$ at $M$ locations, $q(\mathbf{u})\!=\!\mathcal{N}\left(\mathbf{u}|\mathbf{m},\mathbf{S}\right)$, and using the prior conditional for the rest of the GP, i.e.,
$q(\tilde{f},{\mathbf{u}})=p(\tilde{f}|\mathbf{u})q(\mathbf{u}).$ (7)
This leads to an explicit form for the approximate GP posterior
$q(\tilde{f})=\int q(\tilde{f},{\mathbf{u}})\mathrm{d}{\mathbf{u}}={\cal N}(\tilde{f}({\mathbf{x}}_{*})\mid\xi_{p}({\mathbf{x}}_{*}),\Xi_{p}({\mathbf{x}}_{*})),$ (8)
where
$\xi_{p}({\mathbf{x}}_{*})\!=\!\mathbf{K}_{{\mathbf{x}}_{*},\mathbf{Z}}\mathbf{K}_{\mathbf{Z},\mathbf{Z}}^{-\!1}\mathbf{m},$ (9a)
$\Xi_{p}({\mathbf{x}}_{*})\!=\!\mathbf{K}_{{\mathbf{x}}_{*},{\mathbf{x}}_{*}}\!-\!\mathbf{K}_{{\mathbf{x}}_{*},\mathbf{Z}}\mathbf{K}_{\mathbf{Z},\mathbf{Z}}^{-\!1}[\mathbf{K}_{\mathbf{Z},\mathbf{Z}}\!-\!\mathbf{S}]\mathbf{K}_{\mathbf{Z},\mathbf{Z}}^{-\!1}\mathbf{K}_{{\mathbf{x}}_{*},\mathbf{Z}}^{\top},$ (9b)
resulting in a computational complexity that scales as $\mathcal{O}(M^{3})$ instead of $\mathcal{O}(N^{3})$.
Nevertheless, existing variational GPSSMs still face significant computational and modeling challenges, which are listed below.
  • Modeling multi-dimensional state transition function using multiple separate GP results in a computational complexity of $\mathcal{O}(d_{x}M^{3})$ in GPSSMs, where the complexity grows linearly with the latent state dimension $d_{x}$.
  • The variational approximation of the GP posterior introduces a number of variational parameters that scale quadratically with the state dimensionality $d_{x}$. Specifically, the variational parameters in multi-dimensional GPSSMs include inducing inputs $\{\mathbf{Z}_{d}\}_{d=1}^{d_{x}}\in\mathbb{R}^{d_{x}\times(M\times d_{x})}$, variational means $\{\mathbf{m}_{d}\}_{d=1}^{d_{x}}\in\mathbb{R}^{d_{x}\times M}$, and covariance matrices $\{\mathbf{S}_{d}\}_{d=1}^{d_{x}}\in\mathbb{R}^{d_{x}\times(M\times M)}$ for the $d_{x}$-dimensional inducing outputs $\{\mathbf{u}_{d}\}_{d=1}^{d_{x}}$. As the state dimension increases, the quadratic scaling of inducing inputs can lead to computational intractability, while the scaling of the variational means and covariance matrices further compounds the computational burden [38].
  • Existing GPSSMs commonly place a stationary GP prior over the transition function, as shown in Eq. (4a), to model a time-invariant transition function. However, this approach may fail to adequately capture non-stationary dynamics, leading to inaccuracies in scenarios where the transition function evolves over time. Moreover, the use of independent GPs for each dimension of the transition function overlooks dependencies between outputs.
In the next section, we detail our proposed method to resolve these issues.
III Efficient Transformed Gaussian Process State-Space Models
We first introduce the proposed efficient and flexible modeling framework for high-dimensional state spaces in Section III-A. Section III-B then presents the corresponding Bayesian variational inference algorithm. Finally, the details of EnKF-aided variational lower bound evaluation are provided in Section III-C.
III-A Efficient Modeling for High-Dimensional State Space
Figure 2: (2(a)) Two warped GPs, see Eq. (10). $f_{1}(\cdot)$ effectively fits the data, while $f_{2}(\cdot)$ falls short. (2(b)) Independent GPs ($\tilde{f}_{1}(\cdot)$, $\tilde{f}_{2}(\cdot)$) vs our proposed ETGP (${f}_{1}(\cdot)$, ${f}_{2}(\cdot)$) from Fig. 3. Both methods capture the underlying function reasonably well, though the GP $\tilde{f}_{1}(\cdot)$ exhibits slightly worse fit and overestimated uncertainty.
In Section II-C, we discussed the computational and modeling challenges associated with using separate GPs for each dimension of the transition function in high-dimensional state spaces. To address these issues, in [38], we proposed to warp a single shared GP using normalizing flows [43] to fit each dimension of the transition function, that is,
$\tilde{f}(\cdot)\!\sim\!\mathcal{GP}(0,k(\cdot,\cdot)),~{}f_{d}(\cdot)\!=\!{\mathbb{G}}_{{{\bm{\theta}}}_{d}}(\tilde{f}(\cdot)),~{}d\!=\!1,\ldots,d_{x},$ (10)
where each dimension-specific normalizing flow, ${{\mathbb{G}}}_{\bm{\theta}_{d}}(\cdot)$, is an element-wise, invertible and differentiable (bijective) function parameterized by $\bm{\theta}_{d}\in\Theta$. Two specific examples are provided below.
Example 1.
If ${\mathbb{G}}_{{{\bm{\theta}}}_{d}}(\cdot)$ is a simple linear flow, then we have
$f_{d}(\cdot)=\alpha_{d}\cdot\tilde{f}(\cdot)+\beta_{d},$
with ${{\bm{\theta}}}_{d}=[\alpha_{d},\beta_{d}]^{\top}\in\mathbb{R}^{2}$.
Example 2.
If ${\mathbb{G}}_{{{\bm{\theta}}}_{d}}(\cdot)$ is a Sinh-Arcsinh-Linear (SAL) flow [44, 45], we have:
$f_{d}(\cdot)=\alpha_{d}\sinh\left[\varphi_{d}\operatorname{arcsinh}\left(\tilde{f}(\cdot)\right)-\gamma_{d}\right]+\beta_{d},$
with ${{\bm{\theta}}}_{d}=[\alpha_{d},\beta_{d},\gamma_{d},\varphi_{d}]^{\top}\in\mathbb{R}^{4}$.
More complex flows can be achieved by stacking additional layers of SAL flow and/or incorporating more advanced deep neural network-based flows, such as RealNVP [46].
The primary advantage of this methodology lies in its elimination of multiple independent GPs by leveraging computationally efficient flow transformations, thereby enabling circumvent the $\mathcal{O}(d_{x}M^{3})$ computational complexity and quadratic parameter proliferation mentioned in Section II-C. Additionally, the warped GPs are dependent which could potentially alleviate the limitation of independence modeling in conventional GPSSMs.
However, the warped GP in Eq. (10) remains stationary and, being based on a single GP, may still lack the flexibility required to capture complex latent state dynamics, leading to suboptimal learning performance [38]. An intuitive illustrative example is shown in Fig. 2(a), where the learned element-wise flow $\mathbb{G}_{\bm{\theta}_{1}}$ successfully warps $\tilde{f}$ to obtain $f_{1}$, achieving accurate fitting of the underlying function’s simple pattern. However, when faced with more complex, non-stationary data trajectories, the warped GP, $f_{2}$, struggles to capture these patterns, resulting in huge modeling uncertainty. For a more comprehensive comparison, Fig. 2(b) also presents the fitting results obtained using two independent GPs and the ETGP model introduced later in this paper. Both methods effectively capture the underlying functions, although the GP $\tilde{f}_{1}(\cdot)$ demonstrates a slightly inferior fit and inflated uncertainty estimates, due to its inherent stationarity.
To address the limitations in modeling flexibility exhibited by existing methods, we draw inspiration from recent advances in adaptive parameterization [44, 47] and introduce non-stationarity and input dependence into the warped GPs. Specifically, we generate non-stationary processes $\{f_{d}(\cdot)\}_{d=1}^{d_{x}}$ by making the flow parameters depend on the input through a shallow neural network, denoted as $\textsc{nn}:\mathbb{R}^{d_{x}}\mapsto\Theta$ and parameterized by a weight vector $\mathbf{w}$. Formally, in the context of SSM, this is expressed as [^2] Note that ${\mathbf{x}}_{t}$ can always be extended to accommodate control systems incorporating a deterministic control input $\bm{c}_{t}\in\mathbb{R}^{d_{c}}$ by augmenting the latent state to $[{\mathbf{x}}_{t}^{\top},\bm{c}_{t}^{\top}]^{\top}\in\mathbb{R}^{d_{x}+d_{c}}$. For brevity, however, we omit explicit reference to $\bm{c}_{t}$ in our notation throughout this paper.:
${{\bm{\theta}}}_{t}=\textsc{nn}_{\mathbf{w}}({\mathbf{x}}_{t-1}),$ (11)
where ${{\bm{\theta}}}_{t}=\{{{\bm{\theta}}}_{t,d}\}_{d=1}^{d_{x}}$. In this case, the flow parameters vary with the input at each time step $t$, enabling the capture of local variation patterns and facilitating the modeling of time-varying transition functions. Additionally, note that the parameter amortization learning shown in Eq. (11) can avoid that the number of flow parameters increases linearly with time step $t$. Moreover, sharing neural network parameters across flow parameterizations further improves modeling efficiency.
Figure 3: ETGP transition function in SSMs
We may further impose a prior $p_{\bm{\psi}}(\mathbf{w})$ over $\mathbf{w}$ with hyperparameter $\bm{\psi}$ (e.g., $\mathbf{w}\!\sim\!{\cal N}(\mathbf{w}|\bm{0},\operatorname{diag}(\bm{\psi}))$) to regularize the neural network’s parameters and mitigate overfitting issues. In summary, our modeling approach can be expressed mathematically as follows:
$\tilde{f}(\cdot)\sim\mathcal{GP}(0,k(\cdot,\cdot)),\ \ \mathbf{w}\sim p_{\bm{\psi}}(\mathbf{w}),$ (12a)
${{\bm{\theta}}}_{t}=\textsc{nn}_{\mathbf{w}}({\mathbf{x}}_{t-1}),\ \ f_{d}(\cdot)={\mathbb{G}}_{{{\bm{\theta}}}_{t,d}}(\tilde{f}(\cdot)),\ d\!=\!1,...,{d_{x}},$ (12b)
where $\{f_{d}(\cdot)\}_{d=1}^{d_{x}}$ implicitly characterize $d_{x}$ random processes [42]. To better understand the modeling, Fig. 3 illustrates the corresponding graphical representation with $p_{\bm{\psi}}(\mathbf{w})=\delta(\mathbf{w}-\bar{\mathbf{w}})$, where $\delta(\cdot)$ is the Dirac delta measure and $\bar{\mathbf{w}}$ is a deterministic vector, and the following example extends the Example 1, providing a more concrete illustration.
Example 3.
In the case of the linear flow, we have
$\bm{f}(\cdot)=\bm{\alpha}_{t}\cdot\tilde{f}(\cdot)+\bm{\beta}_{t},$ (13a)
${{\bm{\theta}}}_{t}=\textsc{nn}_{\mathbf{w}}({\mathbf{x}}_{t-1}),\ \ \ \mathbf{w}\sim p_{\bm{\psi}}(\mathbf{w}).$ (13b)
where ${{\bm{\theta}}}_{t}\!=\![\bm{\alpha}_{t}^{\top},\bm{\beta}_{t}^{\top}]^{\top}$, and $\bm{\alpha}_{t}\!=\!\{\alpha_{t,d}\}_{d=1}^{d_{x}}$, $\bm{\beta}_{t}\!=\!\{\beta_{t,d}\}_{d=1}^{d_{x}}$. An illustrative example of function fitting using the ETGP with linear flows is depicted in Fig. 2(b).
It is noteworthy that this modeling approach, also known as the efficient transformed Gaussian process (ETGP) in the GP literature, was initially proposed to address large-scale multi-class classification problems [48]. The efficiency stems from the use of a single GP rather than multiple GPs for multi-class settings, thereby ensuring that the computational cost of GP operations remains constant with respect to the number of classes. This leads to a substantial reduction in overall computations. In this work, we adapt it to model the high-dimensional transition function in SSMs. Additionally, ETGP implicitly characterizes $d_{x}$ dependent random processes [42], a dependence that has been overlooked in prior GPSSM literature. A more detailed insight is provided below.
Corollary 1.
If $p_{\bm{\psi}}(\mathbf{w})=\delta(\mathbf{w}-\bar{\mathbf{w}})$, where $\bar{\mathbf{w}}$ is a deterministic vector, then the ETGP defined in Example 3 becomes a $d_{x}$-dimensional dependent non-stationary Gaussian process with time-dependent covariance. Specifically, at arbitrary time steps $t$ and $t^{\prime}$, we have
$\begin{bmatrix}\bm{f}({\mathbf{x}}_{t-1})\\ \bm{f}({\mathbf{x}}_{t^{\prime}-1})\end{bmatrix}\sim{\cal N}\left(\begin{bmatrix}\bm{\beta}_{t}\\ \bm{\beta}_{t^{\prime}}\end{bmatrix},\bm{\Lambda}_{t,t^{\prime}}\right),$ (14)
where the covariance matrix $\bm{\Lambda}_{t,t^{\prime}}\in\mathbb{R}^{2d_{x}\times 2d_{x}}$ is given by
$\bm{\Lambda}_{t,t^{\prime}}\!=\!\begin{bmatrix}k({\mathbf{x}}_{t-1},{\mathbf{x}}_{t-1})\cdot\bm{\alpha}_{t}\bm{\alpha}_{t}^{\top}&k({\mathbf{x}}_{t-1},{\mathbf{x}}_{t^{\prime}-1})\cdot\bm{\alpha}_{t}\bm{\alpha}_{t^{\prime}}^{\top}\\ k({\mathbf{x}}_{t^{\prime}-1},{\mathbf{x}}_{t-1})\cdot\bm{\alpha}_{t^{\prime}}\bm{\alpha}_{t}^{\top}&k({\mathbf{x}}_{{t^{\prime}}-1},{\mathbf{x}}_{{t^{\prime}}-1})\cdot\bm{\alpha}_{t^{\prime}}\bm{\alpha}_{t^{\prime}}^{\top}\end{bmatrix}.$ (15)
Proof.
This result can be built upon the Gaussianity and the linearity property of covariance, which can be found in Appendix A. ∎
With the computationally efficient and flexible function prior defined by ETGP $\bm{f}(\cdot)\!=\!\{f_{d}(\cdot)\}_{d=1}^{d_{x}}$, we can define our proposed SSM, termed ETGPSSM, compactly expressed as follows:
$\bm{f}(\cdot)\sim\mathcal{ETGP}(\tilde{f}(\cdot),p_{\bm{\psi}}(\mathbf{w})),$ (16a)
$\mathbf{x}_{0}\sim p({\mathbf{x}}_{0}),$ (16b)
${\mathbf{f}}_{t}=\bm{f}({\mathbf{x}}_{t-1}),$ (16c)
$\mathbf{x}_{t}|{\mathbf{f}}_{t}\sim\mathcal{N}({\mathbf{x}}_{t}|{\mathbf{f}}_{t},\mathbf{Q}),$ (16d)
$\mathbf{y}_{t}|\mathbf{x}_{t}\sim{\cal N}({\mathbf{y}}_{t}|\bm{C}\mathbf{x}_{t},\mathbf{R}),$ (16e)
where the ETGP prior is characterized by the associated GP $\tilde{f}(\cdot)$, neural network weight prior $p_{\bm{\psi}}(\mathbf{w})$, and the normalizing flows, see Eq. (12). In this paper, unless otherwise specified, we focus on the linear flow shown in Example 3 for simplicity and illustrative purposes. Additionally, the linear flow requires learning only $2$ parameters per dimension, making it more efficient compared to other types of normalizing flows. We also assume that the prior distribution of the initial state, $p({\mathbf{x}}_{0})$, is Gaussian and known for simplicity.
Remark 1.
In contrast to existing GPSSMs, the $\{f_{d}(\cdot)\}_{d=1}^{d_{x}}$ in the ETGPSSM are dependent, as they are both transformed random processes derived from a shared GP and common weight priors (see Corollary 1). Furthermore, as these induced random processes are non-stationary (see Corollary 1), they are well-suited for modeling time-varying transition functions.
The primary objective in ETGPSSM, as in GPSSM, is the joint estimation of model hyperparameters $\bm{\vartheta}\!=\!\{{{\bm{\theta}}}_{gp},\bm{\psi},\mathbf{Q},\mathbf{R}\}$ and inference of both the ETGP and latent states. More specifically, given a sequence of observations ${\mathbf{y}}_{1:T}=\{{\mathbf{y}}_{t}\}_{t=1}^{T}$, with $T\in\mathbb{Z}_{+}$, the task is to estimate $\bm{\vartheta}$ and compute the posterior distributions $p(\bm{f}(\cdot)|{\mathbf{y}}_{1:T})$ for the ETGP and $p({\mathbf{x}}_{0:T}|{\mathbf{y}}_{1:T})$ for the latent states, where ${\mathbf{x}}_{0:T}=\{{\mathbf{x}}_{t}\}_{t=0}^{T}$.
However, the marginal likelihood $p(\mathbf{y}_{1:T}|\bm{\vartheta})$ is generally intractable, rendering both hyperparameter estimation via model evidence maximization and posterior inference computationally infeasible. Furthermore, unlike standard GPs that admit closed-form expressions for the prior and posterior distributions (see Eqs. (2) and (3)), the ETGP constitutes an implicit stochastic process. Consequently, the marginal prior [^3] Obtaining $p(\bm{f}(\cdot))$ by marginalizing over the GP and $\mathbf{w}$ is generally intractable. and posterior distributions are analytically intractable [42], posing significant challenges for inference. The next subsections introduce a variational inference framework that enables joint learning of model hyperparameters and Bayesian inference of the ETGP and latent states.
III-B Approximate Bayesian Inference
To address the intractability of the marginal distribution $p({\mathbf{y}}_{1:T}|\bm{\vartheta})$, we resort to variational inference, which involves maximizing an evidence lower bound (ELBO) on the logarithm of the marginal likelihood [49]. The ELBO, denoted as $\mathcal{L}$, can be generally constructed as follows:
$\mathcal{L}=\mathbb{E}_{q({\mathbf{x}}_{0:T},\bm{f}(\cdot))}\left[\log\frac{p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T},\bm{f}(\cdot))}{q({\mathbf{x}}_{0:T},\bm{f}(\cdot))}\right],$ (17)
so that the difference between $\log p({\mathbf{y}}_{1:T}|\bm{\vartheta})$ and $\mathcal{L}$ equals the Kullback-Leibler (KL) divergence between the introduced variational approximation $q({\mathbf{x}}_{0:T},\bm{f}(\cdot))$ and the true posterior $p({\mathbf{x}}_{0:T},\bm{f}(\cdot)|{\mathbf{y}}_{1:T})$. We then can jointly maximize $\mathcal{L}$ with respect to $\bm{\vartheta}$ and the variational distribution $q({\mathbf{x}}_{0:T},\bm{f}(\cdot))$ to obtain the model hyperparameter estimates and the approximation of the posterior distribution $p({\mathbf{x}}_{0:T},\bm{f}(\cdot)|{\mathbf{y}}_{1:T})$.
However, as noted earlier, the absence of closed-form expressions for the ETGP’s prior and posterior distributions prevents direct expansion of the ELBO in Eq. (17). To address this challenge, rather than operating in the function space of the ETGP, we instead follow its generative construction—illustrated in Fig. 3—and infer the posterior distributions over $\tilde{f}$ and $\mathbf{w}$, which collectively characterize the ETGP posterior. Specifically, by explicitly representing $\tilde{f}$ and $\mathbf{w}$, the joint distribution of the proposed ETGPSSM is given by:
$p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T},\mathbf{w},\tilde{f})=p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T}|\mathbf{w},\tilde{f})p_{\bm{\psi}}(\mathbf{w})p(\tilde{f}),$ (18)
where $p(\tilde{f})$ represents the GP prior with a bit of notation abuse [31] and the prior of $\mathbf{w}$ is assumed to be Gaussian, $p_{\bm{\psi}}(\mathbf{w})={\cal N}(\mathbf{w}|\bm{0},\operatorname{diag}(\bm{\psi}))$, with diagonal variance. The term $p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T}|\mathbf{w},\tilde{f})$ is factorized as:
$p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T}|\mathbf{w},\tilde{f})\!=\!p({\mathbf{x}}_{0})\!\prod_{t=1}^{T}\underbracket{p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})}_{\text{likelihood}}\underbracket{p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})}_{\text{transition}},$ (19)
where the transition term is specified as:
$p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})$ $=\int p({\mathbf{x}}_{t}|{\mathbf{f}}_{t})p({\mathbf{f}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})\mathrm{d}{\mathbf{f}}_{t}$
$={\cal N}\left({\mathbf{x}}_{t}\mid\bm{\alpha}_{t}\cdot\tilde{f}({\mathbf{x}}_{t-1})+\bm{\beta}_{t},\ \mathbf{Q}\right),$ (20)
with $p({\mathbf{f}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})=\delta\left({\mathbf{f}}_{t}-(\bm{\alpha}_{t}\cdot\tilde{f}({\mathbf{x}}_{t-1})+\bm{\beta}_{t})\right)$, and $\{\bm{\alpha}_{t},\bm{\beta}_{t}\}$ being the outputs of the neural network parameterized by $\mathbf{w}$.
To infer the latent variables $\{{\mathbf{x}}_{0:T},\mathbf{w}\}$ and the GP $\tilde{f}$, we employ the following variational distribution:
$q({\mathbf{x}}_{0:T},\tilde{f},\mathbf{w})\!=\!q(\tilde{f}),q(\mathbf{w}),q({\mathbf{x}}_{0})\prod_{t=1}^{T}q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1}),$ (21)
which mirrors the generative structure of the model in Eq. (18) and assumes independence between $\tilde{f}$ and $\mathbf{w}$. Here the variational distributions of ${\mathbf{x}}_{0}$ and $\mathbf{w}$ are defined as follows:
[…]
where ${{\mathbf{m}}}_{0}\!\in\!\mathbb{R}^{d_{x}}$, lower-triangular matrix $\mathbf{L}_{0}\!\in\!\mathbb{R}^{d_{x}\times d_{x}}$, and $\{\mathbf{w},\bm{\sigma}^{2}_{\mathrm{w}}\}$ are free variational parameters.
For the variational distribution $q(\tilde{f})$, we adopt a sparse GP approximation [40, 41], as defined in Eq. (8), which helps to reduce the computational complexity from $\mathcal{O}(T^{3})$ to $\mathcal{O}(M^{3})$.
For $q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})$ in Eq. (21), rather than assuming a parameterized form—which would introduce additional variational parameters and hence increase optimization complexity [31]—we draw inspiration from recent works [37, 50] and incorporate the EnKF into the variational learning framework to infer the latent states (see the next subsection for details). Specifically, without explicitly defining the variational distribution, we adopt the following approximation:
Assumption 1.
$q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})\approx p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1},{\mathbf{y}}_{1:t}),$ (23)
which assumes that the variational distribution of ${\mathbf{x}}_{t}$ can be approximated by the corresponding filtering distribution. Additionally, we assume that the following filtering distribution is approximately equal to its one-step backward smoothing counterpart, i.e.,
Assumption 2.
$p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})\approx p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t}),$ (24)
The two assumptions are reasonable and detailed justification can be found in Appendix B.
With these two assumptions, we can significantly streamline the evaluation of the ELBO, which is summarized in the following proposition.
Proposition 1.
Under the two approximations, and all the assumed variational distributions, together with the model joint distribution in Eq. (18), the ELBO can be reformulated as follows:
$\mathcal{L}\approx$ $\mathbb{E}_{q(\mathbf{w},\tilde{f})}\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))$ (25)
$-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0})),$
where the first term (log-likelihood) can be analytically evaluated using the EnKF (detailed in Section III-C). The three KL divergence terms can also be computed in closed form, due to the Gaussian nature of the prior and variational distributions [49].
Proof.
The proof can be found in Appendix B. ∎
III-C EnKF-aided ELBO Evaluation
We now demonstrate how to evaluate $p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})$ in Eq. (25) using the EnKF. The general idea can be derived from Bayesian filtering. We first notice that
$p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})$ $=\int p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})\mathrm{d}{\mathbf{x}}_{t},$ (26)
where the prediction distribution, $p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})=$
$\int\underbracket{p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})}_{\text{transition}}\underbracket{p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})}_{\text{filtering}}\mathrm{d}{\mathbf{x}}_{t-1}.$ (27)
Consequently, we employ the EnKF, a widely used method for handling nonlinearities and high-dimensional state spaces [51], to obtain the filtering distribution $p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t})$ at each time step $t$. Further details are provided below.
III-C1 Prediction
Given the ensemble of states $\{{\mathbf{x}}_{t-1}^{(i)}\}_{i=1}^{N}$ from the posterior distribution $p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})$ at time $t-1$, and conditioning on $\mathbf{w}\sim q(\mathbf{w})$ and $\tilde{f}\sim q(\tilde{f})$, we utilize the state transition in Eq. (20) to perform the prediction step. This generates $N$ predicted samples $\{\bar{{\mathbf{x}}}_{t}^{(i)}\}_{i=1}^{N}$ as follows:
$\bar{{\mathbf{x}}}_{t}^{(i)}\sim{\cal N}\left({\mathbf{x}}_{t}\mid\bm{\alpha}_{t}\cdot\tilde{f}({\mathbf{x}}_{t-1}^{(i)})+\bm{\beta}_{t},\ \mathbf{Q}\right),i=1,2,\ldots,N.$ (28)
Therefore, the prediction distribution in Eq. (27) is approximated as:
$p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})\approx{\cal N}({\mathbf{x}}_{t}\mid\bar{\mathbf{m}}_{t},\ \bar{\mathbf{P}}_{t}),$ (29)
where
$\overline{\mathbf{m}}_{t}$ $=\frac{1}{N}\sum_{i=1}^{N}\overline{\mathbf{x}}_{t}^{(i)},$ (30a)
$\overline{\mathbf{P}}_{t}$ $=\frac{1}{N-1}\sum_{i=1}^{N}\left(\overline{\mathbf{x}}_{t}^{(i)}-\overline{\mathbf{m}}_{t}\right)\left(\overline{\mathbf{x}}_{t}^{(i)}-\overline{\mathbf{m}}_{t}\right)^{\top}.$ (30b)
III-C2 Update
Observations $\mathbf{y}_{t}$ are then assimilated into the predictive ensemble to update the state estimate. Specifically, the predictive ensemble is updated using:
$\ \ \ \ \ \ \text{{update}:}\ \ \ \ \mathbf{x}_{t}^{(i)}=\bar{\mathbf{x}}_{t}^{(i)}+\mathbf{G}_{t}(\mathbf{y}_{t}+\mathbf{e}_{t}^{(i)}-\bm{C}\bar{\mathbf{x}}_{t}^{(i)}),$ (31a)
$\text{{reparameterization}:}\ \ \mathbf{e}_{t}^{(i)}=\bm{0}+\mathbf{R}^{\frac{1}{2}}\bm{\epsilon},\ \bm{\epsilon}\!\sim\!{\cal N}(\bm{0},\bm{I}_{d_{y}}),$ (31b)
where $\mathbf{G}_{t}$ is the Kalman gain matrix given by:
$\mathbf{G}_{t}=\bar{\mathbf{P}}_{t}\bm{C}^{\top}(\bm{C}\bar{\mathbf{P}}_{t}\bm{C}^{\top}+\mathbf{R})^{-1}.$ (32)
Thus we can obtain the set of $N$ updated samples $\{{\mathbf{x}}_{t}^{(i)}\}_{i=1}^{N}$ from the filtering distribution at time step $t$. These updated samples allow us to recursively compute the samples and prediction distribution for subsequent time steps.
Remark 2.
The prediction and update steps are inherently differentiable, as the sampled latent states become differentiable with respect to the state transition function $\bm{f}(\cdot)$ and the noise covariances $\mathbf{Q}$ and $\mathbf{R}$. This property allows for a straightforward gradient-based optimization (e.g. Adam [52] in Algorithm 1), facilitating a practical learning framework for parameter estimation and model tuning [50, 37].
Algorithm 1 ETGPSSM
1:Input: $\bm{\vartheta}=\{{{\bm{\theta}}}_{gp},\bm{\psi},\mathbf{Q},\mathbf{R}\},~{}{\bm{\zeta}},~{}{\mathbf{y}}_{1:T}$, ${\mathbf{x}}_{0}^{1:N}\sim q({\mathbf{x}}_{0})$
2:while iterations not terminated do
3:     $\mathbf{w}\sim q(\mathbf{w})$, $L_{\ell}=0$
4:     for $t=1,2,\ldots,T$ do
5:         Get $\tilde{f}$ from $q(\tilde{f})$ using Eq. (8) with ${\mathbf{x}}_{*}={\mathbf{x}}_{t-1}$
6:         Get prediction samples using Eq. (28)
7:         Get empirical moments $\bar{\mathbf{m}}_{t},\bar{\mathbf{P}}_{t}$ using Eq. (30)
8:         Get Kalman gain: $\bar{\mathbf{G}}_{t}\!=\!\bar{\mathbf{P}}_{t}\bm{C}^{\top}(\bm{C}\bar{\mathbf{P}}_{t}\bm{C}^{\top}\!\!+\!\mathbf{R})^{-1}$
9:         Get updated samples using Eq. (31)
10:         Evaluate the log-likelihood using Eq. (33), and
$L_{\ell}=L_{\ell}+\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})$
11:     end for
12:     Evaluate $\mathcal{L}$ based on $L_{\ell}$ and Eq. (25)
13:     Maximize $\mathcal{L}$ and update $\bm{\vartheta}$, ${\bm{\zeta}}$ using Adam [52]
14:end while
15:Output: EnKF particles $\{{\mathbf{x}}_{0:T}^{(i)}\}_{i=1}^{N}$, model parameters $\bm{\vartheta}$, and variational parameters ${\bm{\zeta}}$.
Based on the outlined prediction and update steps, the likelihood $p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})$ in Eq. (26) can be easily evaluated. Specifically, at each time step $t$, we have:
$p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})\approx{\cal N}({\mathbf{y}}_{t}\mid\bm{C}\bar{\mathbf{m}}_{t},\ \bm{C}\bar{\mathbf{P}}_{t}\bm{C}^{\top}),$ (33)
due to the Gaussian prediction distribution, see Eq. (29), and the linear emission model, see Eq. (16e).
Now, we can evaluate our variational lower bound, $\mathcal{L}$, as defined in Eq. (25). At each step, we first utilize the reparameterization trick (see e.g. Eq. (31b)) to sample from the Gaussian variational distributions $q(\mathbf{w})$ and $q(\tilde{f})$. This allows us to obtain an unbiased estimate of the expected log-likelihood, $\mathbb{E}_{q(\mathbf{w},\tilde{f})}[\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})]$ numerically. Due to the reparameterization trick [53], $\mathcal{L}$ is differentiable with respect to the model parameters $\bm{\vartheta}=\{\bm{\theta}_{gp},\bm{\psi},\mathbf{Q},\mathbf{R}\}$ and the variational parameters ${\bm{\zeta}}=\{\mathbf{m}_{0},\mathbf{L}_{0},\mathbf{m}_{\mathrm{w}},\bm{\sigma}_{\mathrm{w}}^{2},\mathbf{m},\mathbf{S},{\vec{\mathbf{z}}}\}$. Consequently, we leverage modern differentiation tools, such as PyTorch, to automatically compute the gradients through backpropagation through time (BPTT) and apply gradient-based optimization methods (e.g., Adam) to maximize $\mathcal{L}$ [54, 52]. The detailed routine for implementing our EnKF-aided ETGPSSM is summarized in Algorithm 1.
Remark 3 (Efficiency).
The proposed ETGPSSM in Algorithm 1 improves upon existing GPSSMs in two primary aspects of efficiency. First, its computational complexity is reduced from $\mathcal{O}(d_{x}M^{3})$ to $\mathcal{O}(M^{3})$ by utilizing a single GP—while the shallow neural network adds minimal computational overhead. Second, the use of a single GP and shallow neural network prevents the quadratic growth of variational parameters, as discussed in Section II-C. Moreover, by avoiding explicit parameterization of the latent states and instead leveraging the EnKF, the proposed method further reduces parametric complexity. Together, these design choices contribute to both parameterization and computational efficiency. Further empirical validation of these improvements is provided in Section IV-B.
In the next section, we empirically evaluate the performance of ETGPSSM across various datasets.
IV Experiments
We first demonstrate the advantages of ETGPSSM in learning non-stationary dynamical systems in Section IV-A. Next, in Section IV-B, we highlight the computational and parametric efficiency and filtering performance of ETGPSSM in chaotic high-dimensional dynamical systems. Finally, Section IV-C showcases the superiors prediction performance of ETGPSSM using multiple real-world time series datasets. More experimental details can be found in the accompanying source code, which is publicly available online. [^4] https://github.com/zhidilin/gpssmProj/tree/main/high_dim_GPSSM
IV-A Non-Stationary Dynamical System Learning
            Data               AD-EnKF [50] AD-EnKF (bnn)          EnVI [37] ETGPSSM (dnn)      ETGPSSM (bnn)
Figure 4: Non-stationary kink transition function learning performance (mean $\pm$ $2\sigma$) using various methods across different levels of emission noise ($\sigma_{\mathrm{R}}^{2}\in\{0.0008,0.008,0.08,0.8\}$, from top to bottom). The blue curve () represents the learned mean function, while the red line () indicates the true system transition function. The shaded blue region depicts the confidence interval.
This subsection demonstrates the superior modeling capabilities of ETGPSSM over existing GPSSM and neural network-based methods by evaluating the learning performance on the following underlying dynamical system:
${\mathbf{x}}_{t+1}=f({\mathbf{x}}_{t})+\mathbf{v}_{t},$ $\mathbf{v}_{t}\sim{\cal N}(0,\sigma_{\mathrm{Q}}^{2}),$ (34a)
${\mathbf{y}}_{t}={\mathbf{x}}_{t}+\mathbf{e}_{t},$ $\mathbf{e}_{t}\sim{\cal N}(0,\sigma_{\mathrm{R}}^{2}),$ (34b)
where the non-stationary kink function $f({\mathbf{x}})$ is defined as:
$f({\mathbf{x}})\!=\!\underbrace{\left[0.8+({\mathbf{x}}+0.2)\left(1-\frac{5}{1+\exp(-2{\mathbf{x}})}\right)\right]}_{\text{``kink'' function}}s({\mathbf{x}})-o({\mathbf{x}})$ (35)
with a slope function:
$s({\mathbf{x}})=\begin{cases}1-0.5\exp(-0.5{\mathbf{x}}),&{\mathbf{x}}>0,\\ 1,&{\mathbf{x}}\leq 0,\end{cases}$ (36)
and an oscillatory function:
$o({\mathbf{x}})=\begin{cases}0.5\sin(8{\mathbf{x}}),&{\mathbf{x}}>0,\\ 0.5\sin(2{\mathbf{x}}),&{\mathbf{x}}\leq 0.\end{cases}$ (37)
The slope and oscillatory functions create non-stationary pattern across different regions of the “kink” function (see Fig. 4).
Using this underlying SSM, we generate training data ${\mathbf{y}}_{1:T}$ by fixing $\sigma_{\mathrm{Q}}^{2}$ at 0.05 and varying $\sigma_{\mathrm{R}}^{2}$ systematically across the set $\{0.0008,0.008,0.08,0.8\}$, resulting in four sets of $T=600$ observations each for training purposes. To demonstrate the advantages of ETGPSSM, we compare it with the state-of-the-art (SOTA) neural network-based AD-EnKFs [50] and GP-based EnVI [37]. Consistent with previous works [31, 37], we retain the true emission model and allow the transition model to be learned in all methods, ensuring a fair comparison within the latent space. The system dynamics learning results across different methods are shown in Fig. 4, where our method is represented in the last two columns. The suffixes “dnn” and “bnn” in brackets indicate the network types—Bayesian neural network (BNN) or deep neural network (DNN)—used in our model (see Fig. 3).
The results in Fig. 4 demonstrate that ETGPSSMs consistently achieve superior system dynamics learning performance. Under small observational noise levels ($\sigma_{\mathrm{R}}^{2}=\{0.0008,0.008\}$), ETGPSSMs—particularly ETGPSSM with DNN—excel in learning the dynamics of non-stationary systems. In contrast, their competitor, AD-EnKFs, fails to capture the smooth left segment of the kink function and performs well only on the more zigzag right segment. Meanwhile, EnVI, the SOTA GPSSM, fails to capture the non-stationary behavior of the system and is unable to model the zigzag region effectively. These results highlight the strength of ETGPSSMs in integrating the representational flexibility of neural networks with the GPs, making them particularly well-suited for modeling non-stationary dynamics.
When the observational noise interference becomes more significant $(\sigma_{\mathrm{R}}^{2}=\{0.08,0.8\})$, the learning performance of all methods deteriorates. This is due to the higher estimation errors in the latent states, which hinder accurate learning of the system dynamics. Nevertheless, ETGPSSMs continue to demonstrate superior learning performance. Notably, in this scenario, BNN-based methods, particularly ETGPSSM (bnn), exhibit superior performance and more reliable uncertainty quantification compared to their DNN-based counterparts. These results highlight not only the effectiveness of ETGPSSM’s non-stationary modeling capabilities but also the regularization benefits inherent in Bayesian methods, which contribute to enhanced robustness against significant observational noise.
IV-B High-Dimensional Latent State Estimation
In this subsection, we focus on latent state estimation in the Lorenz-96 system, a mathematical model for studying chaotic behavior in weather and climate dynamics [55]. The system consists of $d_{x}\in\mathbb{Z}_{+}$ coupled ordinary differential equations (ODEs):
[…]
Here the constant forcing parameter $F$ is set to $8$, resulting a fully chaotic dynamic [55], where small variations in initial conditions lead to vastly different trajectories. The dynamics model is discretized using the Euler method, and at each time step $t$, observations are generated through a linear emission model ($\bm{C}\!=\!\bm{I}$) with Gaussian additive noise, where the noise level is significant with the covariance given by $\mathbf{R}=4\bm{I}$.
We first empirically demonstrate the efficiency of ETGPSSM in high-dimensional latent state spaces, as illustrated in Fig. 5. This figure highlights both the computational efficiency and the parameter scalability of ETGPSSM in comparison to conventional GPSSMs across varying latent dimensions. The shallow neural network used in ETGPSSM is a three-layer fully connected architecture with ReLU activation and hidden dimensions of $128$ and $64$, which maps from $\mathbb{R}^{d_{x}}$ to the flow parameter space $\Theta\!\subseteq\!\mathbb{R}^{2d_{x}}$. As shown in the left panel, the number of parameters in ETGPSSM increases linearly with the latent state dimension $d_{x}$, in contrast to the quadratic parameter growth observed in GPSSMs. This linear scalability is achieved through the use of a single shared GP combined with normalizing flows, which drastically reduces the total number of parameters compared to GPSSMs that rely on independent GPs for each state dimension.
Figure 5: (Left) The number of parameters in ETGPSSM grows linearly with the state dimension $d_{x}$, whereas in GPSSM, it scales quadratically. (Right) ETGPSSM maintains low computational costs as $d_{x}$ increases, while GPSSM exhibits linear growth.
Figure 6: Latent state estimation performance for 100-dimensional Lorenz-96 system, with 9 dimensions uniformly sampled for visualization. Root mean square error (RMSE) values for different methods: 0.5468 for EnKF, 1.2028 for ETGPSSM, 1.6187 for AD-EnKF. RMSE between latent states and observations: 1.9855.
The right panel of Fig. 5 illustrates the computational efficiency of ETGPSSM by comparing the one-sample execution time for transition evaluation (i.e., Eq. (1a)). Notably, ETGPSSM maintains consistently low running times across increasing state dimensions, whereas GPSSM exhibits rapid growth in computational cost. For instance, at $d_{x}=100$, ETGPSSM completes the evaluation in 0.36 seconds, significantly outperforming GPSSM, which requires 19.56 seconds. This efficiency stems from the reduced computational complexity of ETGPSSM, which scales as $\mathcal{O}(M^{3})$ due to its use of a single GP, in contrast to GPSSM’s $\mathcal{O}(d_{x}M^{3})$ complexity. These results demonstrate that ETGPSSM achieves more scalable parameter growth with increasing state dimensionality while maintaining computational tractability in high-dimensional settings, making it a practical choice for modeling high-dimensional chaotic systems like the Lorenz-96 system.
We next fix the dimension $d_{x}\!=\!100$ in the Lorenz-96 system and generate observations of length $T\!=\!600$ for training ETGPSSM and its deep neural network-based competitor, AD-EnKF [50]. The EnVI [37], which already caused out-of-memory errors [^5] All experiments were conducted on a machine equipped with an Intel Xeon W7-3545 v6 processor (24 cores, 2.7 GHz), 128 GB of DDR5 RAM, and an NVIDIA RTX 4500 Ada GPU (24 GB VRAM), running Ubuntu 20.04 LTS. at $d_{x}\!=\!30$ with an ensemble size $N\!=\!150$, is excluded due to its prohibitive computational cost in this high-dimensional setting. The corresponding state inference performance after training is shown in Fig. 6, where the performance of EnKF is also included for comparison.
As illustrated in Fig. 6, ETGPSSM achieves a root mean square error (RMSE) of 1.2028, outperforming AD-EnKF, which yields an RMSE of 1.6187. This result highlights the superior filtering performance of ETGPSSM in high-dimensional chaotic systems. Additionally, the RMSE of the observations (1.9855) serves as a baseline, demonstrating that ETGPSSM significantly reduces the estimation error compared to the raw observations. Although EnKF yields the lowest RMSE of 0.5468, it relies on full model knowledge and does not learn system dynamics. In contrast, ETGPSSM simultaneously learns dynamics and infers latent states, offering a more versatile and effective solution for high-dimensional complex systems.
TABLE I: Prediction performance in terms of RMSE of the different models on the system identification datasets. Mean and standard deviation of the prediction results are shown as an averaged result over ten different seeds. The three lowest RMSE values are highlighted in shades of green, with a deeper one indicating lower RMSE.
Method Actuator Ball Beam Drive Dryer Gas Furnace
DKF [7] $1.204\pm 0.250$ $0.144\pm 0.005$ $0.735\pm 0.001$ $1.465\pm 0.087$ $5.589\pm 0.066$
AD-EnKF [50] $0.705\pm 0.117$ $0.057\pm 0.006$ $0.756\pm 0.114$ $0.182\pm 0.053$ $1.408\pm 0.090$
AD-EnKF (bnn) $0.705\pm 0.088$ $0.053\pm 0.007$ $0.896\pm 0.088$ $0.155\pm 0.030$ $1.361\pm 0.061$
vGPSSM [29] $1.640\pm 0.011$ $0.268\pm 0.414$ $0.740\pm 0.010$ $0.822\pm 0.002$ $3.676\pm 0.145$
CO-GPSSM [36] $0.803\pm 0.011$ $0.079\pm 0.018$ $0.736\pm 0.007$ $0.366\pm 0.146$ $1.898\pm 0.157$
PRSSM [30] $0.691\pm 0.148$ $0.074\pm 0.010$ $0.647\pm 0.057$ $0.174\pm 0.013$ $1.503\pm 0.196$
ODGPSSM [39] $0.666\pm 0.074$ $0.068\pm 0.006$ $0.708\pm 0.052$ $0.171\pm 0.011$ $1.704\pm 0.560$
VCDT [31] $0.815\pm 0.012$ $0.065\pm 0.005$ $0.735\pm 0.005$ $0.667\pm 0.266$ $2.052\pm 0.163$
EnVI [37] $0.657\pm 0.095$ $0.055\pm 0.002$ $0.703\pm 0.050$ $0.125\pm 0.017$ $1.388\pm 0.123$
ETGPSSM (dnn) $0.646\pm 0.081$ $0.050\pm 0.003$ $0.668\pm 0.092$ $0.137\pm 0.030$ $1.300\pm 0.052$
ETGPSSM (bnn) $0.646\pm 0.095$ $0.053\pm 0.002$ $0.703\pm 0.028$ $0.154\pm 0.026$ $1.313\pm 0.048$
IV-C Real-World Time Series Prediction
In this subsection, we evaluate the prediction performance of ETGPSSM on several real-world time series datasets commonly used for system identification (see Table I). The prediction RMSE results are summarized in Table I, comparing ETGPSSM against SOTA methods, including neural network-based approaches—DKF [7] and AD-EnKF [50]—as well as various GPSSMs: vGPSSM [29], CO-GPSSM [36], PRSSM [30], ODGPSSM [39], VCDT [31], and EnVI [37]. Following previous works, the latent space dimension is set to $d_{x}=4$ for all the methods.
As shown in Table I, ETGPSSM demonstrates competitive performance across all datasets. Notably, on the Actuator dataset, ETGPSSM (dnn) and ETGPSSM (bnn) both achieve an improved RMSE of 0.646, outperforming EnVI (0.657) and significantly improving upon AD-EnKFs, which record an RMSE of 0.705. Similarly, on the Ball Beam and Gas Furnace datasets, ETGPSSMs attain the first and second lowest RMSEs, highlighting their ability to accurately model system dynamics. However, on datasets such as Drive and Dryer, other GPSSMs like PRSSM and EnVI achieve slightly superior performance. This can be attributed to the relatively low latent state dimension ($d_{x}\!=\!4$), where traditional GPSSMs, such as PRSSM, ODGPSSM and EnVI, remain highly competitive, as discussed in [37]. Nonetheless, we emphasize that the proposed ETGPSSM remains highly competitive, achieving very comparable performance even in low-dimensional settings, while significantly reducing computational complexity and parameter overhead—thereby ensuring both scalability and efficiency in high-dimensional scenarios.
In comparison to neural network-based methods like AD-EnKF, the superior performance of ETGPSSM is driven by its ability to integrate the strengths of GPs and neural networks. This hybrid approach allows ETGPSSM to effectively model complex, non-stationary system dynamics while providing reliable predictions—an area where pure deep neural network-based methods often struggle. This is evident from the results, where neural network-based methods exhibit higher prediction errors. Notably, AD-EnKF (bnn) generally outperforms other network-based approaches, corroborating the benefits of Bayesian treatments, particularly their inherent regularization, which helps mitigate overfitting and improve generalization. In summary, the experimental results demonstrate that ETGPSSM is a powerful and versatile tool for real-world time series prediction.
V Conclusions
In this paper, we presented a novel approach for learning and inference in GPSSMs tailored to non-stationary, high-dimensional dynamical systems. Our key contribution is the introduction of the ETGP prior, which integrates a shared GP with input-dependent normalizing flows modeled via neural networks. This ETGP enhances flexibility in capturing complex, non-stationary transitions while significantly reducing the computational and parametric complexity of traditional GPSSMs in high-dimensional settings. To enable scalable inference, we developed a variational framework that approximates the posterior over the ETGP by separately modeling the underlying GP and neural network parameters. Additionally, we leverage the EnKF to circumvent the need for explicit parameterization of state posteriors, further improving computational efficiency. Empirical evaluations across dynamical system learning, high-dimensional state estimation, and time-series prediction tasks demonstrate that our method outperforms existing GPSSMs and deep neural network-based SSMs, while maintaining lower computational and parametric complexity.
Appendix A Proof of Corollary 1
Proof.
Similar proof of this corollary has been discussed in the literature; see [48]. For ease of reference, we also detail here. Consider two arbitrary time steps $t$ and $t^{\prime}$, we have
$\bm{f}({\mathbf{x}}_{t-1})=\bm{\alpha}_{t}\cdot\tilde{f}({\mathbf{x}}_{t-1})+\bm{\beta}_{t},$ (39)
$\bm{f}({\mathbf{x}}_{t^{\prime}-1})=\bm{\alpha}_{t^{\prime}}\cdot\tilde{f}({\mathbf{x}}_{t^{\prime}-1})+\bm{\beta}_{t^{\prime}}.$ (40)
Since ${{\bm{\theta}}}_{t}$ and ${{\bm{\theta}}}_{t^{\prime}}$ are deterministic, the joint distribution of $\bm{f}({\mathbf{x}}_{t-1})$ and $\bm{f}({\mathbf{x}}_{t^{\prime}-1})$ is Gaussian with mean $[\bm{\beta}_{t}^{\top},\bm{\beta}_{t^{\prime}}^{\top}]^{\top}$.
The covariance between $\bm{f}({\mathbf{x}}_{t-1})$ and $\bm{f}({\mathbf{x}}_{t^{\prime}-1})$ is:
$\bm{\Lambda}_{t,t^{\prime}}$ $=\bm{\alpha}_{t}\cdot\text{Cov}(\tilde{f}({\mathbf{x}}_{t-1}),\tilde{f}({\mathbf{x}}_{t^{\prime}-1}))\cdot\bm{\alpha}_{t^{\prime}}^{\top}$ (41)
$=k({\mathbf{x}}_{t-1},{\mathbf{x}}_{t^{\prime}-1})\cdot\bm{\alpha}_{t}\bm{\alpha}_{t^{\prime}}^{\top}.$ (42)
Similarly, the covariance of $\bm{f}({\mathbf{x}}_{t-1})$ with itself is:
$\text{Cov}(\bm{f}({\mathbf{x}}_{t-1}),\bm{f}({\mathbf{x}}_{t-1}))=k({\mathbf{x}}_{t-1},{\mathbf{x}}_{t-1})\cdot\bm{\alpha}_{t}\bm{\alpha}_{t}^{\top}.$ (43)
The same logic applies to $\bm{f}({\mathbf{x}}_{t^{\prime}-1})$. Therefore, the covariance matrix $\bm{\Lambda}_{t,t^{\prime}}$ is:
$\bm{\Lambda}_{t,t^{\prime}}\!=\!\begin{bmatrix}k({\mathbf{x}}_{t-1},{\mathbf{x}}_{t-1})\cdot\bm{\alpha}_{t}\bm{\alpha}_{t}^{\top}&k({\mathbf{x}}_{t-1},{\mathbf{x}}_{t^{\prime}-1})\cdot\bm{\alpha}_{t}\bm{\alpha}_{t^{\prime}}^{\top}\\ k({\mathbf{x}}_{t^{\prime}-1},{\mathbf{x}}_{t-1})\cdot\bm{\alpha}_{t^{\prime}}\bm{\alpha}_{t}^{\top}&k({\mathbf{x}}_{{t^{\prime}}-1},{\mathbf{x}}_{{t^{\prime}}-1})\cdot\bm{\alpha}_{t^{\prime}}\bm{\alpha}_{t^{\prime}}^{\top}\end{bmatrix}.$ (44)
Appendix B ELBO Derivations and Approximations
Given the joint distribution of our model in Eq. (18) and the variational distribution in Eq. (21), augmented by the sparse GP, we can formulate the ELBO according to the general definition provided in Eq. (17). The detailed derivations are presented in Eq. (45).
From Eq. (45d) to Eq. (45e), we have made the following two assumptions, similar to the work in [37]:
  • 1)
    $p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})\approx p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t})$,
  • 2)
    $q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})\approx p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1},{\mathbf{y}}_{1:t})$.
The transition from Eq. (45f) to Eq. (45g) is derived straightforwardly by applying Bayes’ theorem.
The two approximations used here are similar to those in [37], with the key difference being the focus on the ETGP in this paper rather than the GP. The rationale behind these approximations is as follows:
  • Assumption 1) assumes that the posterior distribution of the latent state at time $t-1$ based on observations up to $t-1$ is nearly identical to that based on observations up to $t$. This holds particularly well when $t$ is large, as the additional observation at time $t$ has minimal impact on the state estimate due to the accumulated information over time.
  • Assumption 2) posits that the variational distribution $q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})$ is approximated by the filtering distribution $p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1},{\mathbf{y}}_{1:t})$. Note that in variational inference, the variational distribution serves as an approximation of the true smoothing distribution $p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1},{\mathbf{y}}_{1:T})$. While there may be some estimation loss, the approximation is reasonable for long observation sequences where future observations minimally affect state estimates. This trade-off balances accuracy and computational efficiency, a crucial consideration in high-dimensional settings.
$\mathcal{L}$ $=\mathbb{E}_{q}\left[\log\frac{p({\vec{\mathbf{y}}},{\vec{\mathbf{x}}},\mathbf{w},\tilde{f},{\mathbf{u}})}{q({\vec{\mathbf{x}}},\mathbf{w},\tilde{f},{\mathbf{u}})}\right]$ (45a)
$=\mathbb{E}_{q}\left[\log\frac{p(\tilde{f},{\mathbf{u}})p(\mathbf{w})p({\mathbf{x}}_{0})\prod_{t=1}^{T}p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})}{q(\tilde{f},{\mathbf{u}})q(\mathbf{w})q({\mathbf{x}}_{0})\prod_{t=1}^{T}q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})}\right]$ (45b)
$=\mathbb{E}_{q}\left[\sum_{t=1}^{T}\log\frac{p({\mathbf{y}}_{t},{\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})}{q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})}\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0}))$ (45c)
$=\mathbb{E}_{q}\left[\sum_{t=1}^{T}\log\frac{p({\mathbf{y}}_{t},{\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})}{q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})}\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0}))$ (45d)
$\approx\mathbb{E}_{q}\left[\sum_{t=1}^{T}\log\frac{p({\mathbf{y}}_{t},{\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})}{\underbracket{p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1},{\mathbf{y}}_{1:t})}_{\text{assumption 1}}\underbracket{p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t})}_{\text{assumption 2}}}\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0}))$ (45e)
$=\mathbb{E}_{q}\left[\sum_{t=1}^{T}\log\frac{p({\mathbf{y}}_{t},{\mathbf{x}}_{t},{\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})}{p({\mathbf{x}}_{t},{\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1},{\mathbf{y}}_{t})}\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0}))$ (45f)
$=\mathbb{E}_{q(\mathbf{w},\tilde{f})}\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0}))$ (45g)

### v3

@@
 I Introduction
-State-space models (SSMs) offer a versatile framework for capturing dynamical systems where latent states evolve based on internal dynamics and external inputs [1].
-For systems with well-defined dynamics, Bayesian filtering methods—including the Kalman filter (KF), extended Kalman filter (EKF), ensemble Kalman filter (EnKF), and particle filter (PF)—enable sequential state estimation by integrating prior knowledge with new observations, thus facilitating reliable predictions in uncertain conditions.
-This capability makes SSMs highly applicable in fields such as robotics, finance, control systems, and signal processing, where modeling temporal dependencies is essential [2, 3, 4].
-Yet, in many complex real-world scenarios, the lack of precise knowledge about system dynamics significantly limits traditional SSM methods, particularly for nonlinear, high-dimensional systems like climate models, physiological processes, robotics, and neural dynamics [5].
-To address this challenge, numerous data-driven modeling and system identification approaches have been proposed.
+State-space models (SSMs) provide a flexible framework for modeling dynamical systems, where latent states evolve based on internal dynamics and external inputs [1].
+When system dynamics are fully known, Bayesian filtering methods—such as the Kalman filter (KF), extended Kalman filter (EKF), ensemble Kalman filter (EnKF), and particle filter (PF)—enable sequential state estimation by combining prior knowledge with incoming observations, yielding reliable predictions under uncertainty.
+This capability makes SSMs widely applicable in robotics, finance, and signal processing, where modeling temporal dependencies is crucial [2, 3, 4].
+However, in complex real-world systems, including nonlinear and high-dimensional problems such as climate modeling, physiological processes, and neural dynamics, the underlying dynamics may be too intricate or poorly understood to specify a physics-informed or explicitly parameterized model.
+In these cases, traditional filtering methods cannot be applied in their standard form and must instead rely on data-driven approaches to learn the dynamics from observations, complicating the task of accurate and robust state estimation [5].
+To address this challenge, a variety of data-driven system identification approaches have been proposed.
 Deterministic models, such as deep state-space models (DSSMs), employ neural networks to learn system dynamics and have gained popularity for their ability to capture high-dimensional, nonlinear behaviors [6].
-However, these models often demand extensive training datasets to mitigate overfitting, exhibit limited generalization in uncertain environments, and lack interpretability in their predictions [7].
-Consequently, despite their theoretical potential, DSSMs frequently underperform in practical applications [8].
-In contrast, stochastic models incorporating flexible random process priors offer a robust alternative within a Bayesian learning framework [9].
-A prominent example is the class of Gaussian process state-space models (GPSSMs), which utilize Gaussian processes (GPs) to provide interpretable uncertainty quantification and inherent regularization, making them extraordinary for modeling safety-critical dynamical systems [10, 11].
-Consequently, GPSSMs have been extensively utilized in various applications, including human pose and motion learning [12], robotics and control learning [13], reinforcement learning [14, 15], target tracking and navigation [16], and magnetic-field sensing [17].
+However, these models often require large training datasets to avoid overfitting, and their generalization performance may degrade in uncertain or data-scarce environments.
+Moreover, their predictions can be difficult to interpret, especially when compared to more structured probabilistic models [7].
+As a result, DSSMs may face challenges in practical deployment despite their modeling capacity [8].
+Stochastic models based on flexible random process priors provide a principled alternative within a Bayesian learning framework [9].
+A prominent example is the class of Gaussian process state-space models (GPSSMs), which leverage Gaussian processes (GPs) to capture uncertainty and provide implicit regularization through prior structure.
+These properties make GPSSMs attractive for applications where uncertainty quantification and data efficiency are critical [10, 11].
+GPSSMs have been applied in diverse domains, including human pose and motion learning [12], robotics and control [13], reinforcement learning [14, 15], target tracking and navigation [16], and magnetic-field sensing [17].
 Research on GPSSMs has also focused on advancing learning and inference methodologies.
 Early works either assumed a pre-learned GPSSM and concentrated solely on latent state inference—an impractical assumption for many real-world applications [18, 19, 20]—or employed maximum a posteriori (MAP) estimation to jointly infer the model and the latent states [12, 21], which is prone to overfitting.
-To address these issues, a pivotal development came with the fully Bayesian treatment of GPSSMs using particle Markov chain Monte Carlo (PMCMC) methods [22], though the inherent cubic computational complexity of GP remains prohibitive for long time series.
+To address these issues, a important development came with the fully Bayesian treatment of GPSSMs using particle Markov chain Monte Carlo (PMCMC) methods [22], though the inherent cubic computational complexity of GP remains prohibitive for long time series.
 Subsequent efforts [23, 24, 25, 26] adopted reduced-rank GP approximations to alleviate the computational demands, yet PMCMC scalability remains a challenge.
-This has led to a growing interest in variational inference methods [10, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39], which leverage sparse GP approximations with inducing points to enhance scalability [40, 41].
+This has led to a growing interest in variational inference methods [10, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] that can leverage sparse GP approximations with inducing points to enhance scalability [40, 41].
 However, existing GPSSMs, which employ multiple independent GPs to model system dynamics, remain limited to low-dimensional state spaces and face fundamental challenges when scaling to high-dimensional systems.
-Specifically, in variational GPSSMs, the inherent cubic computational complexity of the GPs scales linearly with the state dimension, while the number of parameters grows quadratically.
-This computational and parametric complexity seriously hinders the applicability of GPSSMs to high-dimensional systems.
-Additionally, all GPSSMs inherit a fundamental limitation, namely the use of stationary random process priors for modeling system dynamics.
-Such stationary assumptions are often inadequate for capturing time-varying and complex transition functions, ultimately limiting GPSSMs’ ability to model complex dynamical systems accurately.
+Specifically, in variational GPSSMs, each output dimension of the state transition function is modeled by a separate GP, leading to a total computational cost that scales linearly with the state dimensionality.
+Since each GP has a cubic complexity in the number of data points due to covariance matrix inversion, the overall computational burden becomes prohibitive.
+Moreover, the number of model parameters grows quadratically with the state dimension, further exacerbating scalability issues.
+Additionally, all GPSSMs rely on stationary GP priors, which assume that the statistical properties of the transition function are invariant under input translation.
+This assumption often fails for systems with time-varying or non-stationary dynamics, limiting the ability of GPSSMs to accurately capture complex real-world behaviors.
 To address these challenges, we propose a novel approach for learning and inference in GPSSMs tailored to non-stationary, high-dimensional dynamical systems.
-Our primary contribution is the development and integration of an efficient yet flexible implicit process prior [42] for the state transition function.
+Our primary contribution is the development and integration of an efficient yet flexible implicit process prior [42, 43] for the state transition function.
 This prior combines the strengths of GPs, normalizing flows [44], and (Bayesian) neural networks, significantly enhancing modeling capabilities in high-dimensional state spaces.
 Our key contributions are summarized as follows:
-  • Instead of employing multiple independent GPs for the transition function, we propose applying an efficient transformed Gaussian process (ETGP), an implicit random process that combines a shared GP with normalizing flows.
+  • Instead of employing multiple independent GPs for the transition function, we propose applying an efficient transformed Gaussian process (ETGP) [42], an implicit random process that combines a shared GP with normalizing flows.
 Crucially, the flow parameters are made input-dependent through neural network modeling.
@@
 To address this challenge, we explicitly approximate the posterior over the underlying GP and the associated neural network parameters, which together define the ETGP posterior.
-Additionally, we incorporate the EnKF into the variational inference framework, eliminating the need to parameterize the variational distributions over latent states.
+Additionally, we derive a generic and assumption-free variational lower bound that can incorporate an EnKF into the variational inference framework, eliminating the need to parameterize the variational distributions over latent states.
 This hybrid formulation thus not only offers accurate state estimation but also substantially reduces the computational cost of the parameter optimization.
@@
 Experimental results are detailed in Section IV, followed by concluding remarks in Section V.
-Technical proofs and derivations are deferred to the Appendix.
+Supporting results and proofs are deferred to the Appendix and supplementary materials.
 II Preliminaries
@@
 II-A State-Space Models (SSMs)
-A generic SSM describes the probabilistic relationship between the latent state ${\mathbf{x}}_{t}\in\mathbb{R}^{d_{x}}$ and the observation ${\mathbf{y}}_{t}\in\mathbb{R}^{d_{y}}$.
-It can be mathematically expressed as:
+An SSM defines a probabilistic dynamical system by specifying how latent states ${\mathbf{x}}_{t}\in\mathbb{R}^{d_{x}}$ evolve over time and generate observations ${\mathbf{y}}_{t}\in\mathbb{R}^{d_{y}}$.
+It is commonly expressed as:
 (Transition): ${\mathbf{x}}_{t+1}=\bm{f}({\mathbf{x}}_{t})+\mathbf{v}_{t},$ $\mathbf{v}_{t}\sim{\cal N}(\bm{0},\mathbf{Q}),$ (1a)
 (Emission): ${\mathbf{y}}_{t}=\bm{C}{\mathbf{x}}_{t}+\mathbf{e}_{t},$ $\mathbf{e}_{t}\sim{\cal N}(\bm{0},\mathbf{R}).$ (1b)
-Here, the latent states follow a Markov process, meaning that at any time step $t\in\mathbb{N}$, the next state ${\mathbf{x}}_{t+1}$ depends only on the current state ${\mathbf{x}}_{t}$ and the transition function $\bm{f}(\cdot):\mathbb{R}^{d_{x}}\mapsto\mathbb{R}^{d_{x}}$.
-In this paper, the emission function is assumed to be linear, with the coefficient matrix $\bm{C}\in\mathbb{R}^{d_{y}\times d_{x}}$.
-For systems with nonlinear emission functions, the latent state can be augmented to a higher dimension, transforming the system into one with a linear emission function [30], which mitigates the non-identifiability issues commonly encountered in data-driven SSMs [10].
-Both the latent states and observations are assumed to be perturbed by zero-mean Gaussian noise, with covariance matrices $\mathbf{Q}$ and $\mathbf{R}$, respectively.
+The latent states evolve according to a first-order Markov process, meaning that at any time step $t\in\mathbb{N}$, the next state ${\mathbf{x}}_{t+1}$ depends only on the current state ${\mathbf{x}}_{t}$ through the transition function $\bm{f}(\cdot):\mathbb{R}^{d_{x}}\mapsto\mathbb{R}^{d_{x}}$.
+In this work, the emission function is assumed to be linear, parameterized by a matrix $\bm{C}\in\mathbb{R}^{d_{y}\times d_{x}}$.
+For systems with nonlinear emissions, the latent state can be augmented to a higher-dimensional representation, effectively rendering the emission function linear [30].
+This reparameterization helps mitigate non-identifiability issues frequently encountered in data-driven SSMs [10].
+Both latent states and observations are perturbed by zero-mean Gaussian noise, with covariance matrices $\mathbf{Q}$ and $\mathbf{R}$, respectively.
 II-B Gaussian Process State-Space Models (GPSSMs)
@@
 The GP is a generalization of the Gaussian distribution over infinite index sets, enabling the specification of distributions over functions $\tilde{f}(\cdot):\mathbb{R}^{d_{x}}\mapsto\mathbb{R}$.
-A GP is fully characterized by its mean function, often set to zero, and its covariance (kernel) function $k({\mathbf{x}},{\mathbf{x}}^{\prime})$, which includes a set of hyperparameters $\bm{\theta}_{gp}$ that need to be optimized for model selection [9].
-According to the definition of a GP, the function values $\tilde{{\mathbf{f}}}=\{\tilde{f}({\mathbf{x}}_{i})\}_{i=1}^{N}\in\mathbb{R}^{N}$ at any finite set of points $\mathbf{X}=\{{\mathbf{x}}_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times d_{x}}$ follow a joint Gaussian distribution, i.e.,
+A GP is fully characterized by its mean function, often set to zero, and its covariance (kernel) function $k({\mathbf{x}},{\mathbf{x}}^{\prime})$, which includes a set of hyperparameters $\bm{\theta}_{gp}$ optimized during model selection [9].
+By definition, the function values $\tilde{{\mathbf{f}}}=\{\tilde{f}({\mathbf{x}}_{i})\}_{i=1}^{N}\in\mathbb{R}^{N}$, evaluated at a finite set of inputs $\mathbf{X}=\{{\mathbf{x}}_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times d_{x}}$, follow a joint Gaussian distribution:
 $p(\tilde{{\mathbf{f}}}\mid\mathbf{X})=\mathcal{N}(\tilde{{\mathbf{f}}}\mid\bm{0},\mathbf{K}),$ (2)
-where $\mathbf{K}\in\mathbb{R}^{N\times N}$ is the covariance matrix evaluated on the finite input $\mathbf{X}$ with $[\mathbf{K}]_{i,j}\!=\!k({{{\mathbf{x}}}}_{i},{{\mathbf{x}}}_{j})$.
-Given the observed function values $\tilde{{\mathbf{f}}}$ at the input $\mathbf{X}$, the GP prediction distribution, $p(\tilde{f}({\mathbf{x}}_{*})|\tilde{{\mathbf{f}}},\mathbf{X})$, at any new input ${{\mathbf{x}}}_{*}$, is also Gaussian, and fully characterized by the posterior mean $\xi({{\mathbf{x}}}_{*})$ and the posterior variance $\Xi({{\mathbf{x}}}_{*})$.
-Concretely,
+where $\mathbf{K}\in\mathbb{R}^{N\times N}$ is the covariance matrix with entries $[\mathbf{K}]_{i,j}=k({\mathbf{x}}_{i},{\mathbf{x}}_{j})$.
+Given observed function values $\tilde{{\mathbf{f}}}$ at inputs $\mathbf{X}$, the GP prediction at a new input ${\mathbf{x}}_{*}$, denoted $p(\tilde{f}({\mathbf{x}}_{*})\mid\tilde{{\mathbf{f}}},\mathbf{X})$, is also Gaussian, with posterior mean $\xi({\mathbf{x}}_{*})$ and posterior variance $\Xi({\mathbf{x}}_{*})$ given by:
 $\xi(\mathbf{x}_{*})=\mathbf{K}_{\mathbf{x}_{*},\mathbf{X}}\mathbf{K}^{-1}\tilde{{\mathbf{f}}},$ (3a)
 $\Xi(\mathbf{x}_{*})=k(\mathbf{x}_{*},\mathbf{x}_{*})-\mathbf{K}_{\mathbf{x}_{*},\mathbf{X}}\mathbf{K}^{-1}\mathbf{K}_{\mathbf{x}_{*},\mathbf{X}}^{\top},$ (3b)
-where $\mathbf{K}_{\mathbf{x}_{*},\mathbf{X}}$ is the cross covariance matrix evaluated on the new input ${{\mathbf{x}}}_{*}$ and the observed input ${\mathbf{X}}$.
+where $\mathbf{K}_{{\mathbf{x}}_{*},\mathbf{X}}$ is the cross-covariance vector between the test input ${\mathbf{x}}_{*}$ and the training inputs $\mathbf{X}$.
 II-B2 Gaussian process state-space model
-Placing a GP prior over the transition function $\bm{f}(\cdot)$ in the classic SSM (see Eq. (1)) leads to the following salient GPSSM [10]:
+Placing a GP prior over the transition function $\bm{f}(\cdot)$ in the classic SSM (see Eq. (1)) yields the GPSSM [10]:
 $\bm{f}(\cdot)\sim\mathcal{GP}\left(\bm{0},\bm{k}(\cdot,\cdot)\right),$ (4a)
-${{\mathbf{f}}}_{t}=\bm{f}(\mathbf{x}_{t-1}),$ (4b)
+${{\mathbf{f}}}_{t}\triangleq\bm{f}(\mathbf{x}_{t-1}),$ (4b)
 $\mathbf{x}_{0}\sim p(\mathbf{x}_{0}),$ (4c)
@@
 $\mathbf{y}_{t}|\mathbf{x}_{t}\sim{\cal N}\left({\mathbf{y}}_{t}|\bm{C}\mathbf{x}_{t},\mathbf{R}\right),$ (4e)
-where the initial state prior distribution $p({\mathbf{x}}_{0})$ is assumed to be known and follows a Gaussian distribution. The graphical representation is depicted in Fig. 1.
-In GPSSMs, when the state dimension $d_{x}>1$, the transition $\bm{f}(\cdot):\mathbb{R}^{d_{x}}\mapsto\mathbb{R}^{d_{x}}$ is typically modeled using $d_{x}$ mutually independent GPs.
-Each independent GP represents a dimension-specific function ${f}_{d}(\cdot):\mathbb{R}^{d_{x}}\mapsto\mathbb{R}$, and the multivariate output is expressed as
+where the initial state $p({\mathbf{x}}_{0})$ is assumed known and Gaussian. A graphical representation is shown in Fig. 1.
+When the state dimension $d_{x}>1$, the transition function $\bm{f}(\cdot):\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{x}}$ is typically modeled using $d_{x}$ independent GPs.
+Each component function $f_{d}(\cdot):\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}$ maps to one output dimension:
 ${\mathbf{f}}_{t}=\bm{f}({\mathbf{x}}_{t-1})=\{{f}_{d}({\mathbf{x}}_{t-1})\}_{d=1}^{d_{x}}=\{\mathrm{f}_{t,d}\}_{d=1}^{d_{x}},$ (5)
-where each independent GP is characterized by a distinct kernel function and associated hyperparameters [30].
-The main challenge in GPSSMs is the simultaneous learning of the transition function and noise parameters, i.e., learning $[{{\bm{\theta}}}_{gp},\mathbf{Q},\mathbf{R}]$, while inferring both the GPs and the latent states of interest.
+where each GP has its own kernel function and hyperparameters [30].
+The central challenge in GPSSMs lies in jointly learning the transition dynamics and noise parameters, i.e., learning $[{{\bm{\theta}}}_{gp},\mathbf{Q},\mathbf{R}]$, while simultaneously inferring both the latent states and the underlying functions.
 ${{\mathbf{x}}}_{0}$ ${{\mathbf{x}}}_{1}$cdots $\!\!{{\mathbf{x}}}_{t-1}\!\!$ ${{\mathbf{x}}}_{t}$cdots $\cdots$ ${{\mathbf{f}}}_{1}$cdots $\!\!{{\mathbf{f}}}_{t-1}\!\!$ ${{\mathbf{f}}}_{t}$cdots ${{\mathbf{y}}}_{1}$cdots $\!\!{{\mathbf{y}}}_{t-1}\!\!$ ${{\mathbf{y}}}_{t}$cdots
@@
 Note that in Eq.
-(3), evaluating the GP posterior involves inverting the covariance matrix $\mathbf{K}$, resulting in a computational cost of $\mathcal{O}(N^{3})$, which becomes prohibitive for large $N$.
-To address this issue, variational GPSSMs typically adopt sparse GP approximations [40, 41], a ubiquitous technique in GP-based models [44, 10] due to its analytical and computational tractability.
-The general idea of the sparse GP involves augmenting the corresponding GP by introducing a set of inducing points $\mathbf{Z}\!=\!\{{\mathbf{z}}_{i}\}_{i=1}^{M}$ and ${\mathbf{u}}\!=\!\{\mathrm{u}_{i}\}_{i=1}^{M},M\ll N,$ that will serve as a surrogate for the GP.
-Here, ${\mathbf{u}}\!\sim\!{\cal N}(\bm{0},\mathbf{K}_{\mathbf{Z},\mathbf{Z}})$ and ${\mathbf{z}}_{i}\!\in\!\mathbb{R}^{d_{x}}$.
-Specifically, the augmented GP prior is given by [^1] With slight notation abuse, we denote the distribution of GP prior as $p(\tilde{f})$.,
+(3), evaluating the GP posterior requires inverting the covariance matrix $\mathbf{K}$, incurring a computational cost of $\mathcal{O}(N^{3})$, which becomes prohibitive for large $N$.
+To address this, variational GPSSMs typically employ sparse GP approximations [40, 41], a widely used technique in GP-based models [45, 10] due to its analytical tractability and reduced computational cost.
+The core idea is to augment the GP with a smaller set of inducing points, which act as a sparse approximation of the full GP.
+More specifically, we can introduce $M\!\ll\!N$ inducing inputs $\mathbf{Z}=\{{\mathbf{z}}_{i}\}_{i=1}^{M}$ (where ${\mathbf{z}}_{i}\in\mathbb{R}^{d_{x}}$) and their corresponding inducing outputs ${\mathbf{u}}=\{\mathrm{u}_{i}\}_{i=1}^{M}=\{\tilde{f}({\mathbf{z}}_{i})\}_{i=1}^{M}$, which inherit the same GP prior as the original process, i.e., ${\mathbf{u}}\sim\mathcal{N}(\bm{0},\mathbf{K}_{\mathbf{Z},\mathbf{Z}})$.
+The augmented GP prior can be written as: [^1] With slight notation abuse, we denote the distribution of GP prior as $p(\tilde{f})$.
 $p(\tilde{f},{\mathbf{u}})=p(\tilde{f}|\mathbf{u})p(\mathbf{u}).$ (6)
-We can approximate the corresponding GP posterior by explicitly specifying a free Gaussian density on the function values $\mathbf{u}$ at $M$ locations, $q(\mathbf{u})\!=\!\mathcal{N}\left(\mathbf{u}|\mathbf{m},\mathbf{S}\right)$, and using the prior conditional for the rest of the GP, i.e.,
+To approximate the posterior, a variational distribution is defined over the $M$ inducing outputs, $q({\mathbf{u}})=\mathcal{N}({\mathbf{u}}\mid\mathbf{m},\mathbf{S})$, while the prior conditional is used for the rest of the GP:
 $q(\tilde{f},{\mathbf{u}})=p(\tilde{f}|\mathbf{u})q(\mathbf{u}).$ (7)
-This leads to an explicit form for the approximate GP posterior
+This leads to an explicit expression for the variational approximation to the GP posterior:
 $q(\tilde{f})=\int q(\tilde{f},{\mathbf{u}})\mathrm{d}{\mathbf{u}}={\cal N}(\tilde{f}({\mathbf{x}}_{*})\mid\xi_{p}({\mathbf{x}}_{*}),\Xi_{p}({\mathbf{x}}_{*})),$ (8)
@@
 $\Xi_{p}({\mathbf{x}}_{*})\!=\!\mathbf{K}_{{\mathbf{x}}_{*},{\mathbf{x}}_{*}}\!-\!\mathbf{K}_{{\mathbf{x}}_{*},\mathbf{Z}}\mathbf{K}_{\mathbf{Z},\mathbf{Z}}^{-\!1}[\mathbf{K}_{\mathbf{Z},\mathbf{Z}}\!-\!\mathbf{S}]\mathbf{K}_{\mathbf{Z},\mathbf{Z}}^{-\!1}\mathbf{K}_{{\mathbf{x}}_{*},\mathbf{Z}}^{\top},$ (9b)
-resulting in a computational complexity that scales as $\mathcal{O}(M^{3})$ instead of $\mathcal{O}(N^{3})$.
-Nevertheless, existing variational GPSSMs still face significant computational and modeling challenges, which are listed below.
-  • Modeling multi-dimensional state transition function using multiple separate GP results in a computational complexity of $\mathcal{O}(d_{x}M^{3})$ in GPSSMs, where the complexity grows linearly with the latent state dimension $d_{x}$.
-  • The variational approximation of the GP posterior introduces a number of variational parameters that scale quadratically with the state dimensionality $d_{x}$.
-Specifically, the variational parameters in multi-dimensional GPSSMs include inducing inputs $\{\mathbf{Z}_{d}\}_{d=1}^{d_{x}}\in\mathbb{R}^{d_{x}\times(M\times d_{x})}$, variational means $\{\mathbf{m}_{d}\}_{d=1}^{d_{x}}\in\mathbb{R}^{d_{x}\times M}$, and covariance matrices $\{\mathbf{S}_{d}\}_{d=1}^{d_{x}}\in\mathbb{R}^{d_{x}\times(M\times M)}$ for the $d_{x}$-dimensional inducing outputs $\{\mathbf{u}_{d}\}_{d=1}^{d_{x}}$.
-As the state dimension increases, the quadratic scaling of inducing inputs can lead to computational intractability, while the scaling of the variational means and covariance matrices further compounds the computational burden [38].
-  • Existing GPSSMs commonly place a stationary GP prior over the transition function, as shown in Eq.
-(4a), to model a time-invariant transition function.
-However, this approach may fail to adequately capture non-stationary dynamics, leading to inaccuracies in scenarios where the transition function evolves over time.
-Moreover, the use of independent GPs for each dimension of the transition function overlooks dependencies between outputs.
+reducing the computational complexity from $\mathcal{O}(N^{3})$ to $\mathcal{O}(M^{3})$. Nevertheless, existing variational GPSSMs face several computational and modeling challenges, outlined below:
+  • Modeling multi-dimensional state transition function using separate GPs for each output dimension results in a computational complexity of $\mathcal{O}(d_{x}M^{3})$, scaling linearly with the latent state dimension $d_{x}$.
+  • The variational approximation introduces a number of variational parameters that scale quadratically with the state dimensionality $d_{x}$.
+Specifically, variational GPSSMs define inducing inputs $\{\mathbf{Z}_{d}\}_{d=1}^{d_{x}}\in\mathbb{R}^{d_{x}\times(M\times d_{x})}$, variational means $\{\mathbf{m}_{d}\}_{d=1}^{d_{x}}\in\mathbb{R}^{d_{x}\times M}$, and covariance matrices $\{\mathbf{S}_{d}\}_{d=1}^{d_{x}}\in\mathbb{R}^{d_{x}\times(M\times M)}$ for the $d_{x}$-dimensional inducing outputs $\{\mathbf{u}_{d}\}_{d=1}^{d_{x}}$.
+As the state dimension increases, the quadratic scaling of inducing inputs and the associated variational parameters leads to significant computational overhead [38].
+  • Most GPSSMs impose a stationary GP prior over the transition function, as shown in Eq.
+(4a), to model time-invariant dynamics.
+However, such stationary assumptions may fail to capture time-varying or context-dependent behavior [^2] Here, “stationarity” refers to the GP prior over the transition function, not the observed time series.
+A GPSSM with a stationary kernel may still model non-stationary observed data..
+Additionally, modeling each output dimension with an independent GP ignores correlations across dimensions, potentially limiting modeling capacity.
 In the next section, we detail our proposed method to resolve these issues.
 III Efficient Transformed Gaussian Process State-Space Models
-We first introduce the proposed efficient and flexible modeling framework for high-dimensional state spaces in Section III-A.
-Section III-B then presents the corresponding Bayesian variational inference algorithm.
-Finally, the details of EnKF-aided variational lower bound evaluation are provided in Section III-C.
+We first present our proposed efficient and flexible modeling framework for high-dimensional state spaces in Section III-A.
+Next, Section III-B introduces the corresponding EnKF-aided variational inference algorithm.
+Finally, Section III-C details the evaluation of the EnKF-aided variational lower bound.
 III-A Efficient Modeling for High-Dimensional State Space
@@
 $f_{1}(\cdot)$ effectively fits the data, while $f_{2}(\cdot)$ falls short.
-(2(b)) Independent GPs ($\tilde{f}_{1}(\cdot)$, $\tilde{f}_{2}(\cdot)$) vs our proposed ETGP (${f}_{1}(\cdot)$, ${f}_{2}(\cdot)$) from Fig.
+(2(b)) Independent GPs ($\tilde{f}_{1}(\cdot)$, $\tilde{f}_{2}(\cdot)$) vs our proposed ETGP (${f}_{1}(\cdot)$, ${f}_{2}(\cdot)$) depicted in Fig.
 3.
 Both methods capture the underlying function reasonably well, though the GP $\tilde{f}_{1}(\cdot)$ exhibits slightly worse fit and overestimated uncertainty.
-In Section II-C, we discussed the computational and modeling challenges associated with using separate GPs for each dimension of the transition function in high-dimensional state spaces.
-To address these issues, in [38], we proposed to warp a single shared GP using normalizing flows [43] to fit each dimension of the transition function, that is,
+In Section II-C, we discussed the computational and modeling challenges that arise from using separate GPs for each output dimension of the transition function in high-dimensional state spaces.
+To address these limitations, [38] proposed warping a single shared GP using normalizing flows [44] to model each output dimension of the transition function:
 $\tilde{f}(\cdot)\!\sim\!\mathcal{GP}(0,k(\cdot,\cdot)),~{}f_{d}(\cdot)\!=\!{\mathbb{G}}_{{{\bm{\theta}}}_{d}}(\tilde{f}(\cdot)),~{}d\!=\!1,\ldots,d_{x},$ (10)
-where each dimension-specific normalizing flow, ${{\mathbb{G}}}_{\bm{\theta}_{d}}(\cdot)$, is an element-wise, invertible and differentiable (bijective) function parameterized by $\bm{\theta}_{d}\in\Theta$.
-Two specific examples are provided below.
+where each dimension-specific normalizing flow, ${{\mathbb{G}}}_{\bm{\theta}_{d}}(\cdot)$, is an element-wise, bijective, and differentiable transformation parameterized by $\bm{\theta}_{d}\in\Theta$.
+Two concrete examples are presented below.
 Example 1.
@@
 More complex flows can be achieved by stacking additional layers of SAL flow and/or incorporating more advanced deep neural network-based flows, such as RealNVP [47].
-The primary advantage of this methodology lies in its elimination of multiple independent GPs by leveraging computationally efficient flow transformations, thereby enabling circumvent the $\mathcal{O}(d_{x}M^{3})$ computational complexity and quadratic parameter proliferation mentioned in Section II-C.
-Additionally, the warped GPs are dependent which could potentially alleviate the limitation of independence modeling in conventional GPSSMs.
+The main advantage of this approach lies in eliminating the need for multiple independent GPs by using computationally efficient flow transformations.
+This design bypasses the $\mathcal{O}(d_{x}M^{3})$ computational complexity and quadratic parameter growth discussed in Section II-C.
+Moreover, since all dimensions share the same underlying GP, the resulting warped functions are inherently dependent, potentially mitigating the limitations of modeling independent outputs in standard GPSSMs.
 However, the warped GP in Eq.
-(10) remains stationary and, being based on a single GP, may still lack the flexibility required to capture complex latent state dynamics, leading to suboptimal learning performance [38].
-An intuitive illustrative example is shown in Fig.
-2(a), where the learned element-wise flow $\mathbb{G}_{\bm{\theta}_{1}}$ successfully warps $\tilde{f}$ to obtain $f_{1}$, achieving accurate fitting of the underlying function’s simple pattern.
-However, when faced with more complex, non-stationary data trajectories, the warped GP, $f_{2}$, struggles to capture these patterns, resulting in huge modeling uncertainty.
+(10) remains stationary; Also, being derived from a single shared GP, it may lack the expressiveness needed to capture complex, time-varying latent dynamics—resulting in suboptimal performance in practice [38].
+An illustrative example is shown in Fig.
+2(a), where the learned element-wise flow $\mathbb{G}_{\bm{\theta}_{1}}$ successfully warps $\tilde{f}$ to fit a simple target function $f_{1}$.
+In contrast, the warped function $f_{2}$ struggles to capture the time-varying and intricate structure, resulting in high uncertainty.
 For a more comprehensive comparison, Fig.
-2(b) also presents the fitting results obtained using two independent GPs and the ETGP model introduced later in this paper.
-Both methods effectively capture the underlying functions, although the GP $\tilde{f}_{1}(\cdot)$ demonstrates a slightly inferior fit and inflated uncertainty estimates, due to its inherent stationarity.
-To address the limitations in modeling flexibility exhibited by existing methods, we draw inspiration from recent advances in adaptive parameterization [44, 47] and introduce non-stationarity and input dependence into the warped GPs.
-Specifically, we generate non-stationary processes $\{f_{d}(\cdot)\}_{d=1}^{d_{x}}$ by making the flow parameters depend on the input through a shallow neural network, denoted as $\textsc{nn}:\mathbb{R}^{d_{x}}\mapsto\Theta$ and parameterized by a weight vector $\mathbf{w}$.
-Formally, in the context of SSM, this is expressed as [^2] Note that ${\mathbf{x}}_{t}$ can always be extended to accommodate control systems incorporating a deterministic control input $\bm{c}_{t}\in\mathbb{R}^{d_{c}}$ by augmenting the latent state to $[{\mathbf{x}}_{t}^{\top},\bm{c}_{t}^{\top}]^{\top}\in\mathbb{R}^{d_{x}+d_{c}}$.
+2(b) shows the results from both two independent GPs and the ETGP model introduced later in this work.
+While both capture the underlying structure, the stationary GP $\tilde{f}_{1}(\cdot)$ yields inferior fits and inflated uncertainty estimates.
+To overcome these limitations in expressiveness, we draw inspiration from recent advances in adaptive parameterization [45, 48] and introduce non-stationarity and input dependence into the warped GPs.
+Specifically, we generate non-stationary processes $\{f_{d}(\cdot)\}_{d=1}^{d_{x}}$ by conditioning the flow parameters on the input through a shallow neural network $\textsc{nn}:\mathbb{R}^{d_{x}}\mapsto\Theta$ parameterized by weights $\mathbf{w}$.
+In the SSM context, this is formulated as [^3] Note that ${\mathbf{x}}_{t}$ can always be extended to accommodate control systems incorporating a deterministic control input $\bm{c}_{t}\in\mathbb{R}^{d_{c}}$ by augmenting the latent state to $[{\mathbf{x}}_{t}^{\top},\bm{c}_{t}^{\top}]^{\top}\in\mathbb{R}^{d_{x}+d_{c}}$.
 For brevity, however, we omit explicit reference to $\bm{c}_{t}$ in our notation throughout this paper.:
@@
 where ${{\bm{\theta}}}_{t}=\{{{\bm{\theta}}}_{t,d}\}_{d=1}^{d_{x}}$.
-In this case, the flow parameters vary with the input at each time step $t$, enabling the capture of local variation patterns and facilitating the modeling of time-varying transition functions.
+This formulation allows the flow parameters to vary with input at each time step $t$, enabling the model to capture local variations and time-varying dynamics.
 Additionally, note that the parameter amortization learning shown in Eq.
-(11) can avoid that the number of flow parameters increases linearly with time step $t$.
-Moreover, sharing neural network parameters across flow parameterizations further improves modeling efficiency.
-Figure 3: ETGP transition function in SSMs
+(11) prevents the number of flow parameters from growing with time step $t$ while also improving parameter efficiency through weight sharing across dimensions.
 We may further impose a prior $p_{\bm{\psi}}(\mathbf{w})$ over $\mathbf{w}$ with hyperparameter $\bm{\psi}$ (e.g., $\mathbf{w}\!\sim\!{\cal N}(\mathbf{w}|\bm{0},\operatorname{diag}(\bm{\psi}))$) to regularize the neural network’s parameters and mitigate overfitting issues.
@@
 2(b).
-It is noteworthy that this modeling approach, also known as the efficient transformed Gaussian process (ETGP) in the GP literature, was initially proposed to address large-scale multi-class classification problems [48].
-The efficiency stems from the use of a single GP rather than multiple GPs for multi-class settings, thereby ensuring that the computational cost of GP operations remains constant with respect to the number of classes.
-This leads to a substantial reduction in overall computations.
-In this work, we adapt it to model the high-dimensional transition function in SSMs.
-Additionally, ETGP implicitly characterizes $d_{x}$ dependent random processes [42], a dependence that has been overlooked in prior GPSSM literature.
-A more detailed insight is provided below.
+Figure 3: ETGP transition function in SSMs
+It is noteworthy that this modeling strategy, referred to as the efficient transformed Gaussian process (ETGP) in the GP literature, was originally proposed to tackle large-scale multi-class classification problems [42].
+Its efficiency arises from employing a single GP instead of one per class, ensuring that the computational cost of GP operations remains independent of the number of classes.
+This results in a substantial reduction in overall computation.
+In this work, we adapt the ETGP framework to model high-dimensional transition functions in SSMs.
+Notably, ETGP implicitly models $d_{x}$ dependent random processes [43], introducing output correlations that are often neglected in existing GPSSM formulations.
+We elaborate further on this below.
 Corollary 1.
@@
 Proof.
-This result can be built upon the Gaussianity and the linearity property of covariance, which can be found in Appendix A. ∎
-With the computationally efficient and flexible function prior defined by ETGP $\bm{f}(\cdot)\!=\!\{f_{d}(\cdot)\}_{d=1}^{d_{x}}$, we can define our proposed SSM, termed ETGPSSM, compactly expressed as follows:
+This result follows from the Gaussianity of the underlying process and the linearity of covariance, as detailed in Appendix C. ∎
+With the computationally efficient and flexible function prior defined by ETGP $\bm{f}(\cdot)\!=\!\{f_{d}(\cdot)\}_{d=1}^{d_{x}}$, we define our proposed SSM, termed ETGPSSM, as follows:
 $\bm{f}(\cdot)\sim\mathcal{ETGP}(\tilde{f}(\cdot),p_{\bm{\psi}}(\mathbf{w})),$ (16a)
 $\mathbf{x}_{0}\sim p({\mathbf{x}}_{0}),$ (16b)
-${\mathbf{f}}_{t}=\bm{f}({\mathbf{x}}_{t-1}),$ (16c)
+${\mathbf{f}}_{t}\triangleq\bm{f}({\mathbf{x}}_{t-1}),$ (16c)
 $\mathbf{x}_{t}|{\mathbf{f}}_{t}\sim\mathcal{N}({\mathbf{x}}_{t}|{\mathbf{f}}_{t},\mathbf{Q}),$ (16d)
 $\mathbf{y}_{t}|\mathbf{x}_{t}\sim{\cal N}({\mathbf{y}}_{t}|\bm{C}\mathbf{x}_{t},\mathbf{R}),$ (16e)
-where the ETGP prior is characterized by the associated GP $\tilde{f}(\cdot)$, neural network weight prior $p_{\bm{\psi}}(\mathbf{w})$, and the normalizing flows, see Eq.
+where the ETGP prior is defined by the base GP $\tilde{f}(\cdot)$, neural network weight prior $p_{\bm{\psi}}(\mathbf{w})$, and the normalizing flows, see Eq.
 (12).
-In this paper, unless otherwise specified, we focus on the linear flow shown in Example 3 for simplicity and illustrative purposes.
-Additionally, the linear flow requires learning only $2$ parameters per dimension, making it more efficient compared to other types of normalizing flows.
-We also assume that the prior distribution of the initial state, $p({\mathbf{x}}_{0})$, is Gaussian and known for simplicity.
+Unless stated otherwise, in this work we focus on the linear flow described in Example 3 for simplicity and illustration.
+This variant requires learning only two parameters per dimension, offering a more efficient alternative to more expressive but costly flow parameterizations.
+For simplicity, we also assume that the initial state prior $p({\mathbf{x}}_{0})$ is Gaussian and known.
 Remark 1.
@@
 The primary objective in ETGPSSM, as in GPSSM, is the joint estimation of model hyperparameters $\bm{\vartheta}\!=\!\{{{\bm{\theta}}}_{gp},\bm{\psi},\mathbf{Q},\mathbf{R}\}$ and inference of both the ETGP and latent states.
-More specifically, given a sequence of observations ${\mathbf{y}}_{1:T}=\{{\mathbf{y}}_{t}\}_{t=1}^{T}$, with $T\in\mathbb{Z}_{+}$, the task is to estimate $\bm{\vartheta}$ and compute the posterior distributions $p(\bm{f}(\cdot)|{\mathbf{y}}_{1:T})$ for the ETGP and $p({\mathbf{x}}_{0:T}|{\mathbf{y}}_{1:T})$ for the latent states, where ${\mathbf{x}}_{0:T}=\{{\mathbf{x}}_{t}\}_{t=0}^{T}$.
-However, the marginal likelihood $p(\mathbf{y}_{1:T}|\bm{\vartheta})$ is generally intractable, rendering both hyperparameter estimation via model evidence maximization and posterior inference computationally infeasible.
-Furthermore, unlike standard GPs that admit closed-form expressions for the prior and posterior distributions (see Eqs.
-(2) and (3)), the ETGP constitutes an implicit stochastic process.
-Consequently, the marginal prior [^3] Obtaining $p(\bm{f}(\cdot))$ by marginalizing over the GP and $\mathbf{w}$ is generally intractable.
-and posterior distributions are analytically intractable [42], posing significant challenges for inference.
-The next subsections introduce a variational inference framework that enables joint learning of model hyperparameters and Bayesian inference of the ETGP and latent states.
+Given a sequence of observations ${\mathbf{y}}_{1:T}=\{{\mathbf{y}}_{t}\}_{t=1}^{T}$, with $T\in\mathbb{Z}_{+}$, the task is to estimate $\bm{\vartheta}$ and compute the posterior distributions $p(\bm{f}|{\mathbf{y}}_{1:T})$ for the ETGP and $p({\mathbf{x}}_{0:T}|{\mathbf{y}}_{1:T})$ for the latent states, where ${\mathbf{x}}_{0:T}=\{{\mathbf{x}}_{t}\}_{t=0}^{T}$.
+However, the marginal likelihood $p(\mathbf{y}_{1:T}|\bm{\vartheta})$ is generally intractable, making both model evidence maximization and posterior inference computationally challenging.
+Furthermore, unlike standard GPs, which admit closed-form expressions for their prior and posterior (see Eqs.
+(2) and (3)), the ETGP defines an implicit stochastic process.
+As a result, both the marginal prior [^4] Obtaining $p(\bm{f})$ by marginalizing over the GP and $\mathbf{w}$ is generally intractable.
+and posterior distributions are analytically intractable [43], presenting significant challenges for inference.
+In the following subsections, we introduce a variational inference framework that enables joint learning of model hyperparameters and approximate Bayesian inference over the ETGP and latent states.
 III-B Approximate Bayesian Inference
-To address the intractability of the marginal distribution $p({\mathbf{y}}_{1:T}|\bm{\vartheta})$, we resort to variational inference, which involves maximizing an evidence lower bound (ELBO) on the logarithm of the marginal likelihood [49].
-The ELBO, denoted as $\mathcal{L}$, can be generally constructed as follows:
-$\mathcal{L}=\mathbb{E}_{q({\mathbf{x}}_{0:T},\bm{f}(\cdot))}\left[\log\frac{p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T},\bm{f}(\cdot))}{q({\mathbf{x}}_{0:T},\bm{f}(\cdot))}\right],$ (17)
-so that the difference between $\log p({\mathbf{y}}_{1:T}|\bm{\vartheta})$ and $\mathcal{L}$ equals the Kullback-Leibler (KL) divergence between the introduced variational approximation $q({\mathbf{x}}_{0:T},\bm{f}(\cdot))$ and the true posterior $p({\mathbf{x}}_{0:T},\bm{f}(\cdot)|{\mathbf{y}}_{1:T})$.
-We then can jointly maximize $\mathcal{L}$ with respect to $\bm{\vartheta}$ and the variational distribution $q({\mathbf{x}}_{0:T},\bm{f}(\cdot))$ to obtain the model hyperparameter estimates and the approximation of the posterior distribution $p({\mathbf{x}}_{0:T},\bm{f}(\cdot)|{\mathbf{y}}_{1:T})$.
-However, as noted earlier, the absence of closed-form expressions for the ETGP’s prior and posterior distributions prevents direct expansion of the ELBO in Eq.
-(17).
-To address this challenge, rather than operating in the function space of the ETGP, we instead follow its generative construction—illustrated in Fig.
+To address the intractability of the marginal distribution $p({\mathbf{y}}_{1:T}|\bm{\vartheta})$, we employ variational inference.
+This approach maximizes a tractable evidence lower bound (ELBO) on the log-marginal likelihood, which is equivalent to minimizing the Kullback–Leibler (KL) divergence between the variational approximation and the true posterior [49].
+Constructing the ELBO requires specifying both the model joint distribution and the variational distribution.
+Following the existing GPSSM framework and building upon Eq. (16), we factorize the joint distribution as:
+$\!\!\!p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T},\bm{f})$ $\!=\!p({\mathbf{y}}_{1:T},{\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})p(\bm{f})p({\mathbf{x}}_{0})$ (17)
+$\!=\!\prod_{t=1}^{T}p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})p({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1},\bm{f})p(\bm{f})p({\mathbf{x}}_{0}),$
+where $p({\mathbf{x}}_{0})={\cal N}({\mathbf{x}}_{0}|\bm{0},\bm{I})$ and $p(\bm{f})$ denotes the ETGP prior (under slight abuse of notation).
+A common variational family in the GPSSM literature takes the form:
+$q({\mathbf{x}}_{0:T},\bm{f})=q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})q(\bm{f})q({\mathbf{x}}_{0}),$ (18)
+which yields the ELBO, denoted by $\mathcal{L}_{2}$:
+$\mathcal{L}_{2}=\mathbb{E}_{q({\mathbf{x}}_{0:T},\bm{f})}\left[\log\frac{p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T},\bm{f})}{q({\mathbf{x}}_{0:T},\bm{f})}\right].$ (19)
+While general, this formulation requires explicit parameterization of the variational distribution $q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})$, which can introduce a large number of parameters and increase optimization complexity [37, 31].
+To improve learning efficiency, in this work, we instead adopt an alternative approach based on marginalizing out ${\mathbf{x}}_{1:T}$ and infer latent states using an EnKF, avoiding the need to explicitly model this variational term.
+Specifically, marginalizing over ${\mathbf{x}}_{1:T}$ leads to the following joint distribution:
+$\!\!\!p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0},\bm{f})$ $\!=\!p({\mathbf{x}}_{0})\,p(\bm{f})\int p({\mathbf{y}}_{1:T},{\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})\,\mathrm{d}{\mathbf{x}}_{1:T}$
+$\!=\!p({\mathbf{x}}_{0})\,p(\bm{f})\,p({\mathbf{y}}_{1:T}|{\mathbf{x}}_{0},\bm{f}),$ (20)
+where $p({\mathbf{y}}_{1:T}|{\mathbf{x}}_{0},\bm{f})$ will be approximated sequentially using an EnKF, as detailed in Section III-C.
+Since only ${\mathbf{x}}_{0}$ and $\bm{f}$ remain latent, we adopt the factorized variational distribution:
+$q({\mathbf{x}}_{0},\bm{f})=q({\mathbf{x}}_{0})\,q(\bm{f}).$ (21)
+This leads to the marginalized ELBO:
+$\mathcal{L}=\mathbb{E}_{q({\mathbf{x}}_{0},\bm{f})}\left[\log\frac{p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0},\bm{f})}{q({\mathbf{x}}_{0})\,q(\bm{f})}\right].$ (22)
+Compared to $\mathcal{L}_{2}$, the alternative bound $\mathcal{L}$ avoids direct parameterization of $q({\mathbf{x}}_{1:T}\mid\bm{f},{\mathbf{x}}_{0})$, helping reduce optimization overhead and yielding a tighter variational lower bound (see Proposition 1).
+Proposition 1.
+The bound $\mathcal{L}$ in Eq.
+(22) satisfies $\mathcal{L}\geq\mathcal{L}_{2}$.
+Equality holds if and only if the variational approximation $q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})$ matches the true posterior $p({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0},{\mathbf{y}}_{1:T})$.
+Proof.
+Proof can be found in Appendix A
+Theoretically, this tighter lower bound can improve inference accuracy and convergence [49].
+We now proceed to derive the ELBO $\mathcal{L}$.
+As noted earlier, the lack of closed-form expressions for the ETGP’s prior and posterior prevents direct computation of $\mathcal{L}$ in function space.
+To overcome this, we leverage the generative construction of the ETGP—illustrated in Fig.
 3—and infer the posterior distributions over $\tilde{f}$ and $\mathbf{w}$, which collectively characterize the ETGP posterior.
-Specifically, by explicitly representing $\tilde{f}$ and $\mathbf{w}$, the joint distribution of the proposed ETGPSSM is given by:
-$p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T},\mathbf{w},\tilde{f})=p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T}|\mathbf{w},\tilde{f})p_{\bm{\psi}}(\mathbf{w})p(\tilde{f}),$ (18)
-where $p(\tilde{f})$ represents the GP prior with a bit of notation abuse [31] and the prior of $\mathbf{w}$ is assumed to be Gaussian, $p_{\bm{\psi}}(\mathbf{w})={\cal N}(\mathbf{w}|\bm{0},\operatorname{diag}(\bm{\psi}))$, with diagonal variance.
-The term $p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T}|\mathbf{w},\tilde{f})$ is factorized as:
-$p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T}|\mathbf{w},\tilde{f})\!=\!p({\mathbf{x}}_{0})\!\prod_{t=1}^{T}\underbracket{p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})}_{\text{likelihood}}\underbracket{p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})}_{\text{transition}},$ (19)
-where the transition term is specified as:
-$p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})$ $=\int p({\mathbf{x}}_{t}|{\mathbf{f}}_{t})p({\mathbf{f}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})\mathrm{d}{\mathbf{f}}_{t}$
-$={\cal N}\left({\mathbf{x}}_{t}\mid\bm{\alpha}_{t}\cdot\tilde{f}({\mathbf{x}}_{t-1})+\bm{\beta}_{t},\ \mathbf{Q}\right),$ (20)
-with $p({\mathbf{f}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})=\delta\left({\mathbf{f}}_{t}-(\bm{\alpha}_{t}\cdot\tilde{f}({\mathbf{x}}_{t-1})+\bm{\beta}_{t})\right)$, and $\{\bm{\alpha}_{t},\bm{\beta}_{t}\}$ being the outputs of the neural network parameterized by $\mathbf{w}$.
-To infer the latent variables $\{{\mathbf{x}}_{0:T},\mathbf{w}\}$ and the GP $\tilde{f}$, we employ the following variational distribution:
-$q({\mathbf{x}}_{0:T},\tilde{f},\mathbf{w})\!=\!q(\tilde{f}),q(\mathbf{w}),q({\mathbf{x}}_{0})\prod_{t=1}^{T}q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1}),$ (21)
-which mirrors the generative structure of the model in Eq.
-(18) and assumes independence between $\tilde{f}$ and $\mathbf{w}$.
-Here the variational distributions of ${\mathbf{x}}_{0}$ and $\mathbf{w}$ are defined as follows:
+Specifically, by explicitly representing $\tilde{f}$ and $\mathbf{w}$, the marginalized joint distribution becomes:
+$p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0},\mathbf{w},\tilde{f})=p({\mathbf{y}}_{1:T}|{\mathbf{x}}_{0},\mathbf{w},\tilde{f})p({\mathbf{x}}_{0})p_{\bm{\psi}}(\mathbf{w})p(\tilde{f}),$ (23)
+where $p(\tilde{f})$ represents the GP prior and $p_{\bm{\psi}}(\mathbf{w})={\cal N}(\mathbf{w}|\bm{0},\operatorname{diag}(\bm{\psi}))$ is a Gaussian prior over the neural network weights, assuming a Bayesian neural network with diagonal covariance.
+Similarly, instead of using Eq.
+(21), we employ the following variational distribution:
+$q({\mathbf{x}}_{0},\mathbf{w},\tilde{f})=q(\tilde{f})q(\mathbf{w})q({\mathbf{x}}_{0}),$ (24)
+which assumes independence between $\tilde{f}$, $\mathbf{w}$ and ${\mathbf{x}}_{0}$. Here the variational distributions of ${\mathbf{x}}_{0}$ and $\mathbf{w}$ are defined as follows:
 […]
@@
 (8), which helps to reduce the computational complexity from $\mathcal{O}(T^{3})$ to $\mathcal{O}(M^{3})$.
-For $q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})$ in Eq.
-(21), rather than assuming a parameterized form—which would introduce additional variational parameters and hence increase optimization complexity [31]—we draw inspiration from recent works [37, 50] and incorporate the EnKF into the variational learning framework to infer the latent states (see the next subsection for details).
-Specifically, without explicitly defining the variational distribution, we adopt the following approximation:
-Assumption 1.
-$q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})\approx p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1},{\mathbf{y}}_{1:t}),$ (23)
-which assumes that the variational distribution of ${\mathbf{x}}_{t}$ can be approximated by the corresponding filtering distribution.
-Additionally, we assume that the following filtering distribution is approximately equal to its one-step backward smoothing counterpart, i.e.,
-Assumption 2.
-$p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})\approx p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t}),$ (24)
-The two assumptions are reasonable and detailed justification can be found in Appendix B.
-With these two assumptions, we can significantly streamline the evaluation of the ELBO, which is summarized in the following proposition.
+With Eqs. (23) and (24), we can derive our ELBO, $\mathcal{L}$, which is summarized in the following proposition.
 Proposition 2.
-Under the two approximations, and all the assumed variational distributions, together with the model joint distribution in Eq. (18), the ELBO can be reformulated as follows:
-$\mathcal{L}\approx$ $\mathbb{E}_{q(\mathbf{w},\tilde{f})}\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))$ (25)
-$-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0})),$
-where the first term (log-likelihood) can be analytically evaluated using the EnKF (detailed in Section III-C).
+Under all the assumed variational distributions, together with the model joint distribution in Eq. (23), the ELBO can be reformulated as follows:
+$\mathcal{L}=$ $\mathbb{E}_{q(\mathbf{w},\tilde{f})}\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))$
+$-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0})),$ (26)
+where the first terms (log-likelihood) can be analytically evaluated using the EnKF (detailed in Section III-C).
 The three KL divergence terms can also be computed in closed form, due to the Gaussian nature of the prior and variational distributions [49].
@@
 The proof can be found in Appendix B. ∎
+Remark 2.
+The objective in Eq.
+(26) can also be derived from $\mathcal{L}_{2}$ under two specific assumptions (see Supplement A), akin to [37].
+However, our derivation in this work is generic for arbitrary random processes and is assumption-free, offering greater simplicity, conceptual clarity, and broader applicability.
+For further details on the relationship between $\mathcal{L}$ and $\mathcal{L}_{2}$, we refer to Supplement  A.
 III-C EnKF-aided ELBO Evaluation
-We now demonstrate how to evaluate $p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})$ in Eq.
-(25) using the EnKF.
-The general idea can be derived from Bayesian filtering.
-We first notice that
-$p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})$ $=\int p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})\mathrm{d}{\mathbf{x}}_{t},$ (26)
-where the prediction distribution, $p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})=$
+This section describes how to evaluate $p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})$ in Eq. (26) using the EnKF. First, note that
+$p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})$ $=\int\underbracket{p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})}_{\text{likelihood}}\underbracket{p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})}_{\text{prediction}}\mathrm{d}{\mathbf{x}}_{t},$ (27)
+where the prediction distribution $p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})$ reads:
 $\int\underbracket{p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})}_{\text{transition}}\underbracket{p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})}_{\text{filtering}}\mathrm{d}{\mathbf{x}}_{t-1}.$ (28)
-Consequently, we employ the EnKF, a widely used method for handling nonlinearities and high-dimensional state spaces [51], to obtain the filtering distribution $p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t})$ at each time step $t$.
-Further details are provided below.
+The transition term can be expressed as
+$p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})$ $=\int p({\mathbf{x}}_{t}|{\mathbf{f}}_{t})p({\mathbf{f}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})\mathrm{d}{\mathbf{f}}_{t}$
+$={\cal N}\left({\mathbf{x}}_{t}\mid\bm{\alpha}_{t}\cdot\tilde{f}({\mathbf{x}}_{t-1})+\bm{\beta}_{t},\ \mathbf{Q}\right),$ (29)
+with $p({\mathbf{f}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})=\delta\left({\mathbf{f}}_{t}-(\bm{\alpha}_{t}\cdot\tilde{f}({\mathbf{x}}_{t-1})+\bm{\beta}_{t})\right)$, and $\{\bm{\alpha}_{t},\bm{\beta}_{t}\}$ being the outputs of the neural network parameterized by $\mathbf{w}$.
+Consequently, we employ the EnKF, a widely used method for handling nonlinearities and high-dimensional state spaces [50], to estimate the filtering distribution $p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t})$ at each time step $t$.
+The details are as follows.
 III-C1 Prediction
-Given the ensemble of states $\{{\mathbf{x}}_{t-1}^{(i)}\}_{i=1}^{N}$ from the posterior distribution $p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})$ at time $t-1$, and conditioning on $\mathbf{w}\sim q(\mathbf{w})$ and $\tilde{f}\sim q(\tilde{f})$, we utilize the state transition in Eq.
-(20) to perform the prediction step.
-This generates $N$ predicted samples $\{\bar{{\mathbf{x}}}_{t}^{(i)}\}_{i=1}^{N}$ as follows:
-$\bar{{\mathbf{x}}}_{t}^{(i)}\sim{\cal N}\left({\mathbf{x}}_{t}\mid\bm{\alpha}_{t}\cdot\tilde{f}({\mathbf{x}}_{t-1}^{(i)})+\bm{\beta}_{t},\ \mathbf{Q}\right),i=1,2,\ldots,N.$ (28)
-Therefore, the prediction distribution in Eq. (27) is approximated as:
+Given the ensemble of states $\{{\mathbf{x}}_{t-1}^{(i)}\}_{i=1}^{N}$ from the posterior distribution $p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})$ at time $t-1$, and conditioning on $\mathbf{w}\sim q(\mathbf{w})$ and $\tilde{f}\sim q(\tilde{f})$, we use the state transition in Eq.
+(29) to generate $N$ predictive samples:
+$\!\!\bar{{\mathbf{x}}}_{t}^{(i)}\!\sim\!{\cal N}\left({\mathbf{x}}_{t}\mid\bm{\alpha}_{t}\cdot\tilde{f}({\mathbf{x}}_{t-1}^{(i)})+\bm{\beta}_{t},\ \mathbf{Q}\right)$ (30)
+for $1\leq i\leq N$. The prediction distribution in Eq. (28) is then approximated by the predictive ensemble:
 $p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})\approx{\cal N}({\mathbf{x}}_{t}\mid\bar{\mathbf{m}}_{t},\ \bar{\mathbf{P}}_{t}),$ (31)
@@
 III-C2 Update
-Observations $\mathbf{y}_{t}$ are then assimilated into the predictive ensemble to update the state estimate. Specifically, the predictive ensemble is updated using:
+Observations $\mathbf{y}_{t}$ is incorporated to update the predictive ensemble. Each sample is updated via:
 $\ \ \ \ \ \ \text{{update}:}\ \ \ \ \mathbf{x}_{t}^{(i)}=\bar{\mathbf{x}}_{t}^{(i)}+\mathbf{G}_{t}(\mathbf{y}_{t}+\mathbf{e}_{t}^{(i)}-\bm{C}\bar{\mathbf{x}}_{t}^{(i)}),$ (33a)
@@
 $\mathbf{G}_{t}=\bar{\mathbf{P}}_{t}\bm{C}^{\top}(\bm{C}\bar{\mathbf{P}}_{t}\bm{C}^{\top}+\mathbf{R})^{-1}.$ (34)
-Thus we can obtain the set of $N$ updated samples $\{{\mathbf{x}}_{t}^{(i)}\}_{i=1}^{N}$ from the filtering distribution at time step $t$.
-These updated samples allow us to recursively compute the samples and prediction distribution for subsequent time steps.
+This yields the updated ensemble $\{{\mathbf{x}}_{t}^{(i)}\}_{i=1}^{N}$ which is used to recursively propagate forward through time.
 Remark 3.
-The prediction and update steps are inherently differentiable, as the sampled latent states become differentiable with respect to the state transition function $\bm{f}(\cdot)$ and the noise covariances $\mathbf{Q}$ and $\mathbf{R}$.
-This property allows for a straightforward gradient-based optimization (e.g.
-Adam [52] in Algorithm 1), facilitating a practical learning framework for parameter estimation and model tuning [50, 37].
+The prediction and update steps are fully differentiable, as the sampled latent states become differentiable with respect to the state transition function $\bm{f}(\cdot)$ and the noise covariances $\mathbf{Q}$ and $\mathbf{R}$.
+This allows for end-to-end gradient-based optimization (e.g.using Adam [51], as described in Algorithm 1), making the framework practical for joint learning and inference [52, 37].
+Based on the outlined prediction and update steps, the likelihood $p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})$ in Eq. (27) can be easily evaluated. At each time step $t$, we have:
+$p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})\approx{\cal N}({\mathbf{y}}_{t}\mid\bm{C}\bar{\mathbf{m}}_{t},\ \bm{C}\bar{\mathbf{P}}_{t}\bm{C}^{\top}),$ (35)
+due to the Gaussian prediction distribution, see Eq. (31), and the linear emission model, see Eq. (16e).
+We now proceed to evaluate the variational lower bound $\mathcal{L}$, as defined in Eq.
+(26).
+At each iteration, we apply the reparameterization trick (see e.g.
+Eq.
+(33b)) to sample from the Gaussian variational distributions $q(\mathbf{w})$ and $q(\tilde{f})$, enabling unbiased Monte Carlo estimates of the expected log-likelihood term $\mathbb{E}_{q(\mathbf{w},\tilde{f})}[\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})]$.
+Thanks to the reparameterization trick [53], the ELBO $\mathcal{L}$ remains differentiable with respect to both the model parameters $\bm{\vartheta}=\{\bm{\theta}_{gp},\bm{\psi},\mathbf{Q},\mathbf{R}\}$ and the variational parameters ${\bm{\zeta}}=\{\mathbf{m}_{0},\mathbf{L}_{0},\mathbf{m}_{\mathrm{w}},\bm{\sigma}_{\mathrm{w}}^{2},\mathbf{m},\mathbf{S},\mathbf{Z}\}$.
+We leverage modern automatic differentiation tools such as PyTorch to compute gradients via backpropagation through time (BPTT), and apply gradient-based optimization methods (e.g., Adam [51]) to maximize $\mathcal{L}$ [54].
+The complete procedure for training ETGPSSM with EnKF-based latent state inference is summarized in Algorithm 1.
 Algorithm 1 ETGPSSM
@@
 5:         Get $\tilde{f}$ from $q(\tilde{f})$ using Eq. (8) with ${\mathbf{x}}_{*}={\mathbf{x}}_{t-1}$
-6:         Get prediction samples using Eq. (28)
-7:         Get empirical moments $\bar{\mathbf{m}}_{t},\bar{\mathbf{P}}_{t}$ using Eq. (30)
-8:         Get Kalman gain: $\bar{\mathbf{G}}_{t}\!=\!\bar{\mathbf{P}}_{t}\bm{C}^{\top}(\bm{C}\bar{\mathbf{P}}_{t}\bm{C}^{\top}\!\!+\!\mathbf{R})^{-1}$
-9:         Get updated samples using Eq. (31)
-10:         Evaluate the log-likelihood using Eq. (33), and
-$L_{\ell}=L_{\ell}+\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})$
-11:     end for
-12:     Evaluate $\mathcal{L}$ based on $L_{\ell}$ and Eq. (25)
-13:     Maximize $\mathcal{L}$ and update $\bm{\vartheta}$, ${\bm{\zeta}}$ using Adam [52]
-14:end while
-15:Output: EnKF particles $\{{\mathbf{x}}_{0:T}^{(i)}\}_{i=1}^{N}$, model parameters $\bm{\vartheta}$, and variational parameters ${\bm{\zeta}}$.
-Based on the outlined prediction and update steps, the likelihood $p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})$ in Eq.
-(26) can be easily evaluated.
-Specifically, at each time step $t$, we have:
-$p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})\approx{\cal N}({\mathbf{y}}_{t}\mid\bm{C}\bar{\mathbf{m}}_{t},\ \bm{C}\bar{\mathbf{P}}_{t}\bm{C}^{\top}),$ (33)
-due to the Gaussian prediction distribution, see Eq. (29), and the linear emission model, see Eq. (16e).
-Now, we can evaluate our variational lower bound, $\mathcal{L}$, as defined in Eq.
-(25).
-At each step, we first utilize the reparameterization trick (see e.g.
-Eq.
-(31b)) to sample from the Gaussian variational distributions $q(\mathbf{w})$ and $q(\tilde{f})$.
-This allows us to obtain an unbiased estimate of the expected log-likelihood, $\mathbb{E}_{q(\mathbf{w},\tilde{f})}[\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})]$ numerically.
-Due to the reparameterization trick [53], $\mathcal{L}$ is differentiable with respect to the model parameters $\bm{\vartheta}=\{\bm{\theta}_{gp},\bm{\psi},\mathbf{Q},\mathbf{R}\}$ and the variational parameters ${\bm{\zeta}}=\{\mathbf{m}_{0},\mathbf{L}_{0},\mathbf{m}_{\mathrm{w}},\bm{\sigma}_{\mathrm{w}}^{2},\mathbf{m},\mathbf{S},{\vec{\mathbf{z}}}\}$.
-Consequently, we leverage modern differentiation tools, such as PyTorch, to automatically compute the gradients through backpropagation through time (BPTT) and apply gradient-based optimization methods (e.g., Adam) to maximize $\mathcal{L}$ [54, 52].
-The detailed routine for implementing our EnKF-aided ETGPSSM is summarized in Algorithm 1.
+6:    EnKF
+7:         Get prediction samples using Eq. (30)
+8:         Get empirical moments $\bar{\mathbf{m}}_{t},\bar{\mathbf{P}}_{t}$ using Eq. (32)
+9:         Get Kalman gain: $\bar{\mathbf{G}}_{t}\!=\!\bar{\mathbf{P}}_{t}\bm{C}^{\top}(\bm{C}\bar{\mathbf{P}}_{t}\bm{C}^{\top}\!\!+\!\mathbf{R})^{-1}$
+10:         Get updated samples using Eq. (33)
+11:         Evaluate the log-likelihood using Eq. (35), and
+$L_{\ell}\leftarrow L_{\ell}+\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})\vspace{-.1in}$
+12:
+13:     end for
+14:     Evaluate $\mathcal{L}$ based on $L_{\ell}$ and Eq. (26)
+15:     Maximize $\mathcal{L}$ and update $\bm{\vartheta}$, ${\bm{\zeta}}$ using Adam [51]
+16:end while
+17:Output: EnKF particles $\{{\mathbf{x}}_{0:T}^{(i)}\}_{i=1}^{N}$, model parameters $\bm{\vartheta}$, and variational parameters ${\bm{\zeta}}$.
 Remark 4 (Efficiency).
-The proposed ETGPSSM in Algorithm 1 improves upon existing GPSSMs in two primary aspects of efficiency.
-First, its computational complexity is reduced from $\mathcal{O}(d_{x}M^{3})$ to $\mathcal{O}(M^{3})$ by utilizing a single GP—while the shallow neural network adds minimal computational overhead.
-Second, the use of a single GP and shallow neural network prevents the quadratic growth of variational parameters, as discussed in Section II-C.
-Moreover, by avoiding explicit parameterization of the latent states and instead leveraging the EnKF, the proposed method further reduces parametric complexity.
-Together, these design choices contribute to both parameterization and computational efficiency.
-Further empirical validation of these improvements is provided in Section IV-B.
-In the next section, we empirically evaluate the performance of ETGPSSM across various datasets.
-IV Experiments
-We first demonstrate the advantages of ETGPSSM in learning non-stationary dynamical systems in Section IV-A.
-Next, in Section IV-B, we highlight the computational and parametric efficiency and filtering performance of ETGPSSM in chaotic high-dimensional dynamical systems.
-Finally, Section IV-C showcases the superiors prediction performance of ETGPSSM using multiple real-world time series datasets.
-More experimental details can be found in the accompanying source code, which is publicly available online.
-[^4] https://github.com/zhidilin/gpssmProj/tree/main/high_dim_GPSSM
-IV-A Non-Stationary Dynamical System Learning
+The proposed ETGPSSM in Algorithm 1 improves upon existing GPSSMs in two key aspects of efficiency.
+First, it reduces the computational complexity from $\mathcal{O}(d_{x}M^{3})$ to $\mathcal{O}(M^{3})$ by relying on a single GP, while the added cost of a shallow neural network remains negligible.
+Second, by sharing the GP and flow network across dimensions, it avoids the quadratic scaling of variational parameters associated with independent GPs, as discussed in Section II-C.
+Furthermore, by replacing explicit variational parameterization of the latent states with EnKF-based inference, the method significantly reduces parametric overhead.
+Together, these design choices lead to substantial gains in both computational and modeling efficiency.
+Empirical results validating these advantages are presented in Section IV-B.
+In the next section, we empirically evaluate the performance of ETGPSSM across a range of dynamical system modeling tasks.
             Data               AD-EnKF [52] AD-EnKF (bnn)          EnVI [37] ETGPSSM (dnn)      ETGPSSM (bnn)
@@
 The shaded blue region depicts the confidence interval.
-This subsection demonstrates the superior modeling capabilities of ETGPSSM over existing GPSSM and neural network-based methods by evaluating the learning performance on the following underlying dynamical system:
+IV Experiments
+We begin by evaluating ETGPSSM’s ability to model non-stationary dynamical systems in Section IV-A.
+We then assess its computational and parametric efficiency, as well as filtering accuracy, in chaotic high-dimensional settings in Section IV-B.
+Lastly, Section IV-C highlights the model’s predictive performance across a range of real-world time series datasets.
+Unless otherwise stated, all GPSSMs use the squared exponential (SE) kernel [9], with $20$ inducing points initialized at random.
+The shallow neural network used in ETGPSSM is a fully connected three-layer architecture with ReLU activations and hidden dimensions $128$ and $64$, mapping from $\mathbb{R}^{d_{x}}$ to the flow parameter space $\Theta\!\subseteq\!\mathbb{R}^{2d_{x}}$.
+Additional implementation details are available in the publicly released code.
+[^5] https://github.com/zhidilin/gpssmProj/tree/main/high_dim_GPSSM
+IV-A Non-Stationary Dynamical System Learning
+This subsection evaluates the modeling capabilities of ETGPSSM on a non-stationary dynamical system, comparing its performance to state-of-the-art GPSSM and neural network-based methods.
+The underlying system is defined as:
 ${\mathbf{x}}_{t+1}=f({\mathbf{x}}_{t})+\mathbf{v}_{t},$ $\mathbf{v}_{t}\sim{\cal N}(0,\sigma_{\mathrm{Q}}^{2}),$ (36a)
@@
 $f({\mathbf{x}})\!=\!\underbrace{\left[0.8+({\mathbf{x}}+0.2)\left(1-\frac{5}{1+\exp(-2{\mathbf{x}})}\right)\right]}_{\text{``kink'' function}}s({\mathbf{x}})-o({\mathbf{x}})$ (37)
-with a slope function:
+with a slope modulation:
 $s({\mathbf{x}})=\begin{cases}1-0.5\exp(-0.5{\mathbf{x}}),&{\mathbf{x}}>0,\\ 1,&{\mathbf{x}}\leq 0,\end{cases}$ (38)
-and an oscillatory function:
+and an oscillatory component:
 $o({\mathbf{x}})=\begin{cases}0.5\sin(8{\mathbf{x}}),&{\mathbf{x}}>0,\\ 0.5\sin(2{\mathbf{x}}),&{\mathbf{x}}\leq 0.\end{cases}$ (39)
-The slope and oscillatory functions create non-stationary pattern across different regions of the “kink” function (see Fig. 4).
-Using this underlying SSM, we generate training data ${\mathbf{y}}_{1:T}$ by fixing $\sigma_{\mathrm{Q}}^{2}$ at 0.05 and varying $\sigma_{\mathrm{R}}^{2}$ systematically across the set $\{0.0008,0.008,0.08,0.8\}$, resulting in four sets of $T=600$ observations each for training purposes.
-To demonstrate the advantages of ETGPSSM, we compare it with the state-of-the-art (SOTA) neural network-based AD-EnKFs [50] and GP-based EnVI [37].
-Consistent with previous works [31, 37], we retain the true emission model and allow the transition model to be learned in all methods, ensuring a fair comparison within the latent space.
+The kink function is a standard benchmark for evaluating GPSSMs [27, 31, 37].
+Here, we extend it with non-stationary slope and oscillatory terms to increase modeling complexity (see Fig.
+4).
+We generate training data ${\mathbf{y}}_{1:T}$ by fixing $\sigma_{\mathrm{Q}}^{2}$ at $0.05$ and varying $\sigma_{\mathrm{R}}^{2}$ systematically across the set $\{0.0008,0.008,0.08,0.8\}$, producing four noise regimes.
+ETGPSSM is compared to two leading baselines: the neural network-based AD-EnKF [52] and the GP-based EnVI [37].
+Consistent with previous works [31, 37], all methods share the true emission model while learning the transition function, ensuring a fair comparison.
 The system dynamics learning results across different methods are shown in Fig.
@@
 The results in Fig.
-4 demonstrate that ETGPSSMs consistently achieve superior system dynamics learning performance.
-Under small observational noise levels ($\sigma_{\mathrm{R}}^{2}=\{0.0008,0.008\}$), ETGPSSMs—particularly ETGPSSM with DNN—excel in learning the dynamics of non-stationary systems.
-In contrast, their competitor, AD-EnKFs, fails to capture the smooth left segment of the kink function and performs well only on the more zigzag right segment.
-Meanwhile, EnVI, the SOTA GPSSM, fails to capture the non-stationary behavior of the system and is unable to model the zigzag region effectively.
-These results highlight the strength of ETGPSSMs in integrating the representational flexibility of neural networks with the GPs, making them particularly well-suited for modeling non-stationary dynamics.
-When the observational noise interference becomes more significant $(\sigma_{\mathrm{R}}^{2}=\{0.08,0.8\})$, the learning performance of all methods deteriorates.
-This is due to the higher estimation errors in the latent states, which hinder accurate learning of the system dynamics.
-Nevertheless, ETGPSSMs continue to demonstrate superior learning performance.
-Notably, in this scenario, BNN-based methods, particularly ETGPSSM (bnn), exhibit superior performance and more reliable uncertainty quantification compared to their DNN-based counterparts.
-These results highlight not only the effectiveness of ETGPSSM’s non-stationary modeling capabilities but also the regularization benefits inherent in Bayesian methods, which contribute to enhanced robustness against significant observational noise.
+4 show that, across all noise levels, ETGPSSM achieves consistently superior system identification performance.
+In the low-noise regimes($\sigma_{\mathrm{R}}^{2}=\{0.0008,0.008\}$), the DNN-based ETGPSSM excels at capturing both smooth and oscillatory regions of the transition function.
+By contrast, AD-EnKF struggles with the smooth left segment and performs well only on the right, highly nonlinear region.
+EnVI, the strongest GPSSM baseline, fails to capture the non-stationarity and underfits the high-frequency component.
+Under high noise $(\sigma_{\mathrm{R}}^{2}=\{0.08,0.8\})$, performance declines across all methods due to increased latent state uncertainty.
+Nevertheless, ETGPSSM remains robust.
+Notably, in this scenario, BNN-based methods, particularly ETGPSSM (bnn), outperforms the DNN-based version, benefiting from regularization and better uncertainty quantification.
+These results emphasize the importance of modeling flexibility and demonstrate the advantage of combining neural flows with GP priors.
 IV-B High-Dimensional Latent State Estimation
-In this subsection, we focus on latent state estimation in the Lorenz-96 system, a mathematical model for studying chaotic behavior in weather and climate dynamics [55].
+In this subsection, we evaluate latent state estimation in the Lorenz-96 system, a standard benchmark for studying chaotic dynamics in weather and climate modeling [55].
 The system consists of $d_{x}\in\mathbb{Z}_{+}$ coupled ordinary differential equations (ODEs):
 […]
-Here the constant forcing parameter $F$ is set to $8$, resulting a fully chaotic dynamic [55], where small variations in initial conditions lead to vastly different trajectories.
+where $F=8$ induces fully chaotic behavior, with small differences in initial conditions leading to divergent trajectories [55].
 The dynamics model is discretized using the Euler method, and at each time step $t$, observations are generated through a linear emission model ($\bm{C}\!=\!\bm{I}$) with Gaussian additive noise, where the noise level is significant with the covariance given by $\mathbf{R}=4\bm{I}$.
-We first empirically demonstrate the efficiency of ETGPSSM in high-dimensional latent state spaces, as illustrated in Fig.
-5.
-This figure highlights both the computational efficiency and the parameter scalability of ETGPSSM in comparison to conventional GPSSMs across varying latent dimensions.
-The shallow neural network used in ETGPSSM is a three-layer fully connected architecture with ReLU activation and hidden dimensions of $128$ and $64$, which maps from $\mathbb{R}^{d_{x}}$ to the flow parameter space $\Theta\!\subseteq\!\mathbb{R}^{2d_{x}}$.
-As shown in the left panel, the number of parameters in ETGPSSM increases linearly with the latent state dimension $d_{x}$, in contrast to the quadratic parameter growth observed in GPSSMs.
-This linear scalability is achieved through the use of a single shared GP combined with normalizing flows, which drastically reduces the total number of parameters compared to GPSSMs that rely on independent GPs for each state dimension.
-Figure 5: (Left) The number of parameters in ETGPSSM grows linearly with the state dimension $d_{x}$, whereas in GPSSM, it scales quadratically.
-(Right) ETGPSSM maintains low computational costs as $d_{x}$ increases, while GPSSM exhibits linear growth.
-Figure 6: Latent state estimation performance for 100-dimensional Lorenz-96 system, with 9 dimensions uniformly sampled for visualization.
-Root mean square error (RMSE) values for different methods: 0.5468 for EnKF, 1.2028 for ETGPSSM, 1.6187 for AD-EnKF.
-RMSE between latent states and observations: 1.9855.
+We first assess the scalability of ETGPSSM in high-dimensional settings.
+Fig.
+5 compares its parameter growth and runtime to standard GPSSMs, using $100$ inducing points.
+As shown in the left panel, ETGPSSM exhibits linear growth in parameter count with increasing $d_{x}$, while GPSSM grows quadratically.
+This is enabled by ETGPSSM’s use of a shared GP combined with normalizing flows, in contrast to GPSSM’s reliance on independent GPs per dimension.
 The right panel of Fig.
@@
 This efficiency stems from the reduced computational complexity of ETGPSSM, which scales as $\mathcal{O}(M^{3})$ due to its use of a single GP, in contrast to GPSSM’s $\mathcal{O}(d_{x}M^{3})$ complexity.
-These results demonstrate that ETGPSSM achieves more scalable parameter growth with increasing state dimensionality while maintaining computational tractability in high-dimensional settings, making it a practical choice for modeling high-dimensional chaotic systems like the Lorenz-96 system.
+These results demonstrate that ETGPSSM remains computationally and parametrically tractable in high-dimensional latent spaces.
+Figure 5: (Left) The number of parameters in ETGPSSM grows linearly with the state dimension $d_{x}$, whereas in GPSSM, it scales quadratically.
+(Right) ETGPSSM maintains low computational costs as $d_{x}$ increases, while GPSSM exhibits linear growth.
+Figure 6: Latent state estimation performance for 100-dimensional Lorenz-96 system, with 9 dimensions uniformly sampled for visualization.
+Root mean square error (RMSE) values for different methods: 0.5468 for EnKF, 1.2028 for ETGPSSM, 1.6187 for AD-EnKF.
+RMSE between latent states and observations: 1.9855.
 We next fix the dimension $d_{x}\!=\!100$ in the Lorenz-96 system and generate observations of length $T\!=\!600$ for training ETGPSSM and its deep neural network-based competitor, AD-EnKF [52].
@@
 6, where the performance of EnKF is also included for comparison.
-As illustrated in Fig.
-6, ETGPSSM achieves a root mean square error (RMSE) of 1.2028, outperforming AD-EnKF, which yields an RMSE of 1.6187.
+As depicted in Fig.
+6, our ETGPSSM achieves an RMSE of $1.2028$, outperforming AD-EnKF, which yields an RMSE of $1.6187$.
+Additionally, note that the RMSE of the observations (1.9855) serves as a baseline, demonstrating that ETGPSSM significantly reduces the estimation error compared to the raw observations.
 This result highlights the superior filtering performance of ETGPSSM in high-dimensional chaotic systems.
-Additionally, the RMSE of the observations (1.9855) serves as a baseline, demonstrating that ETGPSSM significantly reduces the estimation error compared to the raw observations.
-Although EnKF yields the lowest RMSE of 0.5468, it relies on full model knowledge and does not learn system dynamics.
-In contrast, ETGPSSM simultaneously learns dynamics and infers latent states, offering a more versatile and effective solution for high-dimensional complex systems.
-TABLE I: Prediction performance in terms of RMSE of the different models on the system identification datasets.
+While EnKF achieves the lowest RMSE of $0.5468$, it assumes full knowledge of the dynamics and performs no learning.
+In contrast, ETGPSSM simultaneously learns the system and infers latent states, offering a more general and powerful solution for complex, high-dimensional systems.
+IV-C Real-World Time Series Prediction
+In this subsection, we evaluate the prediction performance of ETGPSSM on several real-world time series datasets [^7] https://homes.esat.kuleuven.be/~smc/daisy/daisydata.html commonly used for system identification (see Table I).
+In every dataset, the first half of the sequence is utilized as training data, with the remaining portion designated for testing.
+Standardization of all datasets is conducted based on the training sequence.
+Table I reports the series prediction results, wherein the RMSE is averaged over 50-step ahead forecasting.
+We compare ETGPSSM against state-of-the-art methods, including neural network-based approaches—DKF [7] and AD-EnKF [52]—as well as various GPSSMs: vGPSSM [29], CO-GPSSM [36], PRSSM [30], ODGPSSM [39], VCDT [31], and EnVI [37].
+Following previous works, the latent space dimension is set to $d_{x}=4$ for all the methods.
+TABLE I: The 50-step ahead prediction performance in terms of RMSE of different models (all with $d_{x}=4$) on system identification datasets.
 Mean and standard deviation of the prediction results are shown as an averaged result over ten different seeds.
@@
 ETGPSSM (bnn) $0.646\pm 0.095$ $0.053\pm 0.002$ $0.703\pm 0.028$ $0.154\pm 0.026$ $1.313\pm 0.048$
-IV-C Real-World Time Series Prediction
-In this subsection, we evaluate the prediction performance of ETGPSSM on several real-world time series datasets commonly used for system identification (see Table I).
-The prediction RMSE results are summarized in Table I, comparing ETGPSSM against SOTA methods, including neural network-based approaches—DKF [7] and AD-EnKF [50]—as well as various GPSSMs: vGPSSM [29], CO-GPSSM [36], PRSSM [30], ODGPSSM [39], VCDT [31], and EnVI [37].
-Following previous works, the latent space dimension is set to $d_{x}=4$ for all the methods.
-As shown in Table I, ETGPSSM demonstrates competitive performance across all datasets.
-Notably, on the Actuator dataset, ETGPSSM (dnn) and ETGPSSM (bnn) both achieve an improved RMSE of 0.646, outperforming EnVI (0.657) and significantly improving upon AD-EnKFs, which record an RMSE of 0.705.
-Similarly, on the Ball Beam and Gas Furnace datasets, ETGPSSMs attain the first and second lowest RMSEs, highlighting their ability to accurately model system dynamics.
-However, on datasets such as Drive and Dryer, other GPSSMs like PRSSM and EnVI achieve slightly superior performance.
-This can be attributed to the relatively low latent state dimension ($d_{x}\!=\!4$), where traditional GPSSMs, such as PRSSM, ODGPSSM and EnVI, remain highly competitive, as discussed in [37].
-Nonetheless, we emphasize that the proposed ETGPSSM remains highly competitive, achieving very comparable performance even in low-dimensional settings, while significantly reducing computational complexity and parameter overhead—thereby ensuring both scalability and efficiency in high-dimensional scenarios.
-In comparison to neural network-based methods like AD-EnKF, the superior performance of ETGPSSM is driven by its ability to integrate the strengths of GPs and neural networks.
-This hybrid approach allows ETGPSSM to effectively model complex, non-stationary system dynamics while providing reliable predictions—an area where pure deep neural network-based methods often struggle.
+As reported in Table I, ETGPSSM demonstrates strong overall performance.
+On the Actuator dataset, both ETGPSSM (dnn) and ETGPSSM (bnn) achieve an RMSE of $0.646$, outperforming EnVI ($0.657$) and significantly improving upon AD-EnKFs, which yield an RMSE of $0.705$.
+Similar trends are observed on the Ball Beam and Gas Furnace datasets, where ETGPSSM ranks first and second, respectively, indicating effective modeling of system dynamics.
+On the Drive and Dryer datasets, however, GPSSMs such as PRSSM and EnVI attain slightly better performance.
+This is likely due to the relatively low latent state dimension ($d_{x}=4$), where conventional GPSSMs, including PRSSM, ODGPSSM, and EnVI, remain competitive, as discussed in [37].
+Nonetheless, we highlight that although the primary contribution of the proposed ETGPSSM lies in its scalability and efficiency in high-dimensional settings, it also remains highly competitive in low-dimensional scenarios, achieving superior or very comparable performance.
+In comparison to neural network-based methods like AD-EnKF, ETGPSSM’s hybrid design leverages the representational power of deep learning while retaining the uncertainty quantification and regularization benefits of GPs.
+This synergy proves especially effective for modeling complex, non-stationary dynamics.
 This is evident from the results, where neural network-based methods exhibit higher prediction errors.
-Notably, AD-EnKF (bnn) generally outperforms other network-based approaches, corroborating the benefits of Bayesian treatments, particularly their inherent regularization, which helps mitigate overfitting and improve generalization.
-In summary, the experimental results demonstrate that ETGPSSM is a powerful and versatile tool for real-world time series prediction.
+AD-EnKF (bnn) consistently outperforms its DNN counterpart, underscoring the robustness provided by Bayesian approaches.
+Still, both lag behind ETGPSSM in predictive accuracy across most datasets.
+In summary, ETGPSSM demonstrates strong predictive performance on diverse real-world time series, matching or outperforming prior methods, while offering a scalable and efficient solution applicable to both low- and high-dimensional dynamical systems.
 V Conclusions
-In this paper, we presented a novel approach for learning and inference in GPSSMs tailored to non-stationary, high-dimensional dynamical systems.
-Our key contribution is the introduction of the ETGP prior, which integrates a shared GP with input-dependent normalizing flows modeled via neural networks.
-This ETGP enhances flexibility in capturing complex, non-stationary transitions while significantly reducing the computational and parametric complexity of traditional GPSSMs in high-dimensional settings.
-To enable scalable inference, we developed a variational framework that approximates the posterior over the ETGP by separately modeling the underlying GP and neural network parameters.
-Additionally, we leverage the EnKF to circumvent the need for explicit parameterization of state posteriors, further improving computational efficiency.
-Empirical evaluations across dynamical system learning, high-dimensional state estimation, and time-series prediction tasks demonstrate that our method outperforms existing GPSSMs and deep neural network-based SSMs, while maintaining lower computational and parametric complexity.
-Appendix A Proof of Corollary 1
+In this paper, we proposed a new framework for learning and inference in GPSSM tailored to non-stationary, high-dimensional dynamical systems.
+Our central contribution is the introduction of the ETGP prior, which combines a shared GP with input-dependent normalizing flows parameterized by neural networks.
+This construction enables expressive, non-stationary modeling while reducing the computational and parametric overhead typical of high-dimensional GPSSMs.
+To support scalable inference, we developed a variational framework that approximates the ETGP posterior by disentangling the GP and neural network components.
+The generic and assumption-free variational lower bound we derive requires no explicit latent state posterior parameterization, but instead uses EnKF for the inference, yielding additional computational efficiency.
+Extensive experiments across system identification, high-dimensional filtering, and time series forecasting show that ETGPSSM consistently achieves strong performance while maintaining favorable computational efficiency.
+Appendix A Proof of Proposition 1
+Proof.
+Our target is to compare the two lower bounds on the log-evidence $\log p({\mathbf{y}}_{1:T})$:
+$\mathcal{L}$ $=\mathbb{E}_{q({\mathbf{x}}_{0},\bm{f})}\left[\log\frac{p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0},\bm{f})}{q({\mathbf{x}}_{0},\bm{f})}\right],$ (41)
+$\mathcal{L}_{2}$ $=\mathbb{E}_{q({\mathbf{x}}_{0:T},\bm{f})}\left[\log\frac{p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T},\bm{f})}{q({\mathbf{x}}_{0:T},\bm{f})}\right].$ (42)
+Since
+$q({\mathbf{x}}_{0:T},\bm{f})=q({\mathbf{x}}_{0},\bm{f})q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0}),$
+we have $\mathcal{L}_{2}=$
+$\mathbb{E}_{q({\mathbf{x}}_{0:T},\bm{f})}\left[\log\frac{p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0},\bm{f})p({\mathbf{x}}_{1:T}|{\mathbf{x}}_{0},\bm{f},{\mathbf{y}}_{1:T})}{q({\mathbf{x}}_{0},\bm{f})q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})}\right]$ (43)
+$\!=\!\mathbb{E}_{q({\mathbf{x}}_{0:T},\bm{f})}\left[\log\frac{p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0},\bm{f})}{q({\mathbf{x}}_{0})q(\bm{f})}+\log\frac{p({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0},{\mathbf{y}}_{1:T})}{q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})}\right]$
+$\!=\!\mathcal{L}-\mathbb{E}_{q(\bm{f},{\mathbf{x}}_{0})}\left[\text{KL}\big{(}q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})\|p({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0},{\mathbf{y}}_{1:T})\big{)}\right]$
+Since KL divergence is non-negative:
+$\text{KL}\big{(}q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})\|p({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0},{\mathbf{y}}_{1:T})\big{)}\geq 0$
+with equality iff $q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})=p({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0},{\mathbf{y}}_{1:T})$. Thus we have:
+$\mathcal{L}\geq\mathcal{L}_{2}.$
+Equality holds when $q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})=p({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0},{\mathbf{y}}_{1:T})$ for all $\bm{f}$ and ${\mathbf{x}}_{0}$. ∎
+Appendix B ELBO Derivation
+Proof.
+We derive the ELBO $\mathcal{L}$ for our ETGPSSM by first considering a general stochastic process prior $\bm{f}$ over the transition function (Eq.
+(22)), then specializing to our ETGP construction.
+The derivation proceeds as follows:
+$\mathcal{L}$ $=\mathbb{E}_{q}\left[\log\frac{p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0},\mathbf{w},\tilde{f})}{q({\mathbf{x}}_{0},\mathbf{w},\tilde{f})}\right],$ (44a)
+$=\mathbb{E}_{q}\left[\log\frac{p({\mathbf{y}}_{1:T}|{\mathbf{x}}_{0},\mathbf{w},\tilde{f})p({\mathbf{x}}_{0})p(\mathbf{w})p(\tilde{f})}{q({\mathbf{x}}_{0})q(\mathbf{w})q(\tilde{f})}\right]$ (44b)
+$=\mathbb{E}_{q}\left[\log\frac{p({\mathbf{y}}_{1:T}|{\mathbf{x}}_{0},\mathbf{w},\tilde{f})p({\mathbf{x}}_{0})p(\mathbf{w})p(\tilde{f}|{\mathbf{u}})q({\mathbf{u}})}{q({\mathbf{x}}_{0})q(\mathbf{w})p(\tilde{f}|{\mathbf{u}})q({\mathbf{u}})}\right]$ (44c)
+$=\mathbb{E}_{q}\left[\log p({\mathbf{y}}_{1:T}|{\mathbf{x}}_{0},\mathbf{w},\tilde{f})\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))$
+$\quad-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0})),$ (44d)
+$=\mathbb{E}_{q}\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))$
+$\quad-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0})),$ (44e)
+where we augment the GP using inducing points and use the variational sparse GP in Eq.
+(44c); We factorize the joint likelihood into sequential terms in Eq.
+(44e), where for $t=1$ we have the log-term as $\log p({\mathbf{y}}_{1}|{\mathbf{x}}_{0},\mathbf{w},\tilde{f})$.
+∎
+Appendix C Proof of Corollary 1
 Proof.
@@
 The same logic applies to $\bm{f}({\mathbf{x}}_{t^{\prime}-1})$. Therefore, the covariance matrix $\bm{\Lambda}_{t,t^{\prime}}$ is:
-$\bm{\Lambda}_{t,t^{\prime}}\!=\!\begin{bmatrix}k({\mathbf{x}}_{t-1},{\mathbf{x}}_{t-1})\cdot\bm{\alpha}_{t}\bm{\alpha}_{t}^{\top}&k({\mathbf{x}}_{t-1},{\mathbf{x}}_{t^{\prime}-1})\cdot\bm{\alpha}_{t}\bm{\alpha}_{t^{\prime}}^{\top}\\ k({\mathbf{x}}_{t^{\prime}-1},{\mathbf{x}}_{t-1})\cdot\bm{\alpha}_{t^{\prime}}\bm{\alpha}_{t}^{\top}&k({\mathbf{x}}_{{t^{\prime}}-1},{\mathbf{x}}_{{t^{\prime}}-1})\cdot\bm{\alpha}_{t^{\prime}}\bm{\alpha}_{t^{\prime}}^{\top}\end{bmatrix}.$ (44)
-Appendix B ELBO Derivations and Approximations
+$\begin{bmatrix}k({\mathbf{x}}_{t-1},{\mathbf{x}}_{t-1})\cdot\bm{\alpha}_{t}\bm{\alpha}_{t}^{\top}&k({\mathbf{x}}_{t-1},{\mathbf{x}}_{t^{\prime}-1})\cdot\bm{\alpha}_{t}\bm{\alpha}_{t^{\prime}}^{\top}\\ k({\mathbf{x}}_{t^{\prime}-1},{\mathbf{x}}_{t-1})\cdot\bm{\alpha}_{t^{\prime}}\bm{\alpha}_{t}^{\top}&k({\mathbf{x}}_{{t^{\prime}}-1},{\mathbf{x}}_{{t^{\prime}}-1})\cdot\bm{\alpha}_{t^{\prime}}\bm{\alpha}_{t^{\prime}}^{\top}\end{bmatrix}.$ (50)
+Supplement D Discussions about ELBO Derivation
+Following the GPSSM literature [37, 31], an alternative derivation of the variational inference algorithm presented in this work can begin with the following ELBO, which explicitly defines the variational distribution over the latent states:
+$\mathcal{L}_{2}=\mathbb{E}_{q({\mathbf{x}}_{0:T},\bm{f}(\cdot))}\left[\log\frac{p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T},\bm{f}(\cdot))}{q({\mathbf{x}}_{0:T},\bm{f}(\cdot))}\right].$ (51)
+It turns out that $\mathcal{L}_{2}$ is a lower bound of $\mathcal{L}$ in Eq.
+(22) as summarized in Proposition 1 in the main text.
+Next, we show how we can use the methodology proposed in [37] to derive the same ELBO as in Eq.
+(26).
+The joint distribution of the proposed ETGPSSM is given by:
+$p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T},\mathbf{w},\tilde{f})=p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T}|\mathbf{w},\tilde{f})p_{\bm{\psi}}(\mathbf{w})p(\tilde{f}).$ (52)
+The term $p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T}|\mathbf{w},\tilde{f})$ is factorized as:
+$p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T}|\mathbf{w},\tilde{f})\!=\!p({\mathbf{x}}_{0})\!\prod_{t=1}^{T}\underbracket{p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})}_{\text{likelihood}}\underbracket{p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})}_{\text{transition}},$ (53)
+where the transition term is specified as:
+$p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})=\int p({\mathbf{x}}_{t}|{\mathbf{f}}_{t})p({\mathbf{f}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})\mathrm{d}{\mathbf{f}}_{t}={\cal N}\left({\mathbf{x}}_{t}\mid\bm{\alpha}_{t}\cdot\tilde{f}({\mathbf{x}}_{t-1})+\bm{\beta}_{t},\ \mathbf{Q}\right),$ (54)
+with $p({\mathbf{f}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})=\delta\left({\mathbf{f}}_{t}-(\bm{\alpha}_{t}\cdot\tilde{f}({\mathbf{x}}_{t-1})+\bm{\beta}_{t})\right)$.
+We can employ the following variational distribution to infer the latent variables $\{{\mathbf{x}}_{0:T},\mathbf{w}\}$ and the GP $\tilde{f}$:
+$q({\mathbf{x}}_{0:T},\tilde{f},\mathbf{w})=q(\tilde{f})q(\mathbf{w})q({\mathbf{x}}_{0})\prod_{t=1}^{T}q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1}),$ (55)
+which mirrors the generative structure of the model in Eq. (52) and assumes independence between $\tilde{f}$ and $\mathbf{w}$.
 Given the joint distribution of our model in Eq.
 (52) and the variational distribution in Eq.
-(21), augmented by the sparse GP, we can formulate the ELBO according to the general definition provided in Eq.
+(55), augmented by the sparse GP, we can formulate the ELBO $\mathcal{L}_{2}$ according to the general definition provided in Eq.
 (51).
-The detailed derivations are presented in Eq.
-(45).
-From Eq. (45d) to Eq. (45e), we have made the following two assumptions, similar to the work in [37]:
-  • 1)
-    $p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})\approx p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t})$,
-  • 2)
-    $q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})\approx p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1},{\mathbf{y}}_{1:t})$.
-The transition from Eq. (45f) to Eq. (45g) is derived straightforwardly by applying Bayes’ theorem.
-The two approximations used here are similar to those in [37], with the key difference being the focus on the ETGP in this paper rather than the GP.
+The detailed derivations are presented as follows.
+$\mathcal{L}_{2}$ $=\mathbb{E}_{q}\left[\log\frac{p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T},\mathbf{w},\tilde{f},{\mathbf{u}})}{q({\mathbf{x}}_{0:T},\mathbf{w},\tilde{f},{\mathbf{u}})}\right]$ (56a)
+$=\mathbb{E}_{q}\left[\log\frac{p(\tilde{f},{\mathbf{u}})p(\mathbf{w})p({\mathbf{x}}_{0})\prod_{t=1}^{T}p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})}{q(\tilde{f},{\mathbf{u}})q(\mathbf{w})q({\mathbf{x}}_{0})\prod_{t=1}^{T}q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})}\right]$ (56b)
+$=\mathbb{E}_{q}\left[\sum_{t=1}^{T}\log\frac{p({\mathbf{y}}_{t},{\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})}{q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})}\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0}))$ (56c)
+$=\mathbb{E}_{q}\left[\sum_{t=1}^{T}\log\frac{p({\mathbf{y}}_{t},{\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})}{q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})}\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0}))$ (56d)
+$\approx\mathbb{E}_{q}\left[\sum_{t=1}^{T}\log\frac{p({\mathbf{y}}_{t},{\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})}{\underbracket{p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1},{\mathbf{y}}_{1:t})}_{\text{assumption 1}}\underbracket{p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t})}_{\text{assumption 2}}}\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0}))$ (56e)
+$=\mathbb{E}_{q}\left[\sum_{t=1}^{T}\log\frac{p({\mathbf{y}}_{t},{\mathbf{x}}_{t},{\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})}{p({\mathbf{x}}_{t},{\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1},{\mathbf{y}}_{t})}\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0}))$ (56f)
+$=\mathbb{E}_{q(\mathbf{w},\tilde{f})}\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0})),$ (56g)
+which is the objective function we have used in our paper. Note that from Eq. (56d) to Eq. (56e), we have made the following two assumptions, which is similar to the work in [37]:
+Assumption 1.
+$q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})\approx p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1},{\mathbf{y}}_{1:t}).$ (57)
+Assumption 2.
+$p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})\approx p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t}).$ (58)
+These two approximations used here are similar to those in [37], with the key difference being the focus on the ETGP in this paper rather than the GP.
 The rationale behind these approximations is as follows:
@@
 This trade-off balances accuracy and computational efficiency, a crucial consideration in high-dimensional settings.
-$\mathcal{L}$ $=\mathbb{E}_{q}\left[\log\frac{p({\vec{\mathbf{y}}},{\vec{\mathbf{x}}},\mathbf{w},\tilde{f},{\mathbf{u}})}{q({\vec{\mathbf{x}}},\mathbf{w},\tilde{f},{\mathbf{u}})}\right]$ (45a)
-$=\mathbb{E}_{q}\left[\log\frac{p(\tilde{f},{\mathbf{u}})p(\mathbf{w})p({\mathbf{x}}_{0})\prod_{t=1}^{T}p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})}{q(\tilde{f},{\mathbf{u}})q(\mathbf{w})q({\mathbf{x}}_{0})\prod_{t=1}^{T}q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})}\right]$ (45b)
-$=\mathbb{E}_{q}\left[\sum_{t=1}^{T}\log\frac{p({\mathbf{y}}_{t},{\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})}{q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})}\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0}))$ (45c)
-$=\mathbb{E}_{q}\left[\sum_{t=1}^{T}\log\frac{p({\mathbf{y}}_{t},{\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})}{q({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})}\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0}))$ (45d)
-$\approx\mathbb{E}_{q}\left[\sum_{t=1}^{T}\log\frac{p({\mathbf{y}}_{t},{\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})}{\underbracket{p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1},{\mathbf{y}}_{1:t})}_{\text{assumption 1}}\underbracket{p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t})}_{\text{assumption 2}}}\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0}))$ (45e)
-$=\mathbb{E}_{q}\left[\sum_{t=1}^{T}\log\frac{p({\mathbf{y}}_{t},{\mathbf{x}}_{t},{\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})}{p({\mathbf{x}}_{t},{\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1},{\mathbf{y}}_{t})}\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0}))$ (45f)
-$=\mathbb{E}_{q(\mathbf{w},\tilde{f})}\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0}))$ (45g)
+Remark 5.
+It can be observed that, by adopting the approach from [37], which relies on two specific assumptions, one can also derive the same objective function as presented in our main text.
+However, our contribution in this paper lies in providing an alternative perspective—one that is more straightforward and assumption-free—by leveraging only the EnKF for inference.
+This not only simplifies the derivation but also enhances the practicality of the method.

### v4

@@
 To address these challenges, we propose an efficient transformed Gaussian process state-space model (ETGPSSM) for scalable and flexible modeling of high-dimensional, non-stationary dynamical systems.
-Specifically, our ETGPSSM integrates a single shared GP with input-dependent normalizing flows, yielding an expressive implicit process prior that captures complex, non-stationary transition dynamics while significantly reducing model complexity.
+Specifically, our ETGPSSM integrates a single shared GP with input-dependent normalizing flows, yielding an expressive non-stationary implicit process prior that can capture complex transition dynamics while significantly reducing model complexity.
 For the inference of the implicit process, we develop a variational inference algorithm that jointly approximates the posterior over the underlying GP and the neural network parameters defining the normalizing flows.
@@
 Extensive empirical evaluations on synthetic and real-world datasets demonstrate the superior performance of our ETGPSSM in system dynamics learning, high-dimensional state estimation, and time-series forecasting, outperforming existing GPSSMs and neural network-based SSMs in terms of computational efficiency and accuracy.
-Index Terms:
-Gaussian process, state-space model, normalizing flows, variational inference, high-dimensional dynamical systems.
 I Introduction
-State-space models (SSMs) provide a flexible framework for modeling dynamical systems, where latent states evolve based on internal dynamics and external inputs [1].
+State-space models (SSMs) provide a flexible framework for modeling dynamical systems, where latent states evolve based on internal dynamics and external inputs [sarkka2013bayesian].
 When system dynamics are fully known, Bayesian filtering methods—such as the Kalman filter (KF), extended Kalman filter (EKF), ensemble Kalman filter (EnKF), and particle filter (PF)—enable sequential state estimation by combining prior knowledge with incoming observations, yielding reliable predictions under uncertainty.
-This capability makes SSMs widely applicable in robotics, finance, and signal processing, where modeling temporal dependencies is crucial [2, 3, 4].
-However, in complex real-world systems, including nonlinear and high-dimensional problems such as climate modeling, physiological processes, and neural dynamics, the underlying dynamics may be too intricate or poorly understood to specify a physics-informed or explicitly parameterized model.
-In these cases, traditional filtering methods cannot be applied in their standard form and must instead rely on data-driven approaches to learn the dynamics from observations, complicating the task of accurate and robust state estimation [5].
-To address this challenge, a variety of data-driven system identification approaches have been proposed.
-Deterministic models, such as deep state-space models (DSSMs), employ neural networks to learn system dynamics and have gained popularity for their ability to capture high-dimensional, nonlinear behaviors [6].
+This capability makes SSMs widely applicable in robotics, finance, and signal processing, where modeling temporal dependencies is crucial [kullberg2021online, khan2008distributing, tobar2015unsupervised].
+However, in complex real-world systems, such as climate modeling, physiological processes, and neural dynamics, the underlying dynamics are often highly nonlinear, high-dimensional, and insufficiently understood to permit a physics-informed or explicitly parameterized model.
+In such settings, traditional filtering methods cannot be applied in their standard form and must instead resort to data-driven approaches that learn the system dynamics directly from observations.
+This shift significantly complicates the task of achieving accurate and robust state estimation [revach2022kalmannet].
+A variety of data-driven approaches for modeling dynamical systems have been proposed.
+Deterministic models, such as deep state-space models (DSSMs), employ neural networks to learn system dynamics and have gained popularity for their ability to capture high-dimensional, nonlinear behaviors [girin2021dynamical].
 However, these models often require large training datasets to avoid overfitting, and their generalization performance may degrade in uncertain or data-scarce environments.
-Moreover, their predictions can be difficult to interpret, especially when compared to more structured probabilistic models [7].
-As a result, DSSMs may face challenges in practical deployment despite their modeling capacity [8].
-Stochastic models based on flexible random process priors provide a principled alternative within a Bayesian learning framework [9].
-A prominent example is the class of Gaussian process state-space models (GPSSMs), which leverage Gaussian processes (GPs) to capture uncertainty and provide implicit regularization through prior structure.
-These properties make GPSSMs attractive for applications where uncertainty quantification and data efficiency are critical [10, 11].
-GPSSMs have been applied in diverse domains, including human pose and motion learning [12], robotics and control [13], reinforcement learning [14, 15], target tracking and navigation [16], and magnetic-field sensing [17].
-Research on GPSSMs has also focused on advancing learning and inference methodologies.
-Early works either assumed a pre-learned GPSSM and concentrated solely on latent state inference—an impractical assumption for many real-world applications [18, 19, 20]—or employed maximum a posteriori (MAP) estimation to jointly infer the model and the latent states [12, 21], which is prone to overfitting.
-To address these issues, a important development came with the fully Bayesian treatment of GPSSMs using particle Markov chain Monte Carlo (PMCMC) methods [22], though the inherent cubic computational complexity of GP remains prohibitive for long time series.
-Subsequent efforts [23, 24, 25, 26] adopted reduced-rank GP approximations to alleviate the computational demands, yet PMCMC scalability remains a challenge.
-This has led to a growing interest in variational inference methods [10, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] that can leverage sparse GP approximations with inducing points to enhance scalability [40, 41].
+Moreover, their predictions can be difficult to interpret, especially when compared to more structured probabilistic models [krishnan2017structured].
+As a result, DSSMs may face challenges in practical deployment despite their modeling capacity [ghosh2024danse].
+Stochastic models based on flexible random process priors provide a principled alternative within a Bayesian learning framework [williams2006gaussian].
+A prominent example is the class of Gaussian process state-space models (GPSSMs), which leverage Gaussian processes (GPs) to encapsulate uncertainty and provide implicit regularization through prior structure.
+These properties make GPSSMs particularly attractive for applications where uncertainty quantification and data efficiency are critical [frigola2015bayesian, suwandi2023sparsityaware].
+GPSSMs have been applied in diverse domains, including human pose and motion learning [wang2007gaussian], robotics and control [deisenroth2013gaussian], reinforcement learning [arulkumaran2017deep, yan2020gaussian], target tracking and navigation [xie2020learning], and magnetic-field sensing [berntorp2023constrained].
+In parallel to their practical applications, significant research efforts have been dedicated to advancing the learning and inference algorithms that underpin GPSSMs.
+Early works either assumed a pre-learned GPSSM and concentrated solely on latent state inference—an impractical assumption for many real-world applications [ko2009gp, deisenroth2009analytic, deisenroth2011robust]—or employed maximum a posteriori (MAP) estimation to jointly infer the model and the latent states [wang2007gaussian, ko2011learning], which is prone to overfitting.
+To address these issues, a important development came with the fully Bayesian treatment of GPSSMs using particle Markov chain Monte Carlo (PMCMC) methods [frigola2013bayesian], though the inherent cubic computational complexity of GP remains prohibitive for long time series.
+Subsequent efforts adopted reduced-rank GP approximations to alleviate the computational demands [svensson2016computationally, svensson2017flexible, berntorp2021online, liu2023sequential], yet PMCMC scalability remains a challenge.
+This has led to a growing interest in variational inference methods [frigola2015bayesian, frigola2014variational, mchutchon2015nonlinear, eleftheriadis2017identification, doerr2018probabilistic, ialongo2019overcoming, curi2020structured, liu2020gpssm, lindinger2022laplace, fan2023free, lin2023towards, lin2023ensemble, lin2023towards_efficient, lin2022output] that can leverage sparse GP approximations with inducing points to enhance scalability [titsias2009variational, hensman2013gaussian].
 However, existing GPSSMs, which employ multiple independent GPs to model system dynamics, remain limited to low-dimensional state spaces and face fundamental challenges when scaling to high-dimensional systems.
@@
 Additionally, all GPSSMs rely on stationary GP priors, which assume that the statistical properties of the transition function are invariant under input translation.
-This assumption often fails for systems with time-varying or non-stationary dynamics, limiting the ability of GPSSMs to accurately capture complex real-world behaviors.
-To address these challenges, we propose a novel approach for learning and inference in GPSSMs tailored to non-stationary, high-dimensional dynamical systems.
-Our primary contribution is the development and integration of an efficient yet flexible implicit process prior [42, 43] for the state transition function.
-This prior combines the strengths of GPs, normalizing flows [44], and (Bayesian) neural networks, significantly enhancing modeling capabilities in high-dimensional state spaces.
-Our key contributions are summarized as follows:
-  • Instead of employing multiple independent GPs for the transition function, we propose applying an efficient transformed Gaussian process (ETGP) [42], an implicit random process that combines a shared GP with normalizing flows.
+This assumption often fails for systems exhibiting time-varying or non-stationary dynamics, limiting the expressiveness and practical applicability of conventional GPSSMs.
+To this end, we propose a novel GPSSM framework for non-stationary, high-dimensional dynamical systems, whose primary contribution is an efficient, flexible implicit process prior for the state transition function.
+By unifying Gaussian processes, input-dependent normalizing flows [kobyzev2020normalizing], and (Bayesian) neural networks, this prior is inherently non-stationary while maintaining computational and parametric complexity that scales gracefully with state dimensionality.
+The key contributions are summarized as follows:
+  • Instead of employing multiple independent GPs for the transition function, we propose applying an efficient transformed Gaussian process (ETGP) [maronas2022efficient], an implicit random process that combines a shared GP with normalizing flows.
 Crucially, the flow parameters are made input-dependent through neural network modeling.
@@
 It achieves significantly lower computational and parametric complexity than conventional GPSSMs while outperforming state-of-the-art approaches, including deep neural network-based SSMs, in terms of learning and inference accuracy.
-Part of this work was presented at IEEE ICASSP 2024 [38].
+Part of this work was presented at IEEE ICASSP 2024 [lin2023towards_efficient].
 The current paper introduces input-dependent normalizing flows to further enhance model flexibility and proposes an EnKF-aided variational inference algorithm for efficient state estimation.
@@
 In this work, the emission function is assumed to be linear, parameterized by a matrix $\bm{C}\in\mathbb{R}^{d_{y}\times d_{x}}$.
-For systems with nonlinear emissions, the latent state can be augmented to a higher-dimensional representation, effectively rendering the emission function linear [30].
-This reparameterization helps mitigate non-identifiability issues frequently encountered in data-driven SSMs [10].
+For systems with nonlinear emissions, the latent state can be augmented to a higher-dimensional representation, effectively rendering the emission function linear [doerr2018probabilistic].
+This reparameterization helps mitigate non-identifiability issues frequently encountered in data-driven SSMs [frigola2015bayesian].
 Both latent states and observations are perturbed by zero-mean Gaussian noise, with covariance matrices $\mathbf{Q}$ and $\mathbf{R}$, respectively.
+${{\mathbf{x}}}_{0}$ ${{\mathbf{x}}}_{1}$ $\cdots$ ${{\mathbf{x}}}_{t-1}\!\!$ ${{\mathbf{x}}}_{t}$ $\cdots$ $\cdots$ ${{\mathbf{f}}}_{1}$ $\cdots$ ${{\mathbf{f}}}_{t-1}\!\!$ ${{\mathbf{f}}}_{t}$ $\cdots$ ${{\mathbf{y}}}_{1}$ $\cdots$ ${{\mathbf{y}}}_{t-1}\!\!$ ${{\mathbf{y}}}_{t}$ $\cdots$
+Figure 1: Graphical representation of GPSSM.
+The white circles represent the latent variables, while the gray circles represent the observable variables.
+The thick horizontal bar represents a set of fully connected, mutually correlated nodes, specifically, the GP.
 II-B Gaussian Process State-Space Models (GPSSMs)
-As discussed in Section I, system dynamics in real-world complex scenarios are often unknown, with only a set of noisy observations, ${\mathbf{y}}_{1:T}\!=\!\{{\mathbf{y}}_{t}\}_{t=1}^{T}$, available.
-Consequently, there has been growing interest in adopting data-driven approaches, such as GPs [9], to model these dynamics and quantify the underlying uncertainties.
+As discussed in Section I, system dynamics in real-world complex scenarios are often unknown, but only a set of noisy observations, ${\mathbf{y}}_{1:T}\!=\!\{{\mathbf{y}}_{t}\}_{t=1}^{T}$, where $T\in\mathbb{Z}_{+}$, is available.
+Consequently, there has been growing interest in adopting data-driven approaches, such as GPs [williams2006gaussian], to model these dynamics and quantify the underlying uncertainties.
 II-B1 Gaussian process (GP)
 The GP is a generalization of the Gaussian distribution over infinite index sets, enabling the specification of distributions over functions $\tilde{f}(\cdot):\mathbb{R}^{d_{x}}\mapsto\mathbb{R}$.
-A GP is fully characterized by its mean function, often set to zero, and its covariance (kernel) function $k({\mathbf{x}},{\mathbf{x}}^{\prime})$, which includes a set of hyperparameters $\bm{\theta}_{gp}$ optimized during model selection [9].
-By definition, the function values $\tilde{{\mathbf{f}}}=\{\tilde{f}({\mathbf{x}}_{i})\}_{i=1}^{N}\in\mathbb{R}^{N}$, evaluated at a finite set of inputs $\mathbf{X}=\{{\mathbf{x}}_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times d_{x}}$, follow a joint Gaussian distribution:
+Definition 1.
+A stochastic process $\{\tilde{f}(\mathbf{x}):\mathbf{x}\in\mathcal{X}\}$ indexed by a domain $\mathcal{X}\subseteq\mathbb{R}^{d_{x}}$ is said to be a GP [williams2006gaussian] if, for every finite set of inputs $\mathbf{X}=\{{\mathbf{x}}_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d_{x}}$, the corresponding random vector $\tilde{{\mathbf{f}}}=\{\tilde{f}({\mathbf{x}}_{i})\}_{i=1}^{N}$ is jointly Gaussian.
+A GP is fully characterized by its mean function, often set to zero, and its covariance (kernel) function $k({\mathbf{x}},{\mathbf{x}}^{\prime})$, which includes a set of hyperparameters $\bm{\theta}_{gp}$ [williams2006gaussian].
+We denote this as
+$\tilde{f}({\mathbf{x}})\sim\mathcal{GP}\big(0,\,k({\mathbf{x}},{\mathbf{x}}^{\prime}\,;{{\bm{\theta}}}_{gp})\big).$
+A widely used family of kernels is the class of stationary kernels [williams2006gaussian], which is formally defined as follows.
+Definition 2.
+A kernel $k$ is stationary if there exists a function $\kappa:\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{x}}\to\mathbb{R}$, such that $k(\mathbf{x},\mathbf{x}^{\prime})=\kappa(\mathbf{x}-\mathbf{x}^{\prime})$ for all ${\mathbf{x}},{\mathbf{x}}^{\prime}\subset\cal X$.
+This assumption encodes translation-invariant properties and is shared by common kernels such as the squared exponential (SE) and Matérn families [williams2006gaussian].
+Together with a constant mean function, a stationary kernel defines a stationary GP, whose finite-dimensional distributions are invariant to translations of the input space.
+Specifically, for any inputs $\mathbf{X}$, the distribution
 $p(\tilde{{\mathbf{f}}}\mid\mathbf{X})=\mathcal{N}(\tilde{{\mathbf{f}}}\mid\bm{0},\,\mathbf{K}),$ (2)
-where $\mathbf{K}\in\mathbb{R}^{N\times N}$ is the covariance matrix with entries $[\mathbf{K}]_{i,j}=k({\mathbf{x}}_{i},{\mathbf{x}}_{j})$.
-Given observed function values $\tilde{{\mathbf{f}}}$ at inputs $\mathbf{X}$, the GP prediction at a new input ${\mathbf{x}}_{*}$, denoted $p(\tilde{f}({\mathbf{x}}_{*})\mid\tilde{{\mathbf{f}}},\mathbf{X})$, is also Gaussian, with posterior mean $\xi({\mathbf{x}}_{*})$ and posterior variance $\Xi({\mathbf{x}}_{*})$ given by:
+where $\mathbf{K}\in\mathbb{R}^{N\times N}$ and $\mathbf{K}]_{i,j}=k({\mathbf{x}}_{i},{\mathbf{x}}_{j}\,;{{\bm{\theta}}}_{gp})$, remains unchanged under translation.
+Given function values $\tilde{{\mathbf{f}}}$, the GP prediction at a new input ${\mathbf{x}}_{*}\!\subset\!\mathcal{X}$, denoted $p(\tilde{f}({\mathbf{x}}_{*})|\tilde{{\mathbf{f}}},\mathbf{X})$, is also Gaussian, with posterior mean $\xi({\mathbf{x}}_{*})$ and posterior variance $\Xi({\mathbf{x}}_{*})$ given by:
 $\xi(\mathbf{x}_{*})=\mathbf{K}_{\mathbf{x}_{*},\mathbf{X}}\mathbf{K}^{-1}\tilde{{\mathbf{f}}}$, (3a)
@@
 II-B2 Gaussian process state-space model
-Placing a GP prior over the transition function $\bm{f}(\cdot)$ in the classic SSM (see Eq. (1)) yields the GPSSM [10]:
+Placing a GP prior over the transition function $\bm{f}(\cdot)$ in the classic SSM (see Eq. (1)) yields the GPSSM [frigola2015bayesian]:
 $\bm{f}(\cdot)\sim\mathcal{GP}\left(\bm{0},\bm{k}(\cdot,\cdot)\right),$ (4a)
@@
 $\mathbf{y}_{t}|\mathbf{x}_{t}\sim{\cal N}\left({\mathbf{y}}_{t}|\bm{C}\mathbf{x}_{t},\mathbf{R}\right),$ (4e)
-where the initial state $p({\mathbf{x}}_{0})$ is assumed known and Gaussian. A graphical representation is shown in Fig. 1.
-When the state dimension $d_{x}>1$, the transition function $\bm{f}(\cdot):\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{x}}$ is typically modeled using $d_{x}$ independent GPs.
-Each component function $f_{d}(\cdot):\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}$ maps to one output dimension:
+where the initial state $p({\mathbf{x}}_{0})$ is assumed known and Gaussian for simplicity. A graphical representation is shown in Fig. 1.
+When the state dimension $d_{x}>1$, the transition function $\bm{f}(\cdot):\mathbb{R}^{d_{x}}\mapsto\mathbb{R}^{d_{x}}$ is typically modeled using $d_{x}$ independent GPs.
+Each component function $f_{d}(\cdot):\mathbb{R}^{d_{x}}\mapsto\mathbb{R}$ maps to one output dimension:
 ${\mathbf{f}}_{t}=\bm{f}({\mathbf{x}}_{t-1})=\{{f}_{d}({\mathbf{x}}_{t-1})\}_{d=1}^{d_{x}}=\{\mathrm{f}_{t,d}\}_{d=1}^{d_{x}},$ (5)
-where each GP has its own kernel function and hyperparameters [30].
-The central challenge in GPSSMs lies in jointly learning the transition dynamics and noise parameters, i.e., learning $[{{\bm{\theta}}}_{gp},\mathbf{Q},\mathbf{R}]$, while simultaneously inferring both the latent states and the underlying functions.
-${{\mathbf{x}}}_{0}$ ${{\mathbf{x}}}_{1}$cdots $\!\!{{\mathbf{x}}}_{t-1}\!\!$ ${{\mathbf{x}}}_{t}$cdots $\cdots$ ${{\mathbf{f}}}_{1}$cdots $\!\!{{\mathbf{f}}}_{t-1}\!\!$ ${{\mathbf{f}}}_{t}$cdots ${{\mathbf{y}}}_{1}$cdots $\!\!{{\mathbf{y}}}_{t-1}\!\!$ ${{\mathbf{y}}}_{t}$cdots
-Figure 1: Graphical representation of GPSSM.
-The white circles represent the latent variables, while the gray circles represent the observable variables.
-The thick horizontal bar represents a set of fully connected, mutually correlated nodes, specifically, the GP.
+where each GP has its own kernel function and hyperparameters [doerr2018probabilistic].
+Given a single observed sequence $\{\mathbf{y}_{t}\}_{t=1}^{T}$, the central task in GPSSMs involves jointly learning the transition dynamics and noise parameters, i.e., learning ${{\bm{\theta}}}_{gp},\mathbf{Q},\mathbf{R}]$, while simultaneously inferring both the corresponding latent states and the underlying functions.
 II-C Challenges in Existing Variational GPSSMs
-Note that in Eq.
-(3), evaluating the GP posterior requires inverting the covariance matrix $\mathbf{K}$, incurring a computational cost of $\mathcal{O}(N^{3})$, which becomes prohibitive for large $N$.
-To address this, variational GPSSMs typically employ sparse GP approximations [40, 41], a widely used technique in GP-based models [45, 10] due to its analytical tractability and reduced computational cost.
+In contrast to the standard GP described in Eq.
+(3), where the GP posterior is computed directly from observed input-output pairs, inferring the GP in a GPSSM is more complex.
+The key challenge is that the inputs to the transition function are the latent states $\{\mathbf{x}_{t}\}_{t=0}^{T-1}$, which are themselves unobserved.
+Consequently, the GP posterior over the transition function must be inferred jointly with the latent state trajectory.
+This is typically done approximately via variational inference [frigola2014variational].
+The variational GPSSMs employ sparse GP approximations [titsias2009variational, hensman2013gaussian], a widely used technique in GP-based models [maronas2021transforming, frigola2015bayesian] due to its analytical tractability and reduced computational cost.
 The core idea is to augment the GP with a smaller set of inducing points, which act as a sparse approximation of the full GP.
-More specifically, we can introduce $M\!\ll\!N$ inducing inputs $\mathbf{Z}=\{{\mathbf{z}}_{i}\}_{i=1}^{M}$ (where ${\mathbf{z}}_{i}\in\mathbb{R}^{d_{x}}$) and their corresponding inducing outputs ${\mathbf{u}}=\{\mathrm{u}_{i}\}_{i=1}^{M}=\{\tilde{f}({\mathbf{z}}_{i})\}_{i=1}^{M}$, which inherit the same GP prior as the original process, i.e., ${\mathbf{u}}\sim\mathcal{N}(\bm{0},\mathbf{K}_{\mathbf{Z},\mathbf{Z}})$.
-The augmented GP prior can be written as: [^1] With slight notation abuse, we denote the distribution of GP prior as $p(\tilde{f})$.
+More specifically, we can introduce $M\!\ll\!N$ inducing inputs $\mathbf{Z}=\{{\mathbf{z}}_{i}\}_{i=1}^{M}\subset\mathbb{R}^{d_{x}}$ and their corresponding inducing outputs ${\mathbf{u}}=\{\mathrm{u}_{i}\}_{i=1}^{M}=\{\tilde{f}({\mathbf{z}}_{i})\}_{i=1}^{M}$, which inherit the same GP prior as the original process, i.e., ${\mathbf{u}}\sim\mathcal{N}(\bm{0},\mathbf{K}_{\mathbf{Z},\mathbf{Z}})$.
+The augmented GP prior can be written as: [^1] With slight notation abuse, we denote the distribution of GP prior as p(f~)p($\tilde${f}).
 $p(\tilde{f},{\mathbf{u}})=p(\tilde{f}|\mathbf{u})p(\mathbf{u}).$ (6)
-To approximate the posterior, a variational distribution is defined over the $M$ inducing outputs, $q({\mathbf{u}})=\mathcal{N}({\mathbf{u}}\mid\mathbf{m},\mathbf{S})$, while the prior conditional is used for the rest of the GP:
+To approximate the posterior, a variational distribution is defined over the $M$ inducing outputs, $q({\mathbf{u}})\!=\!\mathcal{N}({\mathbf{u}}|\mathbf{m},\mathbf{S})$, while the prior conditional is used for the rest of the GP:
 $q(\tilde{f},{\mathbf{u}})=p(\tilde{f}|\mathbf{u})q(\mathbf{u}).$ (7)
@@
 Specifically, variational GPSSMs define inducing inputs $\{\mathbf{Z}_{d}\}_{d=1}^{d_{x}}\in\mathbb{R}^{d_{x}\times(M\times d_{x})}$, variational means $\{\mathbf{m}_{d}\}_{d=1}^{d_{x}}\in\mathbb{R}^{d_{x}\times M}$, and covariance matrices $\{\mathbf{S}_{d}\}_{d=1}^{d_{x}}\in\mathbb{R}^{d_{x}\times(M\times M)}$ for the $d_{x}$-dimensional inducing outputs $\{\mathbf{u}_{d}\}_{d=1}^{d_{x}}$.
-As the state dimension increases, the quadratic scaling of inducing inputs and the associated variational parameters leads to significant computational overhead [38].
-  • Most GPSSMs impose a stationary GP prior over the transition function, as shown in Eq.
-(4a), to model time-invariant dynamics.
-However, such stationary assumptions may fail to capture time-varying or context-dependent behavior [^2] Here, “stationarity” refers to the GP prior over the transition function, not the observed time series.
-A GPSSM with a stationary kernel may still model non-stationary observed data..
-Additionally, modeling each output dimension with an independent GP ignores correlations across dimensions, potentially limiting modeling capacity.
-In the next section, we detail our proposed method to resolve these issues.
+As the state dimension increases, the quadratic scaling of inducing inputs and the associated variational parameters leads to significant computational overhead [lin2023towards_efficient].
+  • Most GPSSMs impose a stationary GP prior on the transition function, as shown in Eq.
+(4a).
+However, this stationarity assumption may be inadequate for capturing time-varying or input-dependent dynamical behavior.
+Additionally, modeling each output dimension with an independent GP ignores potential correlations across state dimensions, which can limit the model’s capacity.
+It is important to clarify that the “stationarity” in this work refers specifically to the properties of the transition function prior, not the observed time series $\mathbf{y}_{1:T}$.
+A non-stationary transition function is one whose characteristics (e.g., amplitude, smoothness, or periodicity) vary with the input state.
+Notably, a GPSSM with a stationary kernel in its transition function can still generate a non-stationary time series.
+In the next section, we present a methodology to resolve these issues.
 III Efficient Transformed Gaussian Process State-Space Models
@@
 III-A Efficient Modeling for High-Dimensional State Space
-Figure 2: (2(a)) Two warped GPs, see Eq.
-(10).
-$f_{1}(\cdot)$ effectively fits the data, while $f_{2}(\cdot)$ falls short.
-(2(b)) Independent GPs ($\tilde{f}_{1}(\cdot)$, $\tilde{f}_{2}(\cdot)$) vs our proposed ETGP (${f}_{1}(\cdot)$, ${f}_{2}(\cdot)$) depicted in Fig.
-3.
-Both methods capture the underlying function reasonably well, though the GP $\tilde{f}_{1}(\cdot)$ exhibits slightly worse fit and overestimated uncertainty.
+Figure 2: Two-dimensional regression example comparing GP models. The non-stationary ETGP accurately captures local variations, outperforming stationary warped GPs and independent GPs.
 In Section II-C, we discussed the computational and modeling challenges that arise from using separate GPs for each output dimension of the transition function in high-dimensional state spaces.
-To address these limitations, [38] proposed warping a single shared GP using normalizing flows [44] to model each output dimension of the transition function:
+To address these limitations, [lin2023towards_efficient] proposed warping a single shared GP using normalizing flows [kobyzev2020normalizing] to model each output dimension of the transition function:
 $\tilde{f}(\cdot)\!\sim\!\mathcal{GP}(0,k(\cdot,\cdot)),~f_{d}(\cdot)\!=\!{\mathbb{G}}_{{{\bm{\theta}}}_{d}}(\tilde{f}(\cdot)),~d\!=\!1,\ldots,d_{x},$ (10)
@@
 Example 2.
-If ${\mathbb{G}}_{{{\bm{\theta}}}_{d}}(\cdot)$ is a Sinh-Arcsinh-Linear (SAL) flow [45, 46], we have:
+If ${\mathbb{G}}_{{{\bm{\theta}}}_{d}}(\cdot)$ is a Sinh-Arcsinh-Linear (SAL) flow [maronas2021transforming, rios2019compositionally], we have:
 $f_{d}(\cdot)=\alpha_{d}\sinh\left[\varphi_{d}\operatorname{arcsinh}\left(\tilde{f}(\cdot)\right)-\gamma_{d}\right]+\beta_{d},$
 with ${{\bm{\theta}}}_{d}=[\alpha_{d},\beta_{d},\gamma_{d},\varphi_{d}]^{\top}\in\mathbb{R}^{4}$.
-More complex flows can be achieved by stacking additional layers of SAL flow and/or incorporating more advanced deep neural network-based flows, such as RealNVP [47].
+More complex flows can be achieved by stacking additional layers of SAL flow and/or incorporating more advanced deep neural network-based flows, such as RealNVP [dinh2017density].
 The main advantage of this approach lies in eliminating the need for multiple independent GPs by using computationally efficient flow transformations.
@@
 However, the warped GP in Eq.
-(10) remains stationary; Also, being derived from a single shared GP, it may lack the expressiveness needed to capture complex, time-varying latent dynamics—resulting in suboptimal performance in practice [38].
-An illustrative example is shown in Fig.
-2(a), where the learned element-wise flow $\mathbb{G}_{\bm{\theta}_{1}}$ successfully warps $\tilde{f}$ to fit a simple target function $f_{1}$.
-In contrast, the warped function $f_{2}$ struggles to capture the time-varying and intricate structure, resulting in high uncertainty.
-For a more comprehensive comparison, Fig.
-2(b) shows the results from both two independent GPs and the ETGP model introduced later in this work.
-While both capture the underlying structure, the stationary GP $\tilde{f}_{1}(\cdot)$ yields inferior fits and inflated uncertainty estimates.
-To overcome these limitations in expressiveness, we draw inspiration from recent advances in adaptive parameterization [45, 48] and introduce non-stationarity and input dependence into the warped GPs.
+(10) remains stationary; Also, being derived from a single shared GP, it may lack the expressiveness needed to capture complex latent dynamics—resulting in suboptimal performance in practice [lin2023towards_efficient].
+A demonstrative example is provided as follows.
+Example 3.
+Consider a two-dimensional regression problem where the outputs are generated as follows:
+$y_{1}$ $=3.5\cdot h(x),$
+$y_{2}$ $=-1.5\cdot h(x)+5\sin(\pi x)+2\cos(2\pi x),$
+where $h(x)=0.8+(x+0.2)\left(1-\frac{5}{1+\exp(-2x)}\right)+\varepsilon$ is a noisy kink function, and noise $\varepsilon\sim\mathcal{N}(0,0.25^{2})$.
+The function $y_{1}$ is linearly dependent on the noisy kink function, and $y_{2}$ introduces additional nonlinear, periodic components and a negative correlation with $y_{1}$.
+A dataset of 100 input-output pairs was generated with inputs in [4,2][-4,2], as shown in Fig.
+2.
+The warped GPs, which are constructed by a base GP with an SE kernel and two SAL flows, were trained by maximizing the variational lower bound, see e.g.
+[maronas2021transforming, Section 4.1] for more details.
+The results in Fig.
+2(a) show that the learned element-wise SAL flow $\mathbb{G}_{\bm{\theta}_{1}}$ successfully warps the base GP $\tilde{f}$ into $f_{1}$, which fits $y_{1}$ well.
+In contrast, applying $\mathbb{G}_{\bm{\theta}_{2}}$ to the same base process yields $f_{2}$, which fails to capture the complex structure of $y_{2}$ due to the limited modeling flexibility of the input-independent SAL flows (see Example 2), resulting in high predictive uncertainty.
+To overcome the limitations in expressiveness, we draw inspiration from recent advances in adaptive parameterization [maronas2021transforming, gu2023mamba] and introduce non-stationarity and input dependence into the warped GPs.
 Specifically, we generate non-stationary processes $\{f_{d}(\cdot)\}_{d=1}^{d_{x}}$ by conditioning the flow parameters on the input through a shallow neural network $\textsc{nn}:\mathbb{R}^{d_{x}}\mapsto\Theta$ parameterized by weights $\mathbf{w}$.
-In the SSM context, this is formulated as [^3] Note that ${\mathbf{x}}_{t}$ can always be extended to accommodate control systems incorporating a deterministic control input $\bm{c}_{t}\in\mathbb{R}^{d_{c}}$ by augmenting the latent state to $[{\mathbf{x}}_{t}^{\top},\bm{c}_{t}^{\top}]^{\top}\in\mathbb{R}^{d_{x}+d_{c}}$.
+In the SSM context, this is formulated as [^2] Note that ${\mathbf{x}}_{t}$ can always be extended to accommodate control systems incorporating a deterministic control input $\bm{c}_{t}\in\mathbb{R}^{d_{c}}$ by augmenting the latent state to ${\mathbf{x}}_{t}^{\top},\bm{c}_{t}^{\top}]^{\top}\in\mathbb{R}^{d_{x}+d_{c}}$.
 For brevity, however, we omit explicit reference to $\bm{c}_{t}$ in our notation throughout this paper.:
 ${{\bm{\theta}}}_{t}=\textsc{nn}_{\mathbf{w}}({\mathbf{x}}_{t-1}),$ (11)
-where ${{\bm{\theta}}}_{t}=\{{{\bm{\theta}}}_{t,d}\}_{d=1}^{d_{x}}$.
-This formulation allows the flow parameters to vary with input at each time step $t$, enabling the model to capture local variations and time-varying dynamics.
+where ${{\bm{\theta}}}_{t}=\{\theta_{t,d}\}_{d=1}^{d_{x}}$.
+This formulation allows the flow parameters to vary with input at each time step $t$, enabling the model to capture local variations and complex dynamics.
 Additionally, note that the parameter amortization learning shown in Eq.
 (11) prevents the number of flow parameters from growing with time step $t$ while also improving parameter efficiency through weight sharing across dimensions.
+Figure 3: Non-stationary ETGP transition function in SSMs
 We may further impose a prior $p_{\bm{\psi}}(\mathbf{w})$ over $\mathbf{w}$ with hyperparameter $\bm{\psi}$ (e.g., $\mathbf{w}\!\sim\!{\cal N}(\mathbf{w}|\bm{0},\operatorname{diag}(\bm{\psi}))$) to regularize the neural network’s parameters and mitigate overfitting issues.
-In summary, our modeling approach can be expressed mathematically as follows:
+In summary, this efficient transformed GP modeling can be mathematically expressed as follows:
 $\tilde{f}(\cdot)\sim\mathcal{GP}(0,k(\cdot,\cdot)),\ \ \mathbf{w}\sim p_{\bm{\psi}}(\mathbf{w}),$ (12a)
-${{\bm{\theta}}}_{t}=\textsc{nn}_{\mathbf{w}}({\mathbf{x}}_{t-1}),\ \ f_{d}(\cdot)={\mathbb{G}}_{{{\bm{\theta}}}_{t,d}}(\tilde{f}(\cdot)),\ d\!=\!1,...,{d_{x}},$ (12b)
-where $\{f_{d}(\cdot)\}_{d=1}^{d_{x}}$ implicitly characterize $d_{x}$ random processes [43].
+${{\bm{\theta}}}_{t}=\textsc{nn}_{\mathbf{w}}({\mathbf{x}}_{t-1}),\ \ f_{d}(\cdot)={\mathbb{G}}_{\theta_{t,d}}(\tilde{f}(\cdot)),\ d\!=\!1,...,{d_{x}},$ (12b)
+where $\{f_{d}(\cdot)\}_{d=1}^{d_{x}}$ implicitly characterize $d_{x}$ random processes [ma2019variational].
 To better understand the modeling, Fig.
-3 illustrates the corresponding graphical representation with $p_{\bm{\psi}}(\mathbf{w})=\delta(\mathbf{w}-\bar{\mathbf{w}})$, where $\delta(\cdot)$ is the Dirac delta measure and $\bar{\mathbf{w}}$ is a deterministic vector, and the following example extends the Example 1, providing a more concrete illustration.
+3 illustrates the corresponding graphical representation with $p_{\bm{\psi}}(\mathbf{w})=\delta(\mathbf{w}-\bar{\mathbf{w}})$, where $\delta(\cdot)$ is the Dirac delta measure and $\bar{\mathbf{w}}$ is a deterministic vector.
+The following example extends the Example 1, providing a more concrete illustration.
 Example 4.
@@
 where ${{\bm{\theta}}}_{t}\!=\![\bm{\alpha}_{t}^{\top},\bm{\beta}_{t}^{\top}]^{\top}$, and $\bm{\alpha}_{t}\!=\!\{\alpha_{t,d}\}_{d=1}^{d_{x}}$, $\bm{\beta}_{t}\!=\!\{\beta_{t,d}\}_{d=1}^{d_{x}}$.
-An illustrative example of function fitting using the ETGP with linear flows is depicted in Fig.
-2(b).
-Figure 3: ETGP transition function in SSMs
-It is noteworthy that this modeling strategy, referred to as the efficient transformed Gaussian process (ETGP) in the GP literature, was originally proposed to tackle large-scale multi-class classification problems [42].
+An illustrative two-dimensional regression example using the resulting transformed GPs, ${f}_{1}$ and ${f}_{2}$, is shown in Fig.
+2(b), where the neural network weights follow a Dirac delta measure.
+The training data consist of 100 input–output pairs consistent with Example 3.
+For comparison, results from two independent GPs, $\tilde{f}_{1}$ and $\tilde{f}_{2}$, with SE kernels are also presented.
+Both models were trained by maximizing their respective variational lower bounds [hensman2013gaussian, maronas2022efficient].
+While both models capture the general structure of the target functions, the independent GP $\tilde{f}_{2}$ exhibits a poorer fit and overestimates uncertainty for the more complex $y_{2}$, primarily due to its stationary SE kernel’s inability to adapt to local variations.
+In contrast, the ETGP achieves a notably better fit for these complex patterns through its input-dependent linear flow, which induces non-stationary behavior.
+Moreover, the ETGP shares a single latent GP across outputs, offering improved parameter efficiency and scalability compared to modeling each output with a separate GP, as discussed in Section II-C.
+It is noteworthy that this efficient transformed Gaussian process (ETGP) modeling strategy was originally proposed to tackle large-scale multi-class classification problems [maronas2022efficient].
 Its efficiency arises from employing a single GP instead of one per class, ensuring that the computational cost of GP operations remains independent of the number of classes.
 This results in a substantial reduction in overall computation.
-In this work, we adapt the ETGP framework to model high-dimensional transition functions in SSMs.
-Notably, ETGP implicitly models $d_{x}$ dependent random processes [43], introducing output correlations that are often neglected in existing GPSSM formulations.
+In this work, we adapt this framework to model high-dimensional transition functions in SSMs.
+Notably, ETGP implicitly models $d_{x}$ dependent random processes [ma2019variational], introducing output correlations that are often neglected in existing GPSSM formulations.
 We elaborate further on this below.
 Corollary 1.
-If $p_{\bm{\psi}}(\mathbf{w})=\delta(\mathbf{w}-\bar{\mathbf{w}})$, where $\bar{\mathbf{w}}$ is a deterministic vector, then the ETGP defined in Example 3 becomes a $d_{x}$-dimensional dependent non-stationary Gaussian process with time-dependent covariance.
+If $p_{\bm{\psi}}(\mathbf{w})=\delta(\mathbf{w}-\bar{\mathbf{w}})$, where $\bar{\mathbf{w}}$ is a deterministic vector, then the ETGP defined in Example 4 becomes a $d_{x}$-dimensional dependent non-stationary Gaussian process with an input-dependent covariance function.
 Specifically, at arbitrary time steps $t$ and $t^{\prime}$, we have
@@
 where the covariance matrix $\bm{\Lambda}_{t,t^{\prime}}\in\mathbb{R}^{2d_{x}\times 2d_{x}}$ is given by
-$\bm{\Lambda}_{t,t^{\prime}}\!=\!\begin{bmatrix}k({\mathbf{x}}_{t-1},{\mathbf{x}}_{t-1})\cdot\bm{\alpha}_{t}\bm{\alpha}_{t}^{\top}&k({\mathbf{x}}_{t-1},{\mathbf{x}}_{t^{\prime}-1})\cdot\bm{\alpha}_{t}\bm{\alpha}_{t^{\prime}}^{\top}\\ k({\mathbf{x}}_{t^{\prime}-1},{\mathbf{x}}_{t-1})\cdot\bm{\alpha}_{t^{\prime}}\bm{\alpha}_{t}^{\top}&k({\mathbf{x}}_{{t^{\prime}}-1},{\mathbf{x}}_{{t^{\prime}}-1})\cdot\bm{\alpha}_{t^{\prime}}\bm{\alpha}_{t^{\prime}}^{\top}\end{bmatrix}.$ (15)
+$\bm{\Lambda}_{t,t^{\prime}}\!=\!\begin{bmatrix}k({\mathbf{x}}_{t-1},{\mathbf{x}}_{t-1})\cdot\bm{\alpha}_{t}\bm{\alpha}_{t}^{\top}&k({\mathbf{x}}_{t-1},{\mathbf{x}}_{t^{\prime}-1})\cdot\bm{\alpha}_{t}\bm{\alpha}_{t^{\prime}}^{\top}\\ k({\mathbf{x}}_{t^{\prime}-1},{\mathbf{x}}_{t-1})\cdot\bm{\alpha}_{t^{\prime}}\bm{\alpha}_{t}^{\top}&k({\mathbf{x}}_{{t^{\prime}}-1},{\mathbf{x}}_{{t^{\prime}}-1})\cdot\bm{\alpha}_{t^{\prime}}\bm{\alpha}_{t^{\prime}}^{\top}\end{bmatrix}$.
+(15)
 Proof.
-This result follows from the Gaussianity of the underlying process and the linearity of covariance, as detailed in Appendix C. ∎
+This result follows from the Gaussianity of the underlying process and the linearity of covariance, as detailed in Appendix A. ∎
 With the computationally efficient and flexible function prior defined by ETGP $\bm{f}(\cdot)\!=\!\{f_{d}(\cdot)\}_{d=1}^{d_{x}}$, we define our proposed SSM, termed ETGPSSM, as follows:
@@
 $\mathbf{y}_{t}|\mathbf{x}_{t}\sim{\cal N}({\mathbf{y}}_{t}|\bm{C}\mathbf{x}_{t},\mathbf{R}),$ (16e)
-where the ETGP prior is defined by the base GP $\tilde{f}(\cdot)$, neural network weight prior $p_{\bm{\psi}}(\mathbf{w})$, and the normalizing flows, see Eq.
-(12).
-Unless stated otherwise, in this work we focus on the linear flow described in Example 3 for simplicity and illustration.
-This variant requires learning only two parameters per dimension, offering a more efficient alternative to more expressive but costly flow parameterizations.
-For simplicity, we also assume that the initial state prior $p({\mathbf{x}}_{0})$ is Gaussian and known.
+[…]
 Remark 1.
-In contrast to existing GPSSMs, the $\{f_{d}(\cdot)\}_{d=1}^{d_{x}}$ in the ETGPSSM are dependent, as they are both transformed random processes derived from a shared GP and common weight priors (see Corollary 1).
-Furthermore, as these induced random processes are non-stationary (see Corollary 1), they are well-suited for modeling time-varying transition functions.
+In contrast to existing GPSSMs with independent, stationary transitions, the ETGPSSM yields dependent and non-stationary transition functions $\{f_{d}(\cdot)\}_{d=1}^{d_{x}}$ due to their shared base GP and weight priors.
+In the case of linear input-dependent flows, the transition function remains a GP—like a GPSSM—but enriched with dependence and non-stationarity, see Corollary 1, offering greater modeling flexibility while retaining GP interpretability.
 The primary objective in ETGPSSM, as in GPSSM, is the joint estimation of model hyperparameters $\bm{\vartheta}\!=\!\{{{\bm{\theta}}}_{gp},\bm{\psi},\mathbf{Q},\mathbf{R}\}$ and inference of both the ETGP and latent states.
-Given a sequence of observations ${\mathbf{y}}_{1:T}=\{{\mathbf{y}}_{t}\}_{t=1}^{T}$, with $T\in\mathbb{Z}_{+}$, the task is to estimate $\bm{\vartheta}$ and compute the posterior distributions $p(\bm{f}|{\mathbf{y}}_{1:T})$ for the ETGP and $p({\mathbf{x}}_{0:T}|{\mathbf{y}}_{1:T})$ for the latent states, where ${\mathbf{x}}_{0:T}=\{{\mathbf{x}}_{t}\}_{t=0}^{T}$.
+Given a sequence of observations ${\mathbf{y}}_{1:T}$, the task is to estimate $\bm{\vartheta}$ and compute the posterior distributions $p(\bm{f}|{\mathbf{y}}_{1:T})$ for the ETGP and $p({\mathbf{x}}_{0:T}|{\mathbf{y}}_{1:T})$ for the latent states, where ${\mathbf{x}}_{0:T}=\{{\mathbf{x}}_{t}\}_{t=0}^{T}$.
 However, the marginal likelihood $p(\mathbf{y}_{1:T}|\bm{\vartheta})$ is generally intractable, making both model evidence maximization and posterior inference computationally challenging.
-Furthermore, unlike standard GPs, which admit closed-form expressions for their prior and posterior (see Eqs.
-(2) and (3)), the ETGP defines an implicit stochastic process.
-As a result, both the marginal prior [^4] Obtaining $p(\bm{f})$ by marginalizing over the GP and $\mathbf{w}$ is generally intractable.
-and posterior distributions are analytically intractable [43], presenting significant challenges for inference.
+Furthermore, unlike GPs, which admit closed-form expressions for their prior and (approximate) posterior (see Eqs.
+(2), (3) & (8)), the ETGP defines an implicit stochastic process.
+As a result, both the prior and posterior are analytically intractable [ma2019variational], presenting significant challenges for inference.
 In the following subsections, we introduce a variational inference framework that enables joint learning of model hyperparameters and approximate Bayesian inference over the ETGP and latent states.
 III-B Approximate Bayesian Inference
-To address the intractability of the marginal distribution $p({\mathbf{y}}_{1:T}|\bm{\vartheta})$, we employ variational inference.
-This approach maximizes a tractable evidence lower bound (ELBO) on the log-marginal likelihood, which is equivalent to minimizing the Kullback–Leibler (KL) divergence between the variational approximation and the true posterior [49].
+Variational inference proceeds by maximizing a more tractable evidence lower bound (ELBO) on the log-marginal likelihood, which is equivalent to minimizing the Kullback–Leibler (KL) divergence between the variational approximation and the true posterior [theodoridis2020machine].
 Constructing the ELBO requires specifying both the model joint distribution and the variational distribution.
-Following the existing GPSSM framework and building upon Eq. (16), we factorize the joint distribution as:
+Following the existing GPSSM framework, we factorize the joint distribution of the ETGPSSM defined in Eq. (16) as:
 $\!\!\!p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T},\bm{f})$ $\!=\!p({\mathbf{y}}_{1:T},{\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})p(\bm{f})p({\mathbf{x}}_{0})$ (17)
 $\!=\!\prod_{t=1}^{T}p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})p({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1},\bm{f})p(\bm{f})p({\mathbf{x}}_{0}),$
-where $p({\mathbf{x}}_{0})={\cal N}({\mathbf{x}}_{0}|\bm{0},\bm{I})$ and $p(\bm{f})$ denotes the ETGP prior (under slight abuse of notation).
+where $p({\mathbf{x}}_{0})={\cal N}({\mathbf{x}}_{0}|\bm{0},\bm{I})$, and under slight abuse of notation [^3] In general a random process does not have density function., $p(\bm{f})$ denotes the ETGP prior, which is generally intractable because it requires marginalizing over the base GP $\tilde{f}$ and the neural network weights $\mathbf{w}$.
 A common variational family in the GPSSM literature takes the form:
@@
 $\mathcal{L}_{2}=\mathbb{E}_{q({\mathbf{x}}_{0:T},\bm{f})}\left[\log\frac{p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T},\bm{f})}{q({\mathbf{x}}_{0:T},\bm{f})}\right].$ (19)
-While general, this formulation requires explicit parameterization of the variational distribution $q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})$, which can introduce a large number of parameters and increase optimization complexity [37, 31].
+While general, this formulation requires explicit parameterization of the variational distribution $q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})$, which can introduce a large number of parameters and increase optimization complexity [lin2023ensemble, ialongo2019overcoming].
 To improve learning efficiency, in this work, we instead adopt an alternative approach based on marginalizing out ${\mathbf{x}}_{1:T}$ and infer latent states using an EnKF, avoiding the need to explicitly model this variational term.
@@
 $\mathcal{L}=\mathbb{E}_{q({\mathbf{x}}_{0},\bm{f})}\left[\log\frac{p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0},\bm{f})}{q({\mathbf{x}}_{0})\,q(\bm{f})}\right].$ (22)
-Compared to $\mathcal{L}_{2}$, the alternative bound $\mathcal{L}$ avoids direct parameterization of $q({\mathbf{x}}_{1:T}\mid\bm{f},{\mathbf{x}}_{0})$, helping reduce optimization overhead and yielding a tighter variational lower bound (see Proposition 1).
+Compared to $\mathcal{L}_{2}$, the alternative bound $\mathcal{L}$ avoids direct parameterization of $q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})$, helping reduce optimization overhead and yielding a tighter variational lower bound (see Proposition 1).
 Proposition 1.
@@
 Proof.
-Proof can be found in Appendix A
-Theoretically, this tighter lower bound can improve inference accuracy and convergence [49].
-We now proceed to derive the ELBO $\mathcal{L}$.
+Proof can be found in Appendix B
+Theoretically, this tighter lower bound can improve inference accuracy and convergence [theodoridis2020machine]. Next, we proceed to derive the ELBO $\mathcal{L}$.
 As noted earlier, the lack of closed-form expressions for the ETGP’s prior and posterior prevents direct computation of $\mathcal{L}$ in function space.
@@
 Specifically, by explicitly representing $\tilde{f}$ and $\mathbf{w}$, the marginalized joint distribution becomes:
-$p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0},\mathbf{w},\tilde{f})=p({\mathbf{y}}_{1:T}|{\mathbf{x}}_{0},\mathbf{w},\tilde{f})p({\mathbf{x}}_{0})p_{\bm{\psi}}(\mathbf{w})p(\tilde{f}),$ (23)
-where $p(\tilde{f})$ represents the GP prior and $p_{\bm{\psi}}(\mathbf{w})={\cal N}(\mathbf{w}|\bm{0},\operatorname{diag}(\bm{\psi}))$ is a Gaussian prior over the neural network weights, assuming a Bayesian neural network with diagonal covariance.
+[…]
+where p(f~)p($\tilde${f}) represents the GP prior and $p_{\bm{\psi}}(\mathbf{w})={\cal N}(\mathbf{w}|\bm{0},\operatorname{diag}(\bm{\psi}))$ is a Gaussian prior over the neural network weights, assuming a Bayesian neural network with diagonal covariance.
 Similarly, instead of using Eq.
@@
 where ${{\mathbf{m}}}_{0}\!\in\!\mathbb{R}^{d_{x}}$, lower-triangular matrix $\mathbf{L}_{0}\!\in\!\mathbb{R}^{d_{x}\times d_{x}}$, and $\{\mathbf{w},\bm{\sigma}^{2}_{\mathrm{w}}\}$ are free variational parameters.
-For the variational distribution $q(\tilde{f})$, we adopt a sparse GP approximation [40, 41], as defined in Eq.
+For the variational distribution q(f~)q($\tilde${f}), we adopt a sparse GP approximation [titsias2009variational, hensman2013gaussian], as defined in Eq.
 (8), which helps to reduce the computational complexity from $\mathcal{O}(T^{3})$ to $\mathcal{O}(M^{3})$.
@@
 where the first terms (log-likelihood) can be analytically evaluated using the EnKF (detailed in Section III-C).
-The three KL divergence terms can also be computed in closed form, due to the Gaussian nature of the prior and variational distributions [49].
+The three KL divergence terms can also be computed in closed form, due to the Gaussian nature of the prior and variational distributions [theodoridis2020machine].
 Proof.
-The proof can be found in Appendix B. ∎
+The proof can be found in Appendix C. ∎
 Remark 2.
 The objective in Eq.
-(26) can also be derived from $\mathcal{L}_{2}$ under two specific assumptions (see Supplement A), akin to [37].
+(26) can also be derived from $\mathcal{L}_{2}$ under two specific assumptions (see Supplement A), akin to [lin2023ensemble].
 However, our derivation in this work is generic for arbitrary random processes and is assumption-free, offering greater simplicity, conceptual clarity, and broader applicability.
-For further details on the relationship between $\mathcal{L}$ and $\mathcal{L}_{2}$, we refer to Supplement  A.
+For further details on the relationship between $\mathcal{L}$ and $\mathcal{L}_{2}$, we refer the reader to Supplement A.
 III-C EnKF-aided ELBO Evaluation
-This section describes how to evaluate $p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})$ in Eq. (26) using the EnKF. First, note that
-$p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})$ $=\int\underbracket{p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})}_{\text{likelihood}}\underbracket{p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})}_{\text{prediction}}\mathrm{d}{\mathbf{x}}_{t},$ (27)
+[…]
+[…]
 where the prediction distribution $p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})$ reads:
@@
 with $p({\mathbf{f}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})=\delta\left({\mathbf{f}}_{t}-(\bm{\alpha}_{t}\cdot\tilde{f}({\mathbf{x}}_{t-1})+\bm{\beta}_{t})\right)$, and $\{\bm{\alpha}_{t},\bm{\beta}_{t}\}$ being the outputs of the neural network parameterized by $\mathbf{w}$.
-Consequently, we employ the EnKF, a widely used method for handling nonlinearities and high-dimensional state spaces [50], to estimate the filtering distribution $p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t})$ at each time step $t$.
+Consequently, we employ the EnKF, a widely used method for handling nonlinearities and high-dimensional state spaces [evensen1994sequential], to estimate the filtering distribution $p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t})$ at each time step $t$.
 The details are as follows.
 III-C1 Prediction
-Given the ensemble of states $\{{\mathbf{x}}_{t-1}^{(i)}\}_{i=1}^{N}$ from the posterior distribution $p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})$ at time $t-1$, and conditioning on $\mathbf{w}\sim q(\mathbf{w})$ and $\tilde{f}\sim q(\tilde{f})$, we use the state transition in Eq.
-(29) to generate $N$ predictive samples:
+[…]
 $\bar{{\mathbf{x}}}_{t}^{(i)}\!\sim\!{\cal N}\left({\mathbf{x}}_{t}\mid\bm{\alpha}_{t}\cdot\tilde{f}({\mathbf{x}}_{t-1}^{(i)})+\bm{\beta}_{t},\ \mathbf{Q}\right)$ (30)
@@
 The prediction and update steps are fully differentiable, as the sampled latent states become differentiable with respect to the state transition function $\bm{f}(\cdot)$ and the noise covariances $\mathbf{Q}$ and $\mathbf{R}$.
-This allows for end-to-end gradient-based optimization (e.g.using Adam [51], as described in Algorithm 1), making the framework practical for joint learning and inference [52, 37].
-Based on the outlined prediction and update steps, the likelihood $p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})$ in Eq. (27) can be easily evaluated. At each time step $t$, we have:
+This allows for end-to-end gradient-based optimization (e.g.using Adam [kingma2015adam], as described in Algorithm 1), making the framework practical for joint learning and inference [chen2022autodifferentiable, lin2023ensemble].
+[…]
 $p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})\approx{\cal N}({\mathbf{y}}_{t}\mid\bm{C}\bar{\mathbf{m}}_{t},\ \bm{C}\bar{\mathbf{P}}_{t}\bm{C}^{\top}),$ (35)
@@
 (26).
-At each iteration, we apply the reparameterization trick (see e.g.
-Eq.
-(33b)) to sample from the Gaussian variational distributions $q(\mathbf{w})$ and $q(\tilde{f})$, enabling unbiased Monte Carlo estimates of the expected log-likelihood term $\mathbb{E}_{q(\mathbf{w},\tilde{f})}[\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})]$.
-Thanks to the reparameterization trick [53], the ELBO $\mathcal{L}$ remains differentiable with respect to both the model parameters $\bm{\vartheta}=\{\bm{\theta}_{gp},\bm{\psi},\mathbf{Q},\mathbf{R}\}$ and the variational parameters ${\bm{\zeta}}=\{\mathbf{m}_{0},\mathbf{L}_{0},\mathbf{m}_{\mathrm{w}},\bm{\sigma}_{\mathrm{w}}^{2},\mathbf{m},\mathbf{S},\mathbf{Z}\}$.
-We leverage modern automatic differentiation tools such as PyTorch to compute gradients via backpropagation through time (BPTT), and apply gradient-based optimization methods (e.g., Adam [51]) to maximize $\mathcal{L}$ [54].
+At each iteration, we apply the reparameterization trick to sample from the Gaussian variational distributions $q(\mathbf{w})$, $q(\mathbf{x}_{0})$, and q(f~)q($\tilde${f}).
+For instance, a sample from q(f~)q($\tilde${f}) at input $\mathbf{x}_{t}$ is generated as
+$\tilde{f}(\mathbf{x}_{t})=\xi_{p}(\mathbf{x}_{t})+\sqrt{\Xi_{p}(\mathbf{x}_{t})}\cdot\epsilon,\quad\epsilon\sim\mathcal{N}(0,\,1),$
+where $\xi_{p}(\cdot)$ and $\Xi_{p}(\cdot)$ are given by Eq.
+(9).
+Similarly, for $q(\mathbf{w})$ and $q(\mathbf{x}_{0})$, we also use the standard Gaussian reparameterization.
+This enables unbiased Monte Carlo estimates of the expected log-likelihood term in $\mathcal{L}$.
+And thanks to the reparameterization trick [kingma2019introduction], the ELBO $\mathcal{L}$ remains differentiable with respect to both the model parameters $\bm{\vartheta}\!=\!\{\bm{\theta}_{gp},\bm{\psi},\mathbf{Q},\mathbf{R}\}$ and the variational parameters ${\bm{\zeta}}\!=\!\{\mathbf{m}_{0},\mathbf{L}_{0},\mathbf{m}_{\mathrm{w}},\bm{\sigma}_{\mathrm{w}}^{2},\mathbf{m},\mathbf{S},\mathbf{Z}\}$.
+We leverage modern automatic differentiation tools such as PyTorch to compute gradients via backpropagation through time, and apply gradient-based optimization methods (e.g., Adam [kingma2015adam]) to maximize $\mathcal{L}$ [paszke2019pytorch].
 The complete procedure for training ETGPSSM with EnKF-based latent state inference is summarized in Algorithm 1.
@@
 4:  for $t=1,2,\ldots,T$ do
-5:         Get $\tilde{f}$ from $q(\tilde{f})$ using Eq. (8) with ${\mathbf{x}}_{*}={\mathbf{x}}_{t-1}$
+5:   Get $\tilde{f}({\mathbf{x}}_{t-1}^{(i)})$ using Eq. (8) with ${\mathbf{x}}_{*}={\mathbf{x}}_{t-1}^{(i)},\forall i$
 6:    EnKF
@@
 11:   Evaluate the log-likelihood using Eq. (35), and
-$L_{\ell}\leftarrow L_{\ell}+\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})\vspace{-.1in}$
+[…]
 12:
@@
 14:  Evaluate $\mathcal{L}$ based on $L_{\ell}$ and Eq. (26)
-15:     Maximize $\mathcal{L}$ and update $\bm{\vartheta}$, ${\bm{\zeta}}$ using Adam [51]
+15:  Maximize $\mathcal{L}$ and update $\bm{\vartheta}$, $\bm{\zeta}$} using Adam [kingma2015adam]
 16:end while
 17:Output: EnKF particles $\{{\mathbf{x}}_{0:T}^{(i)}\}_{i=1}^{N}$, model parameters $\bm{\vartheta}$, and variational parameters $\bm{\zeta}$}.
-Remark 4 (Efficiency).
-The proposed ETGPSSM in Algorithm 1 improves upon existing GPSSMs in two key aspects of efficiency.
+Remark 4.
+The proposed ETGPSSM in Algorithm 1 improves upon existing GPSSMs in several key aspects of efficiency.
 First, it reduces the computational complexity from $\mathcal{O}(d_{x}M^{3})$ to $\mathcal{O}(M^{3})$ by relying on a single GP, while the added cost of a shallow neural network remains negligible.
@@
 Furthermore, by replacing explicit variational parameterization of the latent states with EnKF-based inference, the method significantly reduces parametric overhead.
-Together, these design choices lead to substantial gains in both computational and modeling efficiency.
-Empirical results validating these advantages are presented in Section IV-B.
-In the next section, we empirically evaluate the performance of ETGPSSM across a range of dynamical system modeling tasks.
-            Data               AD-EnKF [52] AD-EnKF (bnn)          EnVI [37] ETGPSSM (dnn)      ETGPSSM (bnn)
+Together, these design choices lead to substantial gain in both computational and modeling efficiency.
+The overall computational complexity is $\mathcal{O}(M^{3}+d_{x}^{3})$, where the $\mathcal{O}(d_{x}^{3})$ term comes from the EnKF operations.
+It is important to note that in practice, $M$ is typically selected based on the complexity of the transition dynamics and the available computational budget, resulting in a trade-off between computation and modeling accuracy.
+Nevertheless, further investigation into how $M$ should scale with $d_{x}$ to preserve approximation accuracy in high-dimensional GPSSMs [burt2020convergence] is warranted for future research.
+      Data       AD-EnKF [chen2022autodifferentiable] AD-EnKF (bnn)     EnVI [lin2023ensemble] ETGPSSM (dnn)    ETGPSSM (bnn)
 Figure 4: Non-stationary kink transition function learning performance (mean $\pm$ $2\sigma$) using various methods across different levels of emission noise ($\sigma_{\mathrm{R}}^{2}\in\{0.0008,0.008,0.08,0.8\}$, from top to bottom).
 The blue curve () represents the learned mean function, while the red line () indicates the true system transition function.
-The shaded blue region depicts the confidence interval.
+The shaded region depicts the total predictive uncertainty for the state transition, which includes the estimated transition noise level $(\pm 2\,\hat{\sigma}_{Q})$ for all methods, plus function uncertainty for probabilistic models (BNN/GP).
 IV Experiments
+In this section, we empirically evaluate the performance of ETGPSSM across a range of tasks.
 We begin by evaluating ETGPSSM’s ability to model non-stationary dynamical systems in Section IV-A.
@@
 Lastly, Section IV-C highlights the model’s predictive performance across a range of real-world time series datasets.
-Unless otherwise stated, all GPSSMs use the squared exponential (SE) kernel [9], with $20$ inducing points initialized at random.
+Unless otherwise stated, all GPSSMs use the SE kernel [williams2006gaussian], with $20$ inducing points initialized at random.
 The shallow neural network used in ETGPSSM is a fully connected three-layer architecture with ReLU activations and hidden dimensions $128$ and $64$, mapping from $\mathbb{R}^{d_{x}}$ to the flow parameter space $\Theta\!\subseteq\!\mathbb{R}^{2d_{x}}$.
-Additional implementation details are available in the publicly released code.
+For further implementation details, we refer the reader to the supplementary material and the public code repository.
 [^4] https://github.com/zhidilin/gpssmProj/tree/main/high_dim_GPSSM
@@
 where the non-stationary kink function $f({\mathbf{x}})$ is defined as:
-$f({\mathbf{x}})\!=\!\underbrace{\left[0.8+({\mathbf{x}}+0.2)\left(1-\frac{5}{1+\exp(-2{\mathbf{x}})}\right)\right]}_{\text{``kink'' function}}s({\mathbf{x}})-o({\mathbf{x}})$ (37)
+[…]
 with a slope modulation:
@@
 $o({\mathbf{x}})=\begin{cases}0.5\sin(8{\mathbf{x}}),&{\mathbf{x}}>0,\\ 0.5\sin(2{\mathbf{x}}),&{\mathbf{x}}\leq 0.\end{cases}$ (39)
-The kink function is a standard benchmark for evaluating GPSSMs [27, 31, 37].
-Here, we extend it with non-stationary slope and oscillatory terms to increase modeling complexity (see Fig.
+The kink function is a standard benchmark for evaluating GPSSMs [frigola2014variational, ialongo2019overcoming, lin2023ensemble], and is readily modeled by a stationary GP with an SE kernel.
+Here, we break this stationarity by extending it with input-dependent slope and oscillatory terms, creating a non-stationary benchmark that presents a greater challenge for dynamical models (see Fig.
 4).
 We generate training data ${\mathbf{y}}_{1:T}$ by fixing $\sigma_{\mathrm{Q}}^{2}$ at $0.05$ and varying $\sigma_{\mathrm{R}}^{2}$ systematically across the set $\{0.0008,0.008,0.08,0.8\}$, producing four noise regimes.
-ETGPSSM is compared to two leading baselines: the neural network-based AD-EnKF [52] and the GP-based EnVI [37].
-Consistent with previous works [31, 37], all methods share the true emission model while learning the transition function, ensuring a fair comparison.
+ETGPSSM is compared to two leading baselines: the neural network-based AD-EnKF [chen2022autodifferentiable] and the GP-based EnVI [lin2023ensemble].
+The AD-EnKF employs a fully-connected network with ReLU activations and hidden layers of dimensions 128 and 64, to model the state transition.
+Further implementation details (e.g., learning rate, optimizer) are provided in Supplement B.
+Consistent with previous works [ialongo2019overcoming, lin2023ensemble], all methods share the true emission model while learning the transition function, ensuring a fair comparison.
 The system dynamics learning results across different methods are shown in Fig.
@@
 3).
+Figure 5: (Left) Comparison of the total number of parameters (including variational and model parameters).
+In ETGPSSM, the count is dominated by the GP ($Md_{x}+3+M+M^{2}$) and the neural network with architecture $d_{x}\!\to\!128\!\to\!64\!\to\!2d_{x}$ ($258\,d_{x}+8384$), resulting in overall linear scaling in $d_{x}$.
+In GPSSM, the parameters are dominated by the $d_{x}$ independent GPs, ($Md_{x}^{2}+(3+M+M^{2})d_{x}$), leading to quadratic scaling in $d_{x}$.
+(Right) ETGPSSM maintains low computational costs as $d_{x}$ increases, while GPSSM exhibits linear growth.
 The results in Fig.
@@
 By contrast, AD-EnKF struggles with the smooth left segment and performs well only on the right, highly nonlinear region.
-EnVI, the strongest GPSSM baseline, fails to capture the non-stationarity and underfits the high-frequency component.
+EnVI, the state-of-the-art GPSSM baseline, fails to capture the non-stationarity and underfits the high-frequency component.
 Under high noise $(\sigma_{\mathrm{R}}^{2}=\{0.08,0.8\})$, performance declines across all methods due to increased latent state uncertainty.
@@
 IV-B High-Dimensional Latent State Estimation
-In this subsection, we evaluate latent state estimation in the Lorenz-96 system, a standard benchmark for studying chaotic dynamics in weather and climate modeling [55].
+In this subsection, we evaluate latent state estimation in the Lorenz-96 system, a standard benchmark for studying chaotic dynamics in weather and climate modeling [lorenz1996predictability].
 The system consists of $d_{x}\in\mathbb{Z}_{+}$ coupled ordinary differential equations (ODEs):
-[…]
-where $F=8$ induces fully chaotic behavior, with small differences in initial conditions leading to divergent trajectories [55].
+$\frac{d\mathrm{x}_{d}}{dt}=(\mathrm{x}_{d+1}-\mathrm{x}_{d-2})\mathrm{x}_{d-1}-\mathrm{x}_{d}+F,\ 1\leq d\leq d_{x}$ (40)
+where $F=8$ induces fully chaotic behavior, with small differences in initial conditions leading to divergent trajectories [lorenz1996predictability].
 The dynamics model is discretized using the Euler method, and at each time step $t$, observations are generated through a linear emission model ($\bm{C}\!=\!\bm{I}$) with Gaussian additive noise, where the noise level is significant with the covariance given by $\mathbf{R}=4\bm{I}$.
 We first assess the scalability of ETGPSSM in high-dimensional settings.
-Fig.
-5 compares its parameter growth and runtime to standard GPSSMs, using $100$ inducing points.
+For simplicity, we focus exclusively on the dnn variant of ETGPSSM.
+The bnn variant achieves comparable performance but doubles the number of parameters by introducing variance parameters for each neural network weight.
+This observation aligns with prior work showing that, on sufficiently large datasets, the bnn variant offers no significant performance gain over its deterministic counterpart [maronas2022efficient, maronas2021transforming].
+Figure 5 shows the parameter count and runtime of ETGPSSM and standard GPSSMs, both using 100 inducing points.
 As shown in the left panel, ETGPSSM exhibits linear growth in parameter count with increasing $d_{x}$, while GPSSM grows quadratically.
@@
 These results demonstrate that ETGPSSM remains computationally and parametrically tractable in high-dimensional latent spaces.
-Figure 5: (Left) The number of parameters in ETGPSSM grows linearly with the state dimension $d_{x}$, whereas in GPSSM, it scales quadratically.
-(Right) ETGPSSM maintains low computational costs as $d_{x}$ increases, while GPSSM exhibits linear growth.
-Figure 6: Latent state estimation performance for 100-dimensional Lorenz-96 system, with 9 dimensions uniformly sampled for visualization.
-Root mean square error (RMSE) values for different methods: 0.5468 for EnKF, 1.2028 for ETGPSSM, 1.6187 for AD-EnKF.
-RMSE between latent states and observations: 1.9855.
-We next fix the dimension $d_{x}\!=\!100$ in the Lorenz-96 system and generate observations of length $T\!=\!600$ for training ETGPSSM and its deep neural network-based competitor, AD-EnKF [52].
-The EnVI [37], which already caused out-of-memory errors [^6] All experiments were conducted on a machine equipped with an Intel Xeon W7-3545 v6 processor (24 cores, 2.7 GHz), 128 GB of DDR5 RAM, and an NVIDIA RTX 4500 Ada GPU (24 GB VRAM), running Ubuntu 20.04 LTS.
-at $d_{x}\!=\!30$ with an ensemble size $N\!=\!150$, is excluded due to its prohibitive computational cost in this high-dimensional setting.
-The corresponding state inference performance after training is shown in Fig.
-6, where the performance of EnKF is also included for comparison.
-As depicted in Fig.
-6, our ETGPSSM achieves an RMSE of $1.2028$, outperforming AD-EnKF, which yields an RMSE of $1.6187$.
-Additionally, note that the RMSE of the observations (1.9855) serves as a baseline, demonstrating that ETGPSSM significantly reduces the estimation error compared to the raw observations.
+Figure 6: Latent state estimation performance for a 100-dimensional Lorenz-96 system, with 3 dimensions sampled for visualization.
+Root mean square error (RMSE) values for different methods are 0.4468 for EnKF, 1.1583 for ETGPSSM, and 1.4553 for AD-EnKF.
+RMSE between the true latent states and the observations is 1.9855, which serves as a baseline, indicating the error if observations were used directly as state estimates.
+TABLE I: Comparison of filtering performance in the pendulum system.
+Method rmse ($\downarrow$) coverage spread crps ($\downarrow$) ad-enkf 1.4553 0.3007 0.2885 0.2681 etgpssm 1.1583 0.3587 0.2774 0.2101 enkf 0.4468 0.5866 0.1741 0.0685
+We next fix the dimension $d_{x}\!=\!100$ in the Lorenz-96 system and generate observations of length T=600T\!=\!600 for training ETGPSSM and its deep neural network-based competitor, AD-EnKF [chen2022autodifferentiable].
+The EnVI [lin2023ensemble], which already caused out-of-memory errors [^5] All experiments were conducted on a machine equipped with an Intel Xeon W7-3545 v6 processor (24 cores, 2.7 GHz), 128 GB of DDR5 RAM, and an NVIDIA RTX 4500 Ada GPU (24 GB VRAM), running Ubuntu 20.04 LTS.
+at $d_{x}\!=\!30$ with an ensemble size N=150N\!=\!150, is excluded due to its prohibitive computational cost in this high-dimensional setting.
+The corresponding state inference performance after training is summarized in Fig.
+6 and Table I.
+Table I presents four standard filtering metrics: RMSE, spread, coverage probability, and CRPS [spantini2022coupling, gneiting2007probabilistic].
+Detailed definitions of these metrics are provided in Supplement B.
+For comparative analysis, the performance of EnKF is also included.
+As shown in Fig.
+6 and Table I, ETGPSSM achieves superior performance compared to AD-EnKF in both mean estimation (see e.g.
+RMSE) and uncertainty quantification (see e.g.
+CRPS).
+Notably, with the observation RMSE of 1.9855 serving as a baseline, ETGPSSM demonstrates substantial error reduction relative to the raw observational data.
 This result highlights the superior filtering performance of ETGPSSM in high-dimensional chaotic systems.
-While EnKF achieves the lowest RMSE of $0.5468$, it assumes full knowledge of the dynamics and performs no learning.
+While EnKF achieves the best filtering performance, e.g., lowest RMSE, it assumes full knowledge of the dynamics and performs no learning.
 In contrast, ETGPSSM simultaneously learns the system and infers latent states, offering a more general and powerful solution for complex, high-dimensional systems.
 IV-C Real-World Time Series Prediction
-In this subsection, we evaluate the prediction performance of ETGPSSM on several real-world time series datasets [^7] https://homes.esat.kuleuven.be/~smc/daisy/daisydata.html commonly used for system identification (see Table I).
+In this subsection, we evaluate the prediction performance of ETGPSSM on several real-world time series datasets [^6] https://homes.esat.kuleuven.be/~smc/daisy/daisydata.html commonly used for system identification (see Table II).
 In every dataset, the first half of the sequence is utilized as training data, with the remaining portion designated for testing.
 Standardization of all datasets is conducted based on the training sequence.
-Table I reports the series prediction results, wherein the RMSE is averaged over 50-step ahead forecasting.
-We compare ETGPSSM against state-of-the-art methods, including neural network-based approaches—DKF [7] and AD-EnKF [52]—as well as various GPSSMs: vGPSSM [29], CO-GPSSM [36], PRSSM [30], ODGPSSM [39], VCDT [31], and EnVI [37].
+Table II reports the series prediction results, wherein the RMSE is averaged over 50-step ahead forecasting.
+We compare ETGPSSM against state-of-the-art methods, including neural network-based approaches—DKF [krishnan2017structured] and AD-EnKF [chen2022autodifferentiable]—as well as various GPSSMs: vGPSSM [eleftheriadis2017identification], CO-GPSSM [lin2023towards], PRSSM [doerr2018probabilistic], ODGPSSM [lin2022output], VCDT [ialongo2019overcoming], and EnVI [lin2023ensemble].
 Following previous works, the latent space dimension is set to $d_{x}=4$ for all the methods.
-TABLE I: The 50-step ahead prediction performance in terms of RMSE of different models (all with $d_{x}=4$) on system identification datasets.
+TABLE II: The 50-step ahead prediction performance in terms of RMSE of different models (all with $d_{x}=4$) on system identification datasets.
 Mean and standard deviation of the prediction results are shown as an averaged result over ten different seeds.
@@
 Method Actuator Ball Beam Drive Dryer Gas Furnace
-DKF [7] $1.204\pm 0.250$ $0.144\pm 0.005$ $0.735\pm 0.001$ $1.465\pm 0.087$ $5.589\pm 0.066$
-AD-EnKF [52] $0.705\pm 0.117$ $0.057\pm 0.006$ $0.756\pm 0.114$ $0.182\pm 0.053$ $1.408\pm 0.090$
-AD-EnKF (bnn) $0.705\pm 0.088$ $0.053\pm 0.007$ $0.896\pm 0.088$ $0.155\pm 0.030$ $1.361\pm 0.061$
-vGPSSM [29] $1.640\pm 0.011$ $0.268\pm 0.414$ $0.740\pm 0.010$ $0.822\pm 0.002$ $3.676\pm 0.145$
-CO-GPSSM [36] $0.803\pm 0.011$ $0.079\pm 0.018$ $0.736\pm 0.007$ $0.366\pm 0.146$ $1.898\pm 0.157$
-PRSSM [30] $0.691\pm 0.148$ $0.074\pm 0.010$ $0.647\pm 0.057$ $0.174\pm 0.013$ $1.503\pm 0.196$
-ODGPSSM [39] $0.666\pm 0.074$ $0.068\pm 0.006$ $0.708\pm 0.052$ $0.171\pm 0.011$ $1.704\pm 0.560$
-VCDT [31] $0.815\pm 0.012$ $0.065\pm 0.005$ $0.735\pm 0.005$ $0.667\pm 0.266$ $2.052\pm 0.163$
-EnVI [37] $0.657\pm 0.095$ $0.055\pm 0.002$ $0.703\pm 0.050$ $0.125\pm 0.017$ $1.388\pm 0.123$
-ETGPSSM (dnn) $0.646\pm 0.081$ $0.050\pm 0.003$ $0.668\pm 0.092$ $0.137\pm 0.030$ $1.300\pm 0.052$
-ETGPSSM (bnn) $0.646\pm 0.095$ $0.053\pm 0.002$ $0.703\pm 0.028$ $0.154\pm 0.026$ $1.313\pm 0.048$
-As reported in Table I, ETGPSSM demonstrates strong overall performance.
+DKF [krishnan2017structured] $1.204\pm 0.250$ $0.144\pm 0.005$ $0.735\pm 0.001$ $1.465\pm 0.087$ $5.589\pm 0.066$
+AD-EnKF [chen2022autodifferentiable] $705\pm 0.117 0.0$$057\pm 0.006$ $0.756\pm 0.114$ $0.182\pm 0.053$ $1.408\pm 0.090$
+AD-EnKF (bnn) $705\pm 0.088 0.0$$053\pm 0.007$ $0.896\pm 0.088$ $0.155\pm 0.030$ $1.361\pm 0.061$
+vGPSSM [eleftheriadis2017identification] $1.640\pm 0.011$ $0.268\pm 0.414$ $0.740\pm 0.010$ $0.822\pm 0.002$ $3.676\pm 0.145$
+CO-GPSSM [lin2023towards] $803\pm 0.011 0.0$$079\pm 0.018$ $0.736\pm 0.007$ $0.366\pm 0.146$ $1.898\pm 0.157$
+PRSSM [doerr2018probabilistic] $691\pm 0.148 0.0$$074\pm 0.010$ $0.647\pm 0.057$ $0.174\pm 0.013$ $1.503\pm 0.196$
+ODGPSSM [lin2022output] $666\pm 0.074 0.0$$068\pm 0.006$ $0.708\pm 0.052$ $0.171\pm 0.011$ $1.704\pm 0.560$
+VCDT [ialongo2019overcoming] $815\pm 0.012 0.0$$065\pm 0.005$ $0.735\pm 0.005$ $0.667\pm 0.266$ $2.052\pm 0.163$
+EnVI [lin2023ensemble] $657\pm 0.095 0.0$$055\pm 0.002$ $0.703\pm 0.050$ $0.125\pm 0.017$ $1.388\pm 0.123$
+ETGPSSM (dnn) $646\pm 0.081 0.0$$050\pm 0.003$ $0.668\pm 0.092$ $0.137\pm 0.030$ $1.300\pm 0.052$
+ETGPSSM (bnn) $646\pm 0.095 0.0$$053\pm 0.002$ $0.703\pm 0.028$ $0.154\pm 0.026$ $1.313\pm 0.048$
+As reported in Table II, ETGPSSM demonstrates strong overall performance.
 On the Actuator dataset, both ETGPSSM (dnn) and ETGPSSM (bnn) achieve an RMSE of $0.646$, outperforming EnVI ($0.657$) and significantly improving upon AD-EnKFs, which yield an RMSE of $0.705$.
@@
 On the Drive and Dryer datasets, however, GPSSMs such as PRSSM and EnVI attain slightly better performance.
-This is likely due to the relatively low latent state dimension ($d_{x}=4$), where conventional GPSSMs, including PRSSM, ODGPSSM, and EnVI, remain competitive, as discussed in [37].
+This is likely due to the relatively low latent state dimension ($d_{x}=4$), where conventional GPSSMs, including PRSSM, ODGPSSM, and EnVI, remain competitive, as discussed in [lin2023ensemble].
 Nonetheless, we highlight that although the primary contribution of the proposed ETGPSSM lies in its scalability and efficiency in high-dimensional settings, it also remains highly competitive in low-dimensional scenarios, achieving superior or very comparable performance.
@@
 Extensive experiments across system identification, high-dimensional filtering, and time series forecasting show that ETGPSSM consistently achieves strong performance while maintaining favorable computational efficiency.
-Appendix A Proof of Proposition 1
+Appendix A Proof of Corollary 1
+Proof.
+Similar proof of this corollary has been discussed in the literature; see [maronas2022efficient].
+For ease of reference, we also detail here.
+Consider two arbitrary time steps $t$ and $t^{\prime}$, we have
+$\bm{f}({\mathbf{x}}_{t-1})=\bm{\alpha}_{t}\cdot\tilde{f}({\mathbf{x}}_{t-1})+\bm{\beta}_{t},$ (41)
+$\bm{f}({\mathbf{x}}_{t^{\prime}-1})=\bm{\alpha}_{t^{\prime}}\cdot\tilde{f}({\mathbf{x}}_{t^{\prime}-1})+\bm{\beta}_{t^{\prime}}.$ (42)
+Since the parameters $\bm{\theta}_{t}$ and $\bm{\theta}_{t^{\prime}}$ are deterministic when conditioned on the inputs $\mathbf{x}_{t}$ and $\mathbf{x}_{t^{\prime}}$, the joint distribution of $\bm{f}({\mathbf{x}}_{t-1})$ and $\bm{f}({\mathbf{x}}_{t^{\prime}-1})$ is Gaussian with mean $\bm{\beta}_{t}^{\top},\bm{\beta}_{t^{\prime}}^{\top}]^{\top}$.
+The covariance between $\bm{f}({\mathbf{x}}_{t-1})$ and $\bm{f}({\mathbf{x}}_{t^{\prime}-1})$ is:
+$\bm{\Lambda}_{t,t^{\prime}}$ $=\bm{\alpha}_{t}\cdot\text{Cov}(\tilde{f}({\mathbf{x}}_{t-1}),\tilde{f}({\mathbf{x}}_{t^{\prime}-1}))\cdot\bm{\alpha}_{t^{\prime}}^{\top}$ (43)
+$=k({\mathbf{x}}_{t-1},{\mathbf{x}}_{t^{\prime}-1})\cdot\bm{\alpha}_{t}\bm{\alpha}_{t^{\prime}}^{\top}.$ (44)
+Similarly, the covariance of $\bm{f}({\mathbf{x}}_{t-1})$ with itself is:
+$\text{Cov}(\bm{f}({\mathbf{x}}_{t-1}),\bm{f}({\mathbf{x}}_{t-1}))=k({\mathbf{x}}_{t-1},{\mathbf{x}}_{t-1})\cdot\bm{\alpha}_{t}\bm{\alpha}_{t}^{\top}.$ (45)
+The same logic applies to $\bm{f}({\mathbf{x}}_{t^{\prime}-1})$. Therefore, the covariance matrix $\bm{\Lambda}_{t,t^{\prime}}$ is:
+$\begin{bmatrix}k({\mathbf{x}}_{t-1},{\mathbf{x}}_{t-1})\cdot\bm{\alpha}_{t}\bm{\alpha}_{t}^{\top}&k({\mathbf{x}}_{t-1},{\mathbf{x}}_{t^{\prime}-1})\cdot\bm{\alpha}_{t}\bm{\alpha}_{t^{\prime}}^{\top}\\ k({\mathbf{x}}_{t^{\prime}-1},{\mathbf{x}}_{t-1})\cdot\bm{\alpha}_{t^{\prime}}\bm{\alpha}_{t}^{\top}&k({\mathbf{x}}_{{t^{\prime}}-1},{\mathbf{x}}_{{t^{\prime}}-1})\cdot\bm{\alpha}_{t^{\prime}}\bm{\alpha}_{t^{\prime}}^{\top}\end{bmatrix}$.
+(46)
+Appendix B Proof of Proposition 1
 Proof.
@@
 Equality holds when $q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})=p({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0},{\mathbf{y}}_{1:T})$ for all $\bm{f}$ and ${\mathbf{x}}_{0}$. ∎
-Appendix B ELBO Derivation
+Appendix C Proposition 2: ELBO Derivation
 Proof.
@@
 $=\mathbb{E}_{q}\left[\log\frac{p({\mathbf{y}}_{1:T}|{\mathbf{x}}_{0},\mathbf{w},\tilde{f})p({\mathbf{x}}_{0})p(\mathbf{w})p(\tilde{f})}{q({\mathbf{x}}_{0})q(\mathbf{w})q(\tilde{f})}\right$] (50b)
-$=\mathbb{E}_{q}\left[\log\frac{p({\mathbf{y}}_{1:T}|{\mathbf{x}}_{0},\mathbf{w},\tilde{f})p({\mathbf{x}}_{0})p(\mathbf{w})p(\tilde{f}|{\mathbf{u}})q({\mathbf{u}})}{q({\mathbf{x}}_{0})q(\mathbf{w})p(\tilde{f}|{\mathbf{u}})q({\mathbf{u}})}\right]$ (44c)
+$=\mathbb{E}_{q}\left[\log\frac{p({\mathbf{y}}_{1:T}|{\mathbf{x}}_{0},\mathbf{w},\tilde{f})p({\mathbf{x}}_{0})p(\mathbf{w})p(\tilde{f}|{\mathbf{u}})p({\mathbf{u}})}{q({\mathbf{x}}_{0})q(\mathbf{w})p(\tilde{f}|{\mathbf{u}})q({\mathbf{u}})}\right]$ (50c)
 $=\mathbb{E}_{q}\left[\log p({\mathbf{y}}_{1:T}|{\mathbf{x}}_{0},\mathbf{w},\tilde{f})\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))$
@@
 $\quad-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0})),$ (50e)
-where we augment the GP using inducing points and use the variational sparse GP in Eq.
-(44c); We factorize the joint likelihood into sequential terms in Eq.
-(44e), where for $t=1$ we have the log-term as $\log p({\mathbf{y}}_{1}|{\mathbf{x}}_{0},\mathbf{w},\tilde{f})$.
-∎
-Appendix C Proof of Corollary 1
-Proof.
-Similar proof of this corollary has been discussed in the literature; see [42]. For ease of reference, we also detail here. Consider two arbitrary time steps $t$ and $t^{\prime}$, we have
-$\bm{f}({\mathbf{x}}_{t-1})=\bm{\alpha}_{t}\cdot\tilde{f}({\mathbf{x}}_{t-1})+\bm{\beta}_{t},$ (45)
-$\bm{f}({\mathbf{x}}_{t^{\prime}-1})=\bm{\alpha}_{t^{\prime}}\cdot\tilde{f}({\mathbf{x}}_{t^{\prime}-1})+\bm{\beta}_{t^{\prime}}.$ (46)
-Since ${{\bm{\theta}}}_{t}$ and ${{\bm{\theta}}}_{t^{\prime}}$ are deterministic, the joint distribution of $\bm{f}({\mathbf{x}}_{t-1})$ and $\bm{f}({\mathbf{x}}_{t^{\prime}-1})$ is Gaussian with mean $[\bm{\beta}_{t}^{\top},\bm{\beta}_{t^{\prime}}^{\top}]^{\top}$.
-The covariance between $\bm{f}({\mathbf{x}}_{t-1})$ and $\bm{f}({\mathbf{x}}_{t^{\prime}-1})$ is:
-$\bm{\Lambda}_{t,t^{\prime}}$ $=\bm{\alpha}_{t}\cdot\text{Cov}(\tilde{f}({\mathbf{x}}_{t-1}),\tilde{f}({\mathbf{x}}_{t^{\prime}-1}))\cdot\bm{\alpha}_{t^{\prime}}^{\top}$ (47)
-$=k({\mathbf{x}}_{t-1},{\mathbf{x}}_{t^{\prime}-1})\cdot\bm{\alpha}_{t}\bm{\alpha}_{t^{\prime}}^{\top}.$ (48)
-Similarly, the covariance of $\bm{f}({\mathbf{x}}_{t-1})$ with itself is:
-$\text{Cov}(\bm{f}({\mathbf{x}}_{t-1}),\bm{f}({\mathbf{x}}_{t-1}))=k({\mathbf{x}}_{t-1},{\mathbf{x}}_{t-1})\cdot\bm{\alpha}_{t}\bm{\alpha}_{t}^{\top}.$ (49)
-The same logic applies to $\bm{f}({\mathbf{x}}_{t^{\prime}-1})$. Therefore, the covariance matrix $\bm{\Lambda}_{t,t^{\prime}}$ is:
-$\begin{bmatrix}k({\mathbf{x}}_{t-1},{\mathbf{x}}_{t-1})\cdot\bm{\alpha}_{t}\bm{\alpha}_{t}^{\top}&k({\mathbf{x}}_{t-1},{\mathbf{x}}_{t^{\prime}-1})\cdot\bm{\alpha}_{t}\bm{\alpha}_{t^{\prime}}^{\top}\\ k({\mathbf{x}}_{t^{\prime}-1},{\mathbf{x}}_{t-1})\cdot\bm{\alpha}_{t^{\prime}}\bm{\alpha}_{t}^{\top}&k({\mathbf{x}}_{{t^{\prime}}-1},{\mathbf{x}}_{{t^{\prime}}-1})\cdot\bm{\alpha}_{t^{\prime}}\bm{\alpha}_{t^{\prime}}^{\top}\end{bmatrix}.$ (50)
+[…]
 Supplement D Discussions about ELBO Derivation
-Following the GPSSM literature [37, 31], an alternative derivation of the variational inference algorithm presented in this work can begin with the following ELBO, which explicitly defines the variational distribution over the latent states:
+Following the GPSSM literature [lin2023ensemble, ialongo2019overcoming], an alternative derivation of the variational inference algorithm presented in this work can begin with the following ELBO, which explicitly defines the variational distribution over the latent states:
 $\mathcal{L}_{2}=\mathbb{E}_{q({\mathbf{x}}_{0:T},\bm{f}(\cdot))}\left[\log\frac{p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T},\bm{f}(\cdot))}{q({\mathbf{x}}_{0:T},\bm{f}(\cdot))}\right].$ (51)
@@
 (22) as summarized in Proposition 1 in the main text.
-Next, we show how we can use the methodology proposed in [37] to derive the same ELBO as in Eq.
+Next, we show how we can use the methodology proposed in [lin2023ensemble] to derive the same ELBO as in Eq.
 (26).
 The joint distribution of the proposed ETGPSSM is given by:
-$p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T},\mathbf{w},\tilde{f})=p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T}|\mathbf{w},\tilde{f})p_{\bm{\psi}}(\mathbf{w})p(\tilde{f}).$ (52)
-The term $p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T}|\mathbf{w},\tilde{f})$ is factorized as:
+[…]
+[…]
 $p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T}|\mathbf{w},\tilde{f})\!=\!p({\mathbf{x}}_{0})\!\prod_{t=1}^{T}\underbracket{p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})}_{\text{likelihood}}\underbracket{p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{x}}_{t-1})}_{\text{transition}},$ (53)
@@
 $=\mathbb{E}_{q(\mathbf{w},\tilde{f})}\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0})),$ (56g)
-which is the objective function we have used in our paper. Note that from Eq. (56d) to Eq. (56e), we have made the following two assumptions, which is similar to the work in [37]:
+which is the objective function we have used in our paper. Note that from Eq. (56d) to Eq. (56e), we have made the following two assumptions, which is similar to the work in [lin2023ensemble]:
 Assumption 1.
@@
 $p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})\approx p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t}).$ (58)
-These two approximations used here are similar to those in [37], with the key difference being the focus on the ETGP in this paper rather than the GP.
+These two approximations used here are similar to those in [lin2023ensemble], with the key difference being the focus on the ETGP in this paper rather than the GP.
 The rationale behind these approximations is as follows:
-  • Assumption 1) assumes that the posterior distribution of the latent state at time $t-1$ based on observations up to $t-1$ is nearly identical to that based on observations up to $t$.
+  • Assumption 1) assumes that the posterior distribution of the latent state at time t1t-1 based on observations up to t1t-1 is nearly identical to that based on observations up to $t$.
 This holds particularly well when $t$ is large, as the additional observation at time $t$ has minimal impact on the state estimate due to the accumulated information over time.
@@
 Remark 5.
-It can be observed that, by adopting the approach from [37], which relies on two specific assumptions, one can also derive the same objective function as presented in our main text.
+It can be observed that, by adopting the approach from [lin2023ensemble], which relies on two specific assumptions, one can also derive the same objective function as presented in our main text.
 However, our contribution in this paper lies in providing an alternative perspective—one that is more straightforward and assumption-free—by leveraging only the EnKF for inference.
 This not only simplifies the derivation but also enhances the practicality of the method.
+Supplement E Implementation Details
+Throughout the paper, we use the network architecture and hyperparameters specified in Table III. For further details, please refer to the code repository available online.
+TABLE III: Implementation details for experimental comparisons.
+Component ETGPSSM AD-EnKF [chen2022autodifferentiable] EnVI [lin2023ensemble]
+Transition Model ETGP (1 GP + NN flows) FCN $d_{x}$ independent GPs
+FCN Architecture $d_{x}\to 128\to 64\to 2d_{x}$ $d_{x}\to 128\to 64\to d_{x}$
+Ensemble Size ($N$) 200 200 200
+Optimizer Adam Adam Adam
+Learning Rate 0.005 0.005 0.005
+Batch Size Full sequence Full sequence Full sequence
+Kernel Function SE SE
+Activation Function ReLU ReLU
+Training Epochs 1000 1000 1000
+Early Stopping Yes Yes Yes
+In Section IV-B of the main paper, we adopt the following four performance metrics [spantini2022coupling] to collectively provide a comprehensive assessment of both point estimation accuracy and uncertainty quantification in filtering.
+  • Root Mean Squared Error (RMSE): Captures the average deviation between the estimated states and the true reference states.
+  • Spread: Defined as the root mean trace of the ensemble covariance, this metric reflects the internal variability of the ensemble.
+  • Coverage Probability: Measures how frequently the true states fall within the 95% empirical confidence intervals (constructed from the 2.5% and 97.5% quantiles) of the ensemble marginals.
+  • Continuous Ranked Probability Score (CRPS): A comprehensive scoring rule that quantifies how well the predictive distribution matches the observed outcomes [gneiting2007probabilistic].
+It evaluates the discrepancy between the forecasted and actual cumulative distribution functions, with lower values signifying more accurate predictions.